### v3

# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

Abstract.
Let $S\subset\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ be finite symmetric and assume $S$ generates a group $G$ which is a Zariski-dense subgroup $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$. We prove that the Cayley graphs
$\{\mathcal{C}ay(G(\operatorname{mod}q),S(\operatorname{mod}q))\}_{q\in\mathbb{Z}}$
form a family of expanders.
Tang and Zhang are supported by ECS grant 27307320, GRF grant 17317222 and NSFC grant 12001457 from the second author.
1. Introduction
Let $G=\langle S\rangle$ be a subgroup of $\text{SL}_{n}(\mathbb{Z})$ with a finite symmetric generating set $S$. For a positive integer $q$, let $G_{q}=G(\operatorname{mod}q)$ and $\Delta_{q}:L^{2}(G_{q})\rightarrow L^{2}(G_{q})$ be defined as, for any $f\in L^{2}(G_{q}),$
$\Delta_{q}(f)(v):=\left(\frac{1}{|S|}\sum_{s\in S}f(\pi_{q}(s)\cdot v)\right)-f(v).$
Each $\Delta_{q}$ is a self-adjoint operator, with spectrum
$0=\lambda_{0,q}>\lambda_{1,q}\geq\lambda_{2,q}\cdots$
For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\epsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\epsilon$, $\forall q\in\mathcal{A}$. The existence of a uniform positive spectral gap only depends on $G$ and is independent of the choice of the finite generating set $S$. If $\mathcal{A}=\mathbb{Z}_{+}$, we simply say $G$ has super approximation.
It is well known that a large class of lattices in $\text{GL}_{n}(\mathbb{Z})$ satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for arithmetic lattices and can not deal with a general non-lattice discrete group. A breakthrough came from Bourgain and Gamburd, who developed the so-called “Bourgain-Gamburd expansion machine”, which is an analytic-combinatorial tool that allows them to prove super approximation property for any Zariski-dense subgroup of $\text{SL}_{2}(\mathbb{Z})$ with respect to prime moduli [BG08b]. A critical ingredient in the proof is Helfgott’s triple product theorem [Hel08]. Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [Gol19], [HdS21].
Regarding the most general possible linear groups having super approximation, in [GV12] Salehi-Golsefidy and Varjú conjectures:
Conjecture 1.1 (Question 2, [GV12]).
Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated. Then $G$ has the super approximation property if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
In the same paper [GV12], Salehi-Golsefidy-Varjú proved that $G$ has super approximation with respect to square free numbers if $\mathbb{G}_{0}$ is perfect. Later Salehi-Golsefidy generalized to bounded powers of square free numbers:
Theorem A.
(Salehi-Golsefidy)[Gol19] Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated, then $G$ has the super approximation property with respect to bounded powers of square free integers coprime to $q_{0}$ if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
In the paper [Gol19], the author also claims super approximation for $G$ over $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$ for a fixed prime $p$. There appears an issue unless $\mathbb{G}_{0}$ has simple closure. To address this issue, we introduce a gluing technique, which turns out to be applicable to various other settings. We view this gluing technique may be the most novel part of this paper. More detailed discussion is given in Section 2.
If we require no restriction on moduli, all known results have restrictions on the Zariski closure. In [BV12] Bourgain-Varjú proved Conjecture 1.1 when the Zariski closure of $\mathbb{G}_{0}$ is $\text{SL}_{d}$, using an extra ingredient from homogeneous dynamics [BFLM11]. Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with simple closure [HdS21].
The full generality of Conjecture 1.1, despite abundant evidence, has still remained a technical challenge. In the paper [Gol20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers. In a separate paper [TZ23], we proved this conjecture:
Theorem B (Tang-Zhang, 2023).
Suppose $0<\delta\ll 1$, $d$ is a positive integer, and $N_{0}\gg_{d,\delta}1$ is a positive integer. Then there are $0<\varepsilon:=\varepsilon(\delta,d)$ and positive integers $C_{1}=C_{1}(\delta,d),C_{2}=C_{2}(\delta,d),C_{3}=C_{3}(\delta,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$. Suppose $\mathfrak{a}$ is an ideal of $\mathcal{O}$ such that $N(\mathfrak{a}):=|\mathcal{O}/\mathfrak{a}|\geq N_{0}$, and suppose $A\subseteq\mathcal{O}$ such that
$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\delta}.$
Then there are an ideal $\mathfrak{a}^{\prime}$ of $\mathcal{O}$, and $a\in\mathcal{O}$ such that
$\mathfrak{a}^{C_{1}}\subseteq\mathfrak{a}^{\prime},$
$\pi_{\mathfrak{a}^{\prime}}(\mathbb{Z}a)\subset\pi_{\mathfrak{a}^{\prime}}\left(\sum_{C_{3}}A^{C_{2}}-\sum_{C_{3}}A^{C_{2}}\right),$
$|\pi_{\mathfrak{a}^{\prime}}(\mathbb{Z}a)|\geq N(\mathfrak{a})^{\varepsilon}.$
Here, $\pi_{\mathfrak{a}}\text{ (resp. }\pi_{\mathfrak{a}^{\prime}})$ is the reduction map $\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}\text{ (resp. }\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
Roughly speaking, Conjecture 1.1 is a quantitative statement that given a set $A\subset\mathcal{O}$ and an ideal $\mathfrak{a}\subset\mathcal{O}$, if $|\pi_{\mathfrak{a}}(A)|$ is not too small, then modulo an ideal $\mathfrak{a}^{\prime}$ comparable to $\mathfrak{a}$, a sum-product set of $A$ contains a thick arithmetic progression. With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the spirit of [BG09]. Combining Theorem B with a gluing technique which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
Theorem 1.2.
Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation property with respect to all positive integers.
Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple clousre case, and the first non-semisimple closure case.
One advantage of our approach using sum-product, compared to the method in [BV12] and [HdS21] is that we can allow denominators for the group. With little modification of the proof, we have the following extension of Theorem 1.2:
Theorem 1.3.
Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\times{\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})$ $({\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\ltimes(\frac{1}{q_{0}}\mathbb{Z})^{2}$, respectively)))), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)))), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
Acknowledgements
We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun and Tran Chieu-Minh for many helpful corrections/suggestions on previous versions of this paper.
2. Notations and methodology
We introduce the following notations which we use throughout this paper.
The unit of any multiplicatively written group is denoted by 1. Occasionally, if a ring structure is present, we denote the additive unit by 0. For given two subsets $A$ and $B$, we denote their product set by $A\cdot B=\{ab|a\in A,b\in B\}$, and their sum set by $A+B=\{a+b|a\in A,b\in B\}$. The $k$-fold product of $A$ is denoted by $A^{k}$, and the $k$-fold sum of $A$ is denoted by $\sum_{k}A$.
If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by $f*g$ their convolution
$f*g(x)=\sum_{y\in G}f(y)g(xy^{-1}).$
We write $f^{(l)}$ for the $k$-fold convolution of $f$ with itself.
For a prime $p$, we write $p^{n}||q$ if $p^{n}|q$ but $p^{n+1}\nmid q$, and we define $v_{p}(q)=n$. Similarly, for a matrix $M$, $v_{p}(M)=n$ if $p^{n}|M$ but $p^{n+1}\nmid M$. For a prime ideal $\mathcal{P}$ in a ring of integers $\mathcal{O}$, we can define $v_{\mathcal{P}}$ in an analogous way. For two integers $q_{1}$ and $q_{2}$, we write $q_{1}||q_{2}$ if for every $p^{n}||q_{1}$, we also have $p^{n}||q_{2}$. The exact division ``||"``||"` ` | | " can be extended to ideals in a natural way. For a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of a ring $\mathcal{O}$, we write $\mathcal{P}^{n}||\mathfrak{a}$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$. Similarly, for two ideals $\mathcal{Q}_{1}\supset\mathcal{Q}_{2}$, we write $\mathcal{Q}_{1}||\mathcal{Q}_{2}$ if for all $\mathcal{P}^{n}||\mathcal{Q}_{1}$, we also have $\mathcal{P}^{n}||\mathcal{Q}_{2}$.
For $q=\prod_{i}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a real positive number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $[n_{i}\alpha]$ is the integer part of $n_{i}\alpha$.
Let $\pi_{q}:\mathbb{Z}\rightarrow q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the reduction map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$. Sometimes we need to reduce the two factors of $\Lambda$ by two different moduli $q_{1},q_{2}$, and we denote this reduction by $\pi_{q_{1},q_{2}}$. Throughout our paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$.
We use the standard asymptotic notations $O,\Omega,\Theta$ to describe growth of functions. For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large. Similarly, $f(q)>q^{c-}$ to mean $f(q)>q^{c-\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large.
Let $\chi_{S}$ be the normalized uniform counting measure supported on $S$, i.e., for $A\subset G$, $\chi_{S}(A)=\frac{|A\cap S|}{|S|}$. Let $\pi_{q}[\chi_{S}]$ be the pushforward of $\chi_{S}$ under the residue map $\pi_{q}$. Let $T_{q}$ be the convolution operator by $\pi_{q}[\chi_{S}]$, i.e., For $f\in L^{2}(G_{q})$,
$T_{q}(f)=\pi_{q}[\chi_{S}]*f.$ (2.1)
Then $T_{q}$ is a self adjoint operator on $L^{2}(G_{q})$ with an invariant subspace $l_{0}^{2}(G_{q})$ consisting of functions with average 0. Denote the set of eigenvalues of $T_{q}$ on $l_{0}^{2}(G_{q})$ by $E_{q}$. It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$. Then following the argument in [BV12] (Proof of Theorem 1, Page 156-158), Theorem 1.2 can be derived from the following proposition:
Proposition 2.2.
Let $S$ be a symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in ${\text{SL}_{2}\times\text{SL}_{2}}$ or ${\text{ASL}_{2}}$. Then for any $\epsilon>0$ there are $10<\delta<$ 1, $C>0$ depending on $\epsilon$ such that the following holds. If $q\in\mathbb{Z}_{+}$ sufficiently large, $A\subset\Lambda$ symmetric, and some integer $l$ satisfying
$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\epsilon},$ (2.3)
then
$|\pi_{q}(A\cdot A\cdot A)|>|\pi_{q}(A)|^{1+\delta}.$ (2.4)
It is noted that the reduction in [BV12] uses bounds for multiplicities of nontrivial irreducible representations appearing in the regular representation of $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$. Similar bounds exist for $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ and $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q\mathbb{Z})^{2}$ as well.
In the rest of the paper we will focus on proving Proposition 2.2. We give a sketch here.
2.1. Sketch of proof for Proposition 2.2
We assume (2.4) does not hold for a sufficiently small $\delta$. We will then show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\epsilon/2}$, which will force (2.4) to hold afterall.
We assume $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and we let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components. For simplicity we assume each prime divisor of $q$ has sufficiently large exponent. On the first stage, we follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q^{\prime}$ of $q$ and some constant $C$ such that one simple factor, say $\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q^{\prime}}$ (Proposition 5.1). In this process, compared to [BG09], we apply Theorem A to take a shortcut to deduce some results on non-concentration of random walks at Archimedean place.
However, for the purpose of proving Proposition 2.2, we need to take $q^{\prime}$ a very large divisor of $q$, and this leads to the following question:
Suppose for two exact divisors $q_{1},q_{2}$ of $q$, we have a set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
We need to deal with the case $q_{1},q_{2}$ coprime for the above question when growing $q^{\prime}$ to a very large divisor of $q$ for one simple factor, and we also need to consider the case $q_{1},q_{2}$ non co-prime when passing from one simple factor to two simple factors.
In the paper [Gol19], the author also claims super approximation for a general group with perfect Zariski closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$. We are reluctant to cite this result unless the group has simple closure, as there seems to be an issue in the paper [SG17] where this claim is based on, and this issue already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$. By first applying sum-product which creates thick one-parameter groups in one simple factor while losing control of the other factor, and then taking commutators, the author essentially arrives at the situation of the aforementioned question for the case $q_{1}=q_{2}=p^{n}$. It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ is large implies $B$ is large, for which simple counter examples can be found. See the sentence “Lemma 71 and Lemma 34 imply Proposition 37” (Pg. 7254, [SG17]) and trace from there.
To address this question, our approach is as follows. For simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma_{q_{2}}$. We consider a section map
$\psi:\Gamma_{q_{1}}\rightarrow B$
such that $\mathbb{P}_{1}\circ\psi$ is the identity map. Take a small parameter $10<\theta<$ 1, e.g. $\theta=\frac{1}{1000}$. According to a dichotomy (Proposition 3.1), there are two scenarios:
  1. (1)
    There exists a large exact divisor $q_{3}$ of $q_{2}$, and $x,y\in\Gamma_{q_{1}}$ such that for any $p^{n}||q_{3}$, $\psi(xy)\neq\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
  2. (2)
    There exists a large exact divisor ${q_{4}}$ of $q_{2}$, and a large subset $S$ of $\Gamma_{q_{1}}$ such that for any $x,y\in S$, any $p^{n}||q_{4}$, $\psi(xy)=\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
In the first scenario, one can take $B^{\prime}$ to be the collection of all conjugates of $\psi(xy)\psi(x)^{-1}\psi(y)^{-1}$ by $B$ so that $\mathbb{P}_{1}(B^{\prime})=1$ and $\pi_{q_{3}}\circ\mathbb{P}_{2}(B^{\prime})$ is large, thus $\pi_{q_{1},q_{3}}(BB^{\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{3}}$.
In the second scenario, we consider two cases. If for a large exact divisor $q_{4}^{\prime}$ of ${q_{4}}$, we have $\pi_{q_{3}^{\{\theta/2\}}}\circ\mathbb{P}_{2}\circ\psi=1$, then by taking commutator of $\psi(\Gamma_{q_{1}})$ iteratively one can create a set $B^{\prime\prime}$ such that $\pi_{q_{4}^{\prime}}\circ\mathbb{P}_{2}(B^{\prime\prime})=1$ and $\mathbb{P}_{1}(B^{\prime\prime})$ is large in $\Gamma_{q_{1}}$, so that $\pi_{q_{1},q_{4}^{\prime}}(BB^{\prime\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime}}$.
If for a large exact divisor $q_{4}^{\prime\prime}$ of ${q_{4}}$ such that for any $p^{n}||q_{4}^{\prime\prime}$, we have $\pi_{q_{3}^{\{{\theta}/{2}\}}}\circ\mathbb{P}_{2}\circ\varphi\neq 1$, then we must have $q_{4}^{\prime\prime}||q_{1}$, and we can find a set $B^{\prime\prime\prime}\subset\text{Im}(\psi)$ such that $B^{\prime\prime\prime}$ contains all the preimage of a set $\mathcal{P}\subset\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$ under the following canonical projection:
$\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}\cong\Gamma_{q_{1}/q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$
where $Q_{1},Q_{2}|q_{4}^{\prime\prime}$ such that for each $p^{n}||q_{4}^{\prime\prime}$, $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta)$, and $\mathcal{P}=\{(1,1)+n(Q_{1}X_{1},Q_{2}X_{2})\}_{n\in\mathbb{Z}}$, where $X_{1},X_{2}$ are two traceless primitive matrices in $\text{Mat}(2,\mathbb{Z})$. Then one can apply quantitative non-concentration of random walks to find elements from $A$ to conjugates $(X_{1},X_{2})$ to different directions. Multiplying these conjugates together and taking commutators, we can recover a large set in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}$.
This whole scheme is carried out in the proof of Proposition 6.1. One can then apply Proposition 6.1 iteratively to create a large subset of $\Lambda_{q}$, which forces (2.4) to hold afterall.
The idea of considering section maps comes from Bourgain’s gluing scheme for $\mathbb{Z}_{q}$ (See Lemma 6.71 of [Bou08]). If the group is non-semismple, this method can also be applied to connect the semisimple part of the group to the unipotent part. See Section 8.
Remark 2.5.
Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can skip Section 5 and start from Section 6, and reference back to more or less standard lemmas in Sections 3 and 4 when necessary. We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3) and for potential further generalization.
3. Preliminaries on combinatorics
The first ingredient is a generalization of Corollary 6.9 in [Bou08] which proved the case $G_{1},G_{2}$ abelian and $\text{gcd}(|G_{1}|,|G_{2}|)=1$.
Proposition 3.1.
Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
|$\left|\left\{(x,y)\in G_{1}\times G_{1}\mid\psi(xy)=\psi(x)\psi(y)\right\}\right|<\left(1-\varepsilon\right)\left|G_{1}\right|^{2},$ (3.2)
or there exists a subset $S\subset G_{1}$ with $|S|>(1-\sqrt{\varepsilon})|G_{1}|$ and a group homomorphism $f:G_{1}\rightarrow G_{2}$ such that
$\psi|_{S}=f|_{S}.$ (3.3)
Remark 3.4.
If $\text{gcd}\left(\left|G_{1}\right|,\left|G_{2}\right|\right)=1$, then (3.3) implies
|$\left|\left\{x\in G_{1}\mid\psi(x)\neq 1\right\}\right|<\sqrt{\varepsilon}\left|G_{1}\right|.$
Proposition 3.1 relies on the following Lemma 3.5 and Theorem 3.6.
Lemma 3.5.
Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $40<\varepsilon<1/$ 4, such that
$|\mathcal{G}|>(1-\varepsilon)|A|^{2}$
Then there exists a subset $A^{\prime}$ of $A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})|A|$
and
$\left|A^{\prime}A^{\prime}\right|<\frac{|A\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}A|^{4}}{(1-\sqrt{\varepsilon})(1-2\sqrt{\varepsilon})^{2}|A|^{3}},$
where $A\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}A:=\{ab\mid a,b\in A,(a,b)\in\mathcal{G}\}$.
Proof of Lemma.
3.5. The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
Theorem 3.6 (Noncommutative Freiman-Kneser theorem for small doubling).
Let $0<\varepsilon\leq 1$, and let $S\subset G$ be a finite non-empty subset of a multiplicative group $G$ such that $|A\cdot S|\leq(2-\varepsilon)|S|$ for some finite set $A$ of cardinality $|A|$ at least $|S|$, where $A\cdot S:=\{as:a\in A,s\in S\}$ is the product set of $A$ and $S$. Then there exists a finite subgroup $H$ of $G$ with cardinality $|H|\leq C(\varepsilon)|S|$, such that $S$ is covered by at most $C^{\prime}(\varepsilon)$ right-cosets $H\cdot x$ of $H$ for some $C(\varepsilon),C^{\prime}(\varepsilon)\leq{2}/{\varepsilon}-1$.
Remark 3.7.
Theorem 3.6 is due to Hamidoune. See the article “Hamidoune’s Freiman-Kneser theorem for nonabelian groups” in Terence Tao’s blog which gives a concise proof of Theorem 3.6.
Proof of Proposition 3.1.
[…]
$\mathcal{G}=\left\{(x,y)\in G_{1}\times G_{1}\mid\psi(xy)=\psi(x)\psi(y)\right\}$
satisfies
$|\mathcal{G}|\geq\left(1-\varepsilon\right)\left|G_{1}\right|^{2}.$
Denote
$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2}$
and let $\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}$. Then $\mathcal{G}^{\prime}$ is a graph on $A$ and $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\epsilon)|G_{1}|^{2}=(1-\epsilon)|A|^{2}$. Apply Lemma 3.5 with $Z=G_{1}\times G_{2}$. We obtain a subset $A^{\prime}\subset A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.8)
and
$\left|A^{\prime}A^{\prime}\right|<\frac{\left|G_{1}\right|}{\left(1-\sqrt{\varepsilon}\right)\left(1-2\sqrt{\varepsilon}\right)^{2}}<\left(1+10\sqrt{\varepsilon}\right)\left|A^{\prime}\right|<\frac{5}{4}\left|A^{\prime}\right|.$ (3.9)
Next, apply Theorem 3.6 to $A^{\prime}\subset G_{1}\times G_{2}$. There is a subgroup $H$ of $G_{1}\times G_{2}$ and $\left(x_{1},x_{2}\right)\in G_{1}\times G_{2}$ such that
$A^{\prime}$ $\subset\left(x_{1},x_{2}\right)\cdot H,$ (3.10)
$|H|$ $<\frac{5}{3}\left|A^{\prime}\right|.$ (3.11)
[…]
Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
[…]
implying that $H_{1}=G_{1}$. Then for any $x\in G_{1},$ there exists $f(x)\in G_{2}$ such that $(x,f(x))\in H$. Assume there exists $(y,z_{1}),(y,z_{2})\in H$ with $z_{1}\neq z_{2}$. Then for any $x\in G_{1}$, we have $(x,f(x)z_{1}z_{2}^{-1})\in H$ with $(x,f(x)z_{1}z_{2}^{-1})\neq(x,f(x))$. So $|H|\geq 2|G_{1}|\geq 2|A^{\prime}|$. Contradiction. Hence the choice of $f(x)$ is unique for all $x$. Since $H$ is a subgroup, we get $(1,1)\in H$ so $f(1)=1$. Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$, so $f$ is a group homomorphism.
Since
$\left|\mathbb{P}_{1}\left(A^{\prime}\right)\right|=|A^{\prime}|>\left(1-\sqrt{\varepsilon}\right)\left|G_{1}\right|,$
we deduce
$\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|\geq|\mathcal{G}|+|\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)|-|G_{1}|^{2}>\left(1-2\sqrt{\varepsilon}\right)\left|G_{1}\right|^{2}$
by inclusion-exclusion; and hence
$\left|\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right|\geq\frac{\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|}{|G_{1}|}>(1-2\sqrt{\varepsilon})\left|G_{1}\right|.$
Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$, and for $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
$x_{2}f(x_{1}^{-1}y_{1}y_{2})=\psi\left(y_{1}y_{2}\right)=\psi\left(y_{1}\right)\psi\left(y_{2}\right)=x_{2}f(x_{1}^{-1}y_{1})x_{2}f(x_{1}^{-1}y_{2}).$
$\Rightarrow f(x_{1})=x_{2}\Rightarrow(x_{1},x_{2})\in H\Rightarrow A^{\prime}\subset H.$
We finish the proof of Proposition 3.1 by taking $S=\mathbb{P}_{1}\left(A^{\prime}\right)$. ∎
We also need the following bounded generation result over $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
Proposition 3.12.
For any $0<\delta<\frac{1}{25}$, the following holds: Let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ be symmetric and $|A|>(q_{1}q_{2})^{3-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{40\delta}$ such that
$A^{2880}\supset\mathrm{SL}_{2}(q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}).$
In [Hel08] Helfgott proved that given $A\subset\text{SL}_{2}(\mathbb{F}_{p})$ with $|A|>p^{\frac{8}{3}}$, then $A\cdot A\cdot A={\text{SL}}_{2}(\mathbb{F}_{p})$ for a sufficiently large prime $p$. Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to adapt well for a general composite modulus. We follow Helfgott’s approach for the proof of Proposition 3.12.
We start with a few lemmas.
Lemma 3.13.
Let $0<\gamma<\frac{1}{4}$ and let $A,B\in\mathbb{Z}/q\mathbb{Z}$ with $|A|,|B|>q^{1-\gamma}$, then there exists $q^{\prime}|q,q^{\prime}<q^{\frac{12\gamma}{5}}$ such that
$q^{\prime}\mathbb{Z}/q\mathbb{Z}\subset\sum_{24}AB-AB.$
Proof.
This is a special case of Corollary A.7 in [TZ23]. ∎
We need a two dimensional analog of Lemma 3.13.
Lemma 3.14.
Let $0<\delta<\frac{1}{8}$. Let $q_{1},q_{2}\in\mathbb{Z}$ and $A,B$ be subsets of $\mathbb{Z}/q_{1}\mathbb{Z}\times\mathbb{Z}/q_{2}\mathbb{Z}$ such that $|A|,|B|>(q_{1}q_{2})^{1-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{10\delta}$, such that
$\sum_{96}AB-AB\supset q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}\times q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}.$ (3.15)
Proof.
Let $\mathbb{P}_{i},i=1,2$ be the projection from $\mathbb{Z}/q_{1}\mathbb{Z}\times\mathbb{Z}/q_{2}\mathbb{Z}$ to the $i$-th component. Without loss of generality we assume $q_{1}=q_{2}^{\alpha}$ for some $0<\alpha\leq 1$. Since $|A|>(q_{1}q_{2})^{1-\delta}$, there is $x_{0}\in A$ such that
|$|\{x\in A:\mathbb{P}_{1}(x)=x_{0}\}|>(q_{1}q_{2})^{1-\delta}/q_{1}>q_{2}^{1-\delta-\alpha\delta},$
so there is $A^{\prime}\subset A-A$, $|A^{\prime}|>q_{2}^{1-\delta-\alpha\delta}$, $\mathbb{P}_{1}(A^{\prime})=\{0\}$.
Similarly, there is $B^{\prime}\subset B-B$, $|B^{\prime}|>(q_{1}q_{2})^{1-\delta}/q_{1}>q_{2}^{1-\delta-\delta\alpha}$, $\mathbb{P}_{1}(B^{\prime})=0$. Applying Lemma 3.13, we obtain $q_{2}^{\prime}|q_{2},q_{2}^{\prime}<q_{2}^{\frac{12(\delta+\alpha\delta)}{5}}$ such that
$\sum_{48}(AB-AB)\supset\sum_{24}(A^{\prime}B^{\prime}-A^{\prime}B^{\prime})\supset\{0\}\times q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}.$ (3.16)
If $\alpha<5\delta$, then one can take $q_{1}^{\prime}=q_{1}$ and we have $q_{1}^{\prime}q_{2}^{\prime}<q_{2}^{10\delta}$, and (3.16) gives Lemma 3.13.
If $\alpha>5\delta$, then there exists $A^{\prime\prime}\subset A-A,B^{\prime\prime}\subset B-B$, such that $|A^{\prime\prime}|,|B^{\prime\prime}|>q_{1}^{1-\delta-\frac{\delta}{\alpha}},\mathbb{P}_{2}(A^{\prime\prime})=\mathbb{P}_{2}(B^{\prime\prime})=\{0\}$. The exponent $1-\delta-\delta/\alpha$ exceeds $3/4$, so applying Lemma 3.13, we obtain $q_{1}^{\prime}|q_{1}$, $q_{1}^{\prime}<q_{1}^{\frac{12(\delta+\frac{\delta}{\alpha})}{5}}$, such that
$\sum_{48}(AB-AB)\supset\sum_{24}(A^{\prime\prime}B^{\prime\prime}-A^{\prime\prime}B^{\prime\prime})\supset q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}\times\{0\}.$ (3.17)
Adding (3.16) and (3.17), we obtain Lemma 3.15, with
$q_{1}^{\prime}q_{2}^{\prime}<q_{1}^{\frac{12(\delta+\frac{\delta}{\alpha})}{5}}q_{2}^{\frac{12(\delta+\alpha\delta)}{5}}=(q_{1}q_{2})^{\frac{24\delta}{5}}.$
Proof of Proposition 3.12.
. Since $|A|>(q_{1}q_{2})^{3-\delta}$, by the pigeon hole principle, there exists $\vec{v}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{2},\vec{w}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$, such that the cardinality of the set
$\{(\gamma_{1},\gamma_{2}):\gamma_{1}\text{ has lower row }\vec{v},\gamma_{2}\text{ has lower row }\vec{w}\}$
exceeds $(q_{1}q_{2})^{1-\delta}$. This implies the cardinality of the set
$A_{1}:=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}1&m\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)\right):m\in\mathbb{Z}/q_{1}\mathbb{Z},n\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}$
exceeds $(q_{1}q_{2})^{1-\delta}$.
Similarly, if we let
$A_{2}:=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}1&0\\ m&1\end{matrix}\right),\left(\begin{matrix}1&0\\ n&1\end{matrix}\right)\right):m\in\mathbb{Z}/q_{1}\mathbb{Z},n\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}.$
Then, $|A_{2}|>(q_{1}q_{2})^{1-\delta}$.
Define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as:
$(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$
if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$. There are at most $q_{1}q_{2}$ many such classes. By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta}$ many elements from $A$. This implies the set
$H_{0}=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},x\in\mathbb{Z}/q_{1}\mathbb{Z},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*},y\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}$
has cardinality $>(q_{1}q_{2})^{2-\delta}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
$H=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*}\right\}$
has cardinality $>(q_{1}q_{2})^{1-\delta}$.
We have the following elementary computation
$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
Using (3.18) and applying Lemma 3.14 to the set
$\left\{(\lambda_{1}^{2},\lambda_{2}^{2}):\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right)\in H\right\}$
and
$\left\{(m,n):\left(\left(\begin{matrix}1&m\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)\right)\in A_{1}\right\}$
with exponent $1-2\delta$, we obtain $Q_{1}|q_{1},Q_{2}|q_{2}$, $Q_{1}Q_{2}<(q_{1}q_{2})^{20\delta}$ such that
$(A\cdot A^{-1})^{288}\supset\left\{\left(\left(\begin{matrix}1&Q_{1}\mathbb{Z}/q_{1}\mathbb{Z}\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&Q_{2}\mathbb{Z}/q_{2}\mathbb{Z}\\ 0&1\end{matrix}\right)\right)\right\}$
Similarly, we can obtain $Q_{1}^{\prime}|q_{1},Q_{2}^{\prime}|q_{2}$, $Q_{1}^{\prime}Q_{2}^{\prime}<(q_{1}q_{2})^{20\delta}$ such that
$(A\cdot A^{-1})^{288}\supset\left\{\left(\left(\begin{matrix}1&0\\ Q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}&1\end{matrix}\right),\left(\begin{matrix}1&0\\ Q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}&1\end{matrix}\right)\right)\right\}$
Let $Q_{1}^{*}=\text{lcm}(Q_{1},Q_{1}^{\prime})$, $Q_{2}^{*}=\text{lcm}(Q_{2},Q_{2}^{\prime})$. It is an elementary exercise to check that for $m\leq{n}/{2}$, any element of the group
$\left\{\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z}):a,d\equiv 1\ (\operatorname{mod}p^{2m}),b,c\equiv 0\ (\operatorname{mod}p^{2m})\right\}$
can be written as a $a_{1}b_{1}a_{2}b_{2}a_{3}$, where $a_{1},a_{2},a_{3}\in\left(\begin{matrix}1&p^{m}\mathbb{Z}/p^{n}\mathbb{Z}\\ 0&1\end{matrix}\right)$ and $b_{1},b_{2}\in\left(\begin{matrix}1&0\\ p^{m}\mathbb{Z}/p^{n}\mathbb{Z}&1\end{matrix}\right)$. From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{*})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{*})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
$(A\cdot A^{-1})^{1440}\supset\mathrm{SL}_{2}(q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}).$
Following the same method, we can prove
Lemma 3.19.
For any $0<\delta<\frac{1}{10}$, the following holds: Let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ be symmetric and $|A|>q^{3-\delta}$. Then there exists $q^{\prime}|q,q^{\prime}<q^{10\delta}$ such that
$A^{720}\supset\Gamma(q^{\prime})/\Gamma(q).$
We need Lemma 3.19 to prove the following bounded generation result for $\text{ASL}_{2}$.
Proposition 3.20.
For any $0<\delta<10^{-4}$, the following holds: Let $q_{2}|q_{1}$ and let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$ be symmetric and $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}<q_{1}^{50\delta},q_{2}^{\prime}<q_{1}^{55\delta}$, such that
$A^{7204}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
Proof.
Define an equivalence relation on $A$:
$(\gamma_{1},v_{1})\sim(\gamma_{2},v_{2})\text{ if and only if }\gamma_{1}^{-1}v_{1}=\gamma_{2}^{-1}v_{2}.$
Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is bounded by $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$. Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 00. Applying Lemma 3.19, we have
$A^{1440}\supset(Bb^{-1})^{720}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
for some $q_{1}^{\prime}<q_{1}^{50\delta}$.
By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$. Take any $b^{\prime}\in B^{\prime}$. It follows that there exists $(1,w)\in B^{\prime}b^{\prime-1}$, $w=q^{\prime}w_{0}$, where $q^{\prime}|q_{2},q^{\prime}\leq q_{1}^{3\delta}q_{2}^{2\delta}$ and $w_{0}$ is a primitive vector in $(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$.
It is an elementary exercise to check that for any primitive vector $v\in(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$, any $m\leq n$,
$\Gamma(p^{m})/\Gamma(p^{n})v-\Gamma(p^{m})/\Gamma(p^{n})v\supset(p^{m}\mathbb{Z}/p^{n}\mathbb{Z})^{2}.$ (3.22)
Let $L\subset A^{2882}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
$A^{5764}\supset LL^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
Multiplying the RHS of (3.21) and the RHS of (3.23) we obtain Proposition 3.20 with $q_{2}^{\prime}=\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})$.
Let $u,v\in\mathbb{Z}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
$f_{u,v}(A,B,C,D,E,F)=\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)-1\right)\left(\begin{matrix}u\\ v\end{matrix}\right)-\left(\begin{matrix}E\\ F\end{matrix}\right).$
It is an elementary exercise to check
$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg{|}f_{u,v}=0\right\}.$
is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
The following lemma shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$.
Proposition 3.24.
Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\Gamma(p^{m})/\Gamma(p^{n})|$. Then there exist some $u,v\in\frac{1}{p^{m}}\mathbb{Z}$ such that
$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})\subset H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})$
Proof.
Lemma 3.24 follows from proofs of a series of claims.
Claim 1: Suppose $(g,w)\in H$. If $p^{k}|g-1$, then $p^{k-m}|w$.
Proof of Claim 1. We prove by backward induction on $k$. When $k=n$, we must have $w=0$ by assumptions on $H$. Now let $k<n$ and assume the statement is true for $k+1$. Write $g=1+p^{k}M$. Since for $t\in\mathbb{N}$, $g^{t}\equiv 1+tp^{k}M\pmod{p^{\min\{2k,n\}}}$, we have
$(g,w)^{p}=(g^{p},(1+g+g^{2}+\cdots+g^{p-1})w)$
with $g^{p}\in\Gamma(p^{k+1})/\Gamma(p^{n})$ and
$(1+g+g^{2}+\cdots+g^{p-1})w\equiv(g^{p},p(1+\frac{(p-1)}{2}p^{k}M)w)(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
By induction hypothesis,
$(1+\frac{p-1}{2}p^{k}M)pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
Since $1+\frac{p-1}{2}p^{k}M$ is non-singular $\operatorname{mod}p$, we get
$pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}\Rightarrow w\in(p^{k-m}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
This proves Claim 1.
Claim 2: If
$A_{1}=(g_{1},w_{1})=\left(\left(\begin{matrix}1&p^{k}\\ 0&1\end{matrix}\right),\left(\begin{matrix}p^{k-m}a_{1}\\ p^{k-m}b_{1}\end{matrix}\right)\right)\in H$
and
$A_{2}=(g_{2},w_{2})=\left(\left(\begin{matrix}1&0\\ p^{k}&1\end{matrix}\right),\left(\begin{matrix}p^{k-m}a_{2}\\ p^{k-m}b_{2}\end{matrix}\right)\right)\in H,$
with $m\leq k\leq\frac{n}{4}$, then $a_{1},b_{2}\equiv 0\pmod{p^{k}}$ if $p\neq 3$, and $a_{1},b_{2}\equiv 0\pmod{p^{k-1}}$ if $p=3$.
Proof of Claim 2. We compute
$A_{3}=A_{1}A_{2}A_{1}^{-1}A_{2}^{-1}\equiv\left(\left(\begin{matrix}1+p^{2k}&0\\ 0&1-p^{2k}\end{matrix}\right),p^{2k-m}\left(\begin{matrix}b_{2}\\ -a_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{3k},p^{3k-m})$
and
$A_{4}=A_{3}A_{1}A_{3}^{-1}A_{1}^{-1}\equiv\left(\left(\begin{matrix}1&2p^{3k}\\ 0&1\end{matrix}\right),p^{3k-m}\left(\begin{matrix}2a_{1}\\ -b_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{4k},p^{4k-m}).$
On the other hand,
$A_{5}=A_{1}^{2p^{2k}}\equiv\left(\left(\begin{matrix}1&2p^{3k}\\ 0&1\end{matrix}\right),p^{3k-m}\left(\begin{matrix}2a_{1}\\ 2b_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{4k},p^{4k-m}).$
Applying Claim 1 to $A_{4}A_{5}^{-1}$, we obtain $3b_{1}\equiv 0(\operatorname{mod}p^{k})$. Similarly, $3a_{2}\equiv 0(\operatorname{mod}p^{k})$. Claim 2 thus follows.
Claim 3: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
Proof of Claim 3. consider
$A_{6}=A_{1}^{p^{t}}=\left(\left(\begin{matrix}1&p^{m+t}\\ 0&1\end{matrix}\right),\left(\begin{matrix}\cdots\\ p^{t}v_{1}\end{matrix}\right)\right)\in H.$
If taking $t=\lfloor\frac{n}{4}\rfloor-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}.$ This finishes the proof of Claim 3.
By the same consideration, we obtain
Claim 4: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$ since
$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor}),$
and that $\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right)$ and $\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right)$ generate $\Gamma(p^{2m})/\Gamma(p^{n})$. ∎
4. Preliminaries on Random Walks
Recall that $S$ be a finitely symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$. The following lemmas are quantitative statements about non-concentration of self-convolutions of $\chi_{S}$ in proper subvarieties.
For the discussion in the following in this section, we prefix a primitive linear form
$L\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right),\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)=X_{1}x_{1}+Y_{1}y_{1}+Z_{1}z_{1}+W_{1}w_{1}+X_{2}x_{2}+Y_{2}y_{2}+Z_{2}z_{2}+W_{2}w_{2},$
i.e., $\text{gcd}(X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2})=1$.
Proposition 4.1.
Let $S$ be a finite symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c}.$
Proposition 4.2.
There are constants $10<c_{S},c<$ 1 depending only on $S$ such that the following holds. Let $Q\in\mathbb{Z}_{+}$ large enough and $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0$
and either for every $p|Q$, $\pi_{p}(\xi_{1}),\pi_{p}(\xi_{2})\neq 0$, or for every $p|Q$, $\pi_{p}(\eta_{1}),\pi_{p}(\eta_{2})\neq 0$. Then for $l>c_{S}\log Q$ and any $n\in\mathbb{Z}$,
$\chi_{S}^{(l)}(\{(g_{1},g_{2})\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|\text{Tr}(g_{1}\xi_{1}g_{1}^{-1}\eta_{1})+\text{Tr}(g_{2}\xi_{2}g_{2}^{-1}\eta_{2})\equiv n(\operatorname{mod}{Q})\})<Q^{-c}.$
In the $\mathrm{ASL}_{2}$ case, we also have the following analogue about non-concentration in proper subvarieties.
Proposition 4.4.
Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense in $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c},$
where $L$ is a primitive linear form on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
We focus on proving Proposition 4.1; Proposition 4.2 and Proposition 4.4 follow in a similar way (see also Proposition 4.1 in [BG09]). We first show that Proposition 4.1 follow from
Lemma 4.5.
There are constants $10<c_{S},c_{0}<$ 1 depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for $l\in\mathbb{Z}_{+}$ with $1\ll_{S}l<c_{S}\log Q$, we have
$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n\pmod{{Q}}\})<e^{-c_{0}l}.$ (4.6)
Proof of Proposition 4.1 assuming Lemma 4.5.
Let $c_{S},c_{0}$ be the constants given by Lemma 4.5.
Let $l_{0}=[c_{S}\log Q]$ and write $\chi_{S}^{(l)}=\chi_{S}^{l_{0}}*\chi_{S}^{(l-l_{0})}$. For any $g^{\prime}$ in the support of $\chi_{S}^{l-l_{0}}$, let $L_{g^{\prime}}(g)=L(gg^{\prime})$. Clearly, $L_{g^{\prime}}$ is also primitive, so Lemma 4.5 is applicable to $L_{g^{\prime}}$. Therefore,
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)$
$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{(g^{\prime})^{-1}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
$<$ $e^{-c_{0}l_{0}}<Q^{-c_{0}c_{S}}$
Proposition 4.1 is thus proved with $c=c_{0}c_{S}$
Now we focus on proving Lemma 4.5.
Lemma 4.7 (Non-concentration at Archimedean place).
There is a constant $c_{0}>0$ depending only on $S$ such that for any $l\gg_{S}1$, any $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z}):L(g)=n\right\}\right)<e^{-c_{0}l}.$
Proof.
By Theorem A , there exists an absolute constant $10<\lambda<$ 1 which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1). So
$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}$ (4.8)
for all prime $p$. Then since $|\pi_{p}(\Gamma)|\approx p^{6}$, if $l>\frac{6\log p}{(\log 1/\lambda)}$,
$\chi_{S}^{(l)}(g)<\dfrac{2}{|\pi_{p}(\Gamma)|}$ (4.9)
for any $g\in\pi_{p}(\Gamma)$. Therefore,
$\chi_{S}^{(l)}\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}):L(g)=n\}$
$\leq\chi_{S}^{(l)}\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}):L(g)\equiv n(\operatorname{mod}{p})\}$
$<\frac{4}{p},$
by counting points in the subvariety of codimension $1$ determined by a primitive linear equation. So we are done by picking any prime $p\in[e^{\frac{l}{8}\log\frac{1}{\lambda}},e^{\frac{l}{7}\log\frac{1}{\lambda}}]$, which exists by the Prime Number Theorem when $l$ is sufficiently large. This finishes the proof of Lemma 4.7. ∎
We follow the method from Bourgain and Gamburd [BG09] to prove Lemma 4.5. The proof requires Effective Bézout Identity [BY91]. For readers’ convenience, we record it here:
Theorem C (The Effective Bézout Identity).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ without common zeros in $\mathbb{C}^{n}$ with $n\geq 2,\operatorname{deg}\mathcal{P}_{j}\leqslant d,d\geqslant 3$, $h\left(\mathcal{P}_{j}\right)\leqslant h$. Here $h(\mathcal{P})=\log\max\limits_{i}\left|a_{i}\right|$ for $\mathcal{P}(x)=a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{n}x^{n}$ is the logarithm height function of polynomials. Then there is an integer $D\in\mathbb{Z}_{+}$, polynomials $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ such that
$\begin{gathered}\mathcal{P}_{1}\mathcal{Q}_{1}+\ldots+\mathcal{P}_{N}\mathcal{Q}_{N}=D,\\ \operatorname{deg}\mathcal{Q}_{j}\leq n(2n+1)d^{n},\\ h\left(\mathcal{Q}_{j}\right)\leqslant\mathfrak{X}(n)d^{8n+3}(h+\log N+d\log d),\\ \log D\leqslant\mathfrak{X}(n)d^{8n+3}(h+\log N+d\log d),\end{gathered}$
where $\mathfrak{X}(n)$ is an effective constant which only depends on $n$ and can be computed explicitly.
Proof of Lemma 4.5.
Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$. We first deal with the case $n=0$.
By the primitivity of $L$, for each $p|Q$, at least one of
$t\in\mathcal{A}:=\{X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2}\}$
must be invertible $\operatorname{mod}\ p$. For each $t\in\mathcal{A}$, let
$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}||Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}$
Since $\prod_{t\in\mathcal{A}}Q_{t}\geq Q$, there exists $t\in\mathcal{A}$ such that
$Q^{\prime}:=Q_{t}\geq Q^{\frac{1}{8}}.$
Assume $t=X_{1}$ without loss of generality. Then $Q^{\prime}||Q$ and $(Q^{\prime},X_{1})=1$.
Now let
$\|(g_{1},g_{2})\|=\max\{\text{absolute values of coefficients of }g_{1}\text{ and }g_{2}\}.$
Let $C_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$. Define
$\mathcal{G}=\left\{g\in\text{supp }[\chi_{S}^{(l)}]\mid L(g)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\right\}.$
[…]
$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl}.$
for some $c>0$. For each $\gamma=\left(\begin{pmatrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{pmatrix},\begin{pmatrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{pmatrix}\right)\in\mathcal{G}$ we introduce a linear polynomial $f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Q}[\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}]$ as follows:
$f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+{\tilde{Y}}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}$ (4.10)
Then we get
$X_{1}f_{\gamma}\left(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\text{ for all }\gamma\in\mathcal{G}.$ (4.11)
Here $\bar{X_{1}}$ is the multiplicative inverse of $X_{1}$ mod $Q^{\prime}$.
Also, by definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely entries of $\gamma$, are bounded by $(2C_{1})^{l}$. Hence,
$h\left(f_{g}\right)<2l\log C_{1},$
We claim that there is a common zero $(\tilde{Y},\tilde{Z},\tilde{W},\tilde{A},\tilde{B},\tilde{C},\tilde{D})\in\mathbb{C}^{7}$ to the following system of equations:
$f_{\gamma}(Y,Z,W,A,B,C,D)=0\quad\text{ for all }\gamma\in\mathcal{G},$ (4.12)
[…]
Assume the claim fails to hold. We invoke Theorem C with $n=7,d=3,h=2l\log C_{1}$. It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{l}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
$M=\sum_{l=1}^{N}F_{l}\varphi_{l}$ (4.13)
with
$0<\log M,h\left(\varphi_{l}\right)<\mathfrak{X}(7)3^{59}((\log C_{1}+\log(2|S|))l+3\log 3)<C_{2}l$ (4.14)
where $C_{2}=3^{61}\mathfrak{X}(7)\log 2C_{1}|S|$.
Now we take $(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\in\mathbb{Z}^{7}$ such that
$(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\equiv(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}})(\operatorname{mod}Q^{\prime}).$
It follows that
$M\equiv\sum_{l=1}^{N}F_{l}\varphi_{l}\Bigg{|}_{(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})}\equiv 0\ (\operatorname{mod}{Q^{\prime}}),$
[…]
$x_{1}+\tilde{Y}y_{1}+\tilde{Z}z_{1}+\tilde{W}w_{1}+\tilde{A}x_{2}+\tilde{B}y_{2}+\tilde{C}z_{2}+\tilde{D}w_{2}=0$
Rescaling to get rid of the common denominator, we obtain $(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Z}^{8}$ such that $\text{gcd}(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=1$, and for all $\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right)\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)\in\mathcal{G}$, we have
$\tilde{X}_{1}x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+\tilde{Y}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}=0.$
We finish the proof of Lemma 4.5 of the case $n=0$ by applying Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}.$ For $1\ll_{S}l<c_{S}\log Q$,
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid L(g)\equiv 0(\operatorname{mod}{Q})\right\}\right)$
[…]
=\ = $\chi_{S}^{(l)}(\mathcal{G})$
[…]
<\ < $e^{-c_{0}l}.$
where $c_{0}$ is the constant given in Lemma 4.7.
Now let us deal with the case $n\neq 0$. Let $C_{3}=3^{70}\mathfrak{X}(8)$, where $\mathfrak{X}(8)$ is the constant given by Theorem C, and take $c_{S}=\frac{1}{C_{3}}$. Write $Q_{1}=(Q,n)$.
Suppose $Q_{1}<Q^{\frac{1}{C_{3}}}$. First we can find $n^{\prime}=Q_{1}m\equiv 0(\operatorname{mod}Q)$ with $(m,Q)=1$. The equation $L(\gamma)=n(\operatorname{mod}Q)$ is the same as $L(\gamma)=Q_{1}m(\operatorname{mod}Q)$. Since $m$ is invertible mod $Q$, we can define
$f_{\gamma}(X,Y,Z,W,A,B,C,D)=xX+yY+zZ+wW+aA+bB+cD+dD-Q_{1}$
for $\gamma=\left(\left(\begin{matrix}x&y\\ z&w\end{matrix}\right),\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)$, and proceed as the $n=0$ case. The reason in the $n=0$ case we do not define $f_{\gamma}$ in this way is that we need to ensure a solution to $f_{\gamma}=0$ is nonzero. But if $n\neq 0$, $f_{\gamma}=0$ never admits a zero solution.
If $Q_{1}\geq Q^{\frac{1}{C_{3}}}$, then we can reduce this case to the $n=0$ case by considering the equation $L(g)\equiv 0(\operatorname{mod}Q_{1})$ instead of $L(g)\equiv n(\operatorname{mod}Q)$, and letting $c_{S}=\frac{1}{8C_{2}C_{3}}$. ∎
5. Bounded generation for simple factors
Let $\mathbb{P}_{0}$ be the projection from $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ to $\text{SL}_{2}(\mathbb{Z})$, and let $\mathbb{P}_{1}$ ($\mathbb{P}_{2}$, resp.) be the projection map from $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ to its first factor (second factor, resp.). The goal of this section is to prove, the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
Proposition 5.1.
Let $S,\epsilon$ be as given in Proposition 2.2. There are constants $c>0$ depending only on the generating set $S$ and $\epsilon$, and a constant $\rho=\rho(\delta)>0$, $C=C(\delta)\in\mathbb{Z}_{+}$, with $\rho=O_{S}(\delta)$, such that for any $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}>\frac{12}{\delta}$, if the conclusion of Proposition 2.2 fails, i.e.,
$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta},$ (5.2)
then we have
$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &\mathbb{(}P_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
where $q_{*}||q$ and $q_{*}\geq q^{c}$.
We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$. The proofs for the other two cases are identical. Our proof mostly follows the steps of [BG09].
5.1. A large set of commuting elements
We use Helfgott’s argument [Hel08] to create a large set of commuting elements under an appropriate modulus.
Let $c_{1},c_{2}$ be the implied constants for $c$ from Proposition 4.1 and Proposition 4.2 respectively. Let $A_{0}=A\cdot A\cap\Lambda(q_{0})$, where $q_{0}=\prod_{p|q_{l}}p$. Then
$\chi_{S}^{2l}(A_{0})>q^{-2\delta}q_{0}^{-12}>q^{-3\delta},$ (5.4)
where we have used the exponent for each prime divisor of $q$ exceeds $\frac{12}{\delta}$.
Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$ and let
$A_{1}=\mathbb{P}_{1}(A_{0}).$
For each $Q|q$, let
$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)^{2}-4\equiv 0(\operatorname{mod}Q))\},$
$\mathcal{E}_{2}(Q)=\left\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\gamma_{21}\equiv 0(\operatorname{mod}Q))\right\}.$
Then Proposition 4.1 gives
$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}}.$
Take $\rho_{1}>0$, and let
$A_{2}=A_{1}-\cup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\{\rho_{1}\}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)),$
We should think of $\rho_{1}$ as a relative level depending only on $S$, so as $\rho_{2}$ in the following; the exact value of $\rho_{1}$ is determined at (5.58). Since the number of divisors of $q$ is bounded by $q^{0+}$, it follows that
$\nu^{(2l)}(A_{2})>q^{-3\delta-},$
if we let
$\delta<\frac{\rho_{1}c_{1}}{4}.$ (5.5)
A simple argument by contradiction shows that for each $\gamma\in A_{2}$, there exists $q_{\gamma}||q$, $q_{\gamma}>q^{\frac{1}{2}}$, such that for any $p^{n}||q_{\gamma}$,
$\text{tr}(\gamma)^{2}-4\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.6)
$\gamma_{21}\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.7)
Since again the number of divisors of $q$ is bounded by $q^{0+}$, there exists $q_{1}||q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{3}\subset A_{2}$, with
$\nu^{(2l)}(A_{3})>q^{-3\delta-},$ (5.8)
[…]
Let
$Q_{1}=q_{1}^{\{\rho_{2}\}},$
where we will take $\rho_{2}\gg\rho_{1}$.
Let
$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}Q_{1})\}.$
By Proposition 4.1, for any $n\in\mathbb{Z}$,
$\nu^{2l}(W(n))<Q_{1}^{-c_{1}}.$ (5.9)
If we take
${\delta}<{\frac{\rho_{2}c_{1}}{12}},$ (5.10)
then $\nu^{(2l)}(A_{3})>Q_{1}^{-\frac{c_{1}}{2}}$, and together with (5.9), it implies that there exists a set $A_{4}\subset A_{3}$, such that
$|A_{4}(\operatorname{mod}Q_{1})|>Q_{1}^{c_{1}/2}$ (5.11)
$\forall\gamma_{1},\gamma_{2}\in A_{4},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\neq\text{tr}(\gamma_{2})(\text{mod }Q_{1})$ (5.12)
Next we need to use our assumption (5.2) to get a control over the growth of $\pi_{Q_{1}}(A_{1})$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
$|\pi_{q}(\prod_{l}A)|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.13)
$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.14)
It follows from Proposition 4.1 that for any $\gamma_{0}\in\Lambda_{q}$,
$\nu^{(2l)}\{\gamma\in\Lambda:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{1})\}<Q_{1}^{-c_{1}}.$ (5.15)
From (5.15) and (5.10), we conclude that
$|\pi_{Q_{1}}(A_{1})|>Q_{1}^{c_{1}}q^{-3\delta}>Q_{1}^{\frac{c_{1}}{2}}.$
Next we show small expansion for $A$ implies small expansion for $A_{1}$:
$|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$ (5.16)
To see this, suppose not, then
$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{\frac{4\delta}{\rho_{2}c_{1}}}\cdot|\pi_{q}(A)|$
$>$ $|\pi_{q}(A)|^{1+\delta}$ (5.17)
which contradicts (5.2).
For each $\alpha\in A_{4}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in A_{1}\}$. Since distinct elements in $A_{4}$ have distinct eigenvalues mod $Q_{1}$, the sets $\{\pi_{Q_{1}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint. Therefore,
$\sum_{\alpha\in A_{4}}|\pi_{Q_{1}}(C_{\alpha})|=|\cup_{\alpha\in A_{4}}\pi_{Q_{1}}(C_{\alpha})|<|\pi_{Q_{1}}(A_{1}^{3})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$
Thus, for some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{4}$,
$|\pi_{Q_{1}}(C_{\alpha_{0}})|\leq|\pi_{Q_{1}}(A)|^{1+\frac{4\delta}{\rho_{2}c_{1}}}Q_{1}^{-c_{1}/2}<|\pi_{Q_{1}}(A)|Q_{1}^{-\frac{c_{1}}{4}}.$ (5.18)
provided that
$\delta<\frac{\rho_{2}c_{1}^{2}}{48}.$ (5.19)
To proceed, we first diagonalise $\alpha_{0}$ in an extended number field, i.e., let
$\lambda_{1}=\frac{\text{tr}(\alpha_{0})+\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2},\hskip 14.22636pt\lambda_{2}=\frac{\text{tr}(\alpha_{0})-\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2}$
be the two eigenvalues of $\alpha_{0}$. Let $K=\mathbb{Q}[\lambda_{1}]$, and so $\mathcal{O}=\mathcal{O}_{K}=\mathbb{Z}[\lambda_{1}]$. Let $M=\left(\begin{matrix}\lambda_{1}-\alpha_{22}&\lambda_{2}-\alpha_{22}\\ \alpha_{21}&\alpha_{21}\end{matrix}\right)$. Then
$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$
Since $\alpha_{0}$ satisfies the condition (5.6) for $\gamma$,
$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]})$ (5.20)
for any $p^{n}||q_{1}$.
From (5.18), there exists $x_{0}\in A_{4}$ and
$A_{5}:=\{x\in A:x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{1})\},$ (5.21)
such that
$|\pi_{Q_{1}}(A_{5})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.22)
Therefore, $A_{5}^{-1}\cdot x_{0}$ commute with $\alpha_{0}(\operatorname{mod}Q_{1})$ and
$|\pi_{Q_{1}}(A_{5}^{-1}\cdot x_{0})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.23)
Let $C_{1},C_{2},C_{3}$ be the implied constants of Theorem B for $\delta=\frac{c_{1}}{17}$.
The condition (5.23) implies that there exists $\upsilon\in[\frac{c_{1}}{24},\frac{1}{2}]$, such that
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{5})|>|\pi_{Q_{1}^{\{\upsilon\}}}(A_{5})|Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.24)
If we let $A_{6}=\left((A_{5}^{-1}\cdot x_{0})\cdot(A_{5}^{-1}\cdot x_{0})^{-1}\right)\cap\Gamma(Q_{1}^{\{\upsilon\}})=A_{5}^{-1}A_{5}\cap\Gamma(Q_{1}^{\{\upsilon\}})$, then
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{6})|>Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.25)
Let
$Q_{2}:=q_{1}^{\{\rho_{2}\upsilon-9\rho_{1}\}},$
where we will take $\rho_{2}\upsilon\gg 9\rho_{1}$. Since $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from (5.7), (5.6), we have
$p^{[6\rho_{1}n]}\nmid\text{Det}(M),\forall p^{n}||q_{1}.$ (5.26)
Combining (5.26) with (5.20), we conclude that the matrix $M$ conjugate every matrix $\sigma$ in $A_{6}$ to a matrix in $\text{SL}_{2}(\mathcal{O})$ which is congruent to identity mod $Q_{2}$ and diagonal mod $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$. In other words, there exists $a_{\sigma}\in\mathcal{O}_{K}$ and $a_{\sigma}\equiv 1(\operatorname{mod}Q_{2})$ such that
$M^{-1}\sigma M\equiv\left(\begin{matrix}a_{\sigma}&0\\ 0&a_{\sigma}^{-1}\end{matrix}\right)(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}).$ (5.27)
It also follows from (5.25) that
$\Big{|}\pi_{Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}(A_{6})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}}.$ (5.28)
if
$\rho_{1}<\frac{\upsilon c_{1}}{2^{5}3^{3}}\rho_{2},$ (5.29)
which is guaranteed if
$\rho_{1}<\frac{c_{1}^{2}}{2^{8}3^{4}}\rho_{2}.$ (5.30)
For each $\gamma\in A_{6}$, let ${q}_{\gamma}$ be the largest divisor of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ such that $\gamma\equiv 1(\operatorname{mod}q_{\gamma})$. Obviously, $Q_{2}|q_{\gamma}$. Then we can write
$A_{6}=\sqcup_{Q_{2}|q^{\prime}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}B_{q^{\prime}},$
where $B_{q^{\prime}}=\{\gamma\in A_{6}:q_{\gamma}=q^{\prime}\}$. Since the number of divisors of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ is bounded by $q^{0+}$, there exists $\tilde{q}$ such that $Q_{2}|\tilde{q}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ and
$\Big{|}B_{\tilde{q}}(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.31)
5.2. Applying sum-product
Let $H=\{h_{\sigma}=(a_{\sigma}-a_{\sigma}^{-1})/\tilde{q}:\sigma\in B_{\tilde{q}}\}\subset\mathcal{O}$, where $a_{\sigma}$ is as given in (5.27). Since the map $a_{\sigma}\mapsto a_{\sigma}-a_{\sigma}^{-1}$ is bounded to one, we have
$\Big{|}H(\operatorname{mod}{Q_{2}^{\{\frac{1}{C_{1}}\}}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.32)
Recall $\mathcal{O}$ is a degree 2 extension of $\mathbb{Z}$. We apply Theorem B to the set $H$ with $\mathfrak{a}=(Q_{2}^{\{\frac{1}{C_{1}}\}})$ and $\delta=\frac{c_{1}}{17}$. Theorem B then produces some $\xi\in\mathcal{O}$, $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$, $c_{3}=c_{3}(c_{1})>0$ such that
$\sum_{C_{3}}H^{C_{2}}\supset\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime}),$ (5.33)
and
$|\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime})|>N(\mathfrak{a})^{c_{3}}>q_{1}^{\frac{\rho_{2}\upsilon c_{3}}{C_{1}}}>q_{1}^{\frac{\rho_{2}c_{1}c_{3}}{24C_{1}}}.$ (5.34)
We pass (5.34) to a local property. From (5.34), it follows that there exists $q_{2}||q_{1}$, $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}$, such that for any $p^{n}||q_{2}$, letting $\overline{\mathfrak{a}_{p}}=\langle\mathfrak{a}^{\prime},p^{n}\rangle$, the ideal of $\mathcal{O}$ generated by $\mathfrak{a}^{\prime}$ and $p^{n}$, we have
$|\mathbb{Z}\xi(\operatorname{mod}\overline{\mathfrak{a}_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{48C_{1}}}.$
Since there are at most two prime ideals in $\mathcal{O}$ lying above $p$, there exists a prime ideal $\mathcal{P}_{p}\supset(p)$ in $\mathcal{O}$, $\mathcal{P}_{p}^{m_{p}}||\bar{\mathfrak{a}_{p}}$, such that
$|\mathbb{Z}\xi(\operatorname{mod}\mathcal{P}_{p}^{m_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{96C_{1}}}.$ (5.35)
For simplicity, we assume for each $p|q_{3}$, the arising prime ideal $\mathcal{P}_{p}$ is unramified, so $v_{p}(\tilde{q})=v_{\mathcal{P}}(\tilde{q})$. The general case adds no additional technical challenge.
We obtain from (5.35), $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$ and the definition of $\mathfrak{a}$ that
$\frac{\rho_{2}c_{1}c_{3}}{96C_{1}}n<m_{p}<\rho_{2}n.$ (5.36)
It also follows from (5.35) that for any $p|q_{2}$,
$v_{p}(Q_{2})\leq v_{p}(\tilde{q})<v_{p}(Q_{2}^{\{1+\frac{1}{C_{1}}\}}).$ (5.37)
5.3. Constructing an appropriate element $\zeta$
We also need to find an element $\zeta=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ from $A_{1}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$. This leads us to work with the linear form
$\mathcal{L}_{1}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$
which gives the 2-1 entry of $\alpha_{0}\zeta-\zeta\alpha_{0}$. If $\mathcal{L}_{1}(\zeta)$ is non-zero, this certainly implies $\zeta$ and $\alpha_{0}$ are not commutative.
Let
$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}_{1}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
By Proposition 4.1, for any $Q|q_{2}$ ,
$\nu^{(2l)}(\mathcal{E}_{3}(Q))<Q^{-c_{1}},$ (5.38)
Define
$A_{7}=A_{1}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
By (5.4), $\nu^{(2l)}(A_{1})>q^{-3\delta}$. We also have $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}>q^{\frac{c_{3}}{4C_{1}}}$. Then
$\nu^{(2l)}(A_{7})>q^{-4\delta}$
if $q_{2}^{-\rho_{1}c_{1}}<q^{-4\delta}$, or
${\delta}<\frac{c_{1}c_{3}\rho_{1}}{16C_{1}}.$ (5.39)
Then we produce some $q_{3}||q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{1}$ such that for any $p^{n}||q_{3}$,
$L({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}).$ (5.40)
From (5.40) and (5.20), we have $M\zeta_{1}M^{-1}=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$ with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}p^{[6n\rho_{1}]})$. Take any element $\zeta_{2}\in B_{\tilde{q}}$. The second inequality of (5.37) ensures that $v_{p}(\zeta_{2})=v_{p}(\tilde{q})$.
If we let $\zeta=\zeta_{1}\zeta_{2}\zeta_{1}^{-1}$ and write $M{\zeta}M^{-1}=\left(\begin{matrix}\zeta_{11},\zeta_{12}\\ \zeta_{21},\zeta_{22}\end{matrix}\right)$. Then
$v_{p}(M\zeta M^{-1}-1)\geq v_{p}(\tilde{q})-[6\rho_{1}n]$ (5.41)
and
$\min\{v_{p}(\zeta_{12}),v_{p}(\zeta_{21})\}\leq v_{p}(\tilde{q})+[10\rho_{1}n].$ (5.42)
5.4. Creating a line
Let
$\left\{\vec{e}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right),\vec{f}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\vec{h}=\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right)\right\}$
[…]
To proceed, we recall the following well known identity.
Lemma 5.43.
Let $m,m^{\prime}\in\mathbb{Z}_{+}$, and let $x,y\in\text{SL}_{2}(\mathcal{O})$, $x\equiv 1(\operatorname{mod}\mathcal{P}^{m}),y\equiv 1(\operatorname{mod}\mathcal{P}^{m^{\prime}})$. Then
$xyx^{-1}y^{-1}\equiv 1(\operatorname{mod}\mathcal{P}^{m+m^{\prime}}),$
and
$xyx^{-1}y^{-1}\equiv 1+xy-yx(\operatorname{mod}\mathcal{P}^{m+m^{\prime}+\min\{m,m^{\prime}\}})$
Lemma 5.43 has an obvious generalization to a general ideal by multiplicativity.
To apply Lemma 5.43, we take $x=M^{-1}\sigma M$ given at (5.27) and $y=M^{-1}\zeta M$, with corresponding moduli $\mathfrak{q}_{1}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})}$ and $\mathfrak{q}_{2}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})-[6\rho_{1}n]}$.
Then
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2})$ (5.44)
and
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1+\tilde{q}h_{\sigma}\left(\zeta_{21}\vec{e}-\zeta_{12}\vec{f}\right)(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2}^{2})$ (5.45)
Applying Lemma 5.43 iteratively, for any $\sigma_{1},\sigma_{2},\cdots,\sigma_{C_{3}}\in A_{6}$, we have
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}-1}\mathfrak{q}_{2}^{2})$ (5.46)
and
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1+\tilde{q}^{C_{3}}\left(\prod_{i=1}^{C_{3}}h_{\sigma_{i}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.47)
where $X=\left(\begin{matrix}0&\zeta_{12}\\ (-1)^{C_{3}}\zeta_{21}&0\end{matrix}\right)$.
The sum structure is clear. For any $\sigma_{1,1},\cdots,\sigma_{C_{2},C_{3}}\in A_{6}$, we have
$(M^{-1}[\sigma_{1,C_{3}},\cdots[\sigma_{1,2},[\sigma_{1,1},\zeta]]M)\cdots(M^{-1}[\sigma_{C_{2},C_{3}},\cdots[\sigma_{C_{2,2}},[\sigma_{C_{{2},1}},\zeta]]\cdots]M)$
$\equiv$ $1+\tilde{q}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.48)
From (5.33) and (5.4), we conclude there is a constant $K_{1}$ depending only on $c_{1}$, such that
$1+\tilde{q}^{C_{3}}\mathbb{Z}\xi X\subset M^{-1}A_{1}^{K_{1}}M(\operatorname{mod}\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}\langle\mathfrak{a}^{\prime},\mathfrak{q}_{2}\rangle)$ (5.49)
Conjugating (5.49) back by $M$, noticing that $A_{1}^{K_{1}}$ is real and also having (5.35), (5.42), and (5.49) in mind, if we take
$\rho_{1}<\frac{c_{1}c_{3}}{28*192C_{1}}\rho_{2},$ (5.50)
then we obtain
Lemma 5.51.
There exists $q_{3}||q_{1}$, $q_{3}>q^{\frac{c_{1}}{8C_{1}}}$, $q_{3}^{\prime}|q_{3}^{\prime\prime}|q_{3}$, some $K_{1}=K_{1}(c_{1})>0$, a matrix $X_{0}\in{{\text{Lie(SL}}}_{2})(\mathbb{Z})$ coprime to $q_{3}$, such that
$1+q_{3}^{\prime}\mathbb{Z}X_{0}\subset A_{1}^{K_{1}}(\operatorname{mod}q_{3}^{\prime\prime})$ (5.52)
where for each $p^{n}||q_{3}$, $v_{p}(q_{3}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and $\frac{c_{1}c_{3}}{192C_{1}}\rho_{2}n<v_{p}(q_{3}^{\prime\prime}/q_{3}^{\prime})<{v_{p}(q_{3}^{\prime\prime})}/2$.
5.5. Creating a segment
Next, we want to find $g_{1},g_{2}$ from a bounded product of $A_{1}$ so that $g_{1},g_{2}$ conjugates $X_{0}$ to different directions.
[…]
If $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that $\text{Tr}(g\xi_{1}g^{-1}\eta_{1})=L_{1}(g)$. If $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, apply Proposition 4.4 with $L=L_{1}$. Let $c_{2}$ be the implied constant from Propositions 4.2 and 4.4. Then we obtain, for any $Q>q_{3}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{1}\circ T_{1}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
where $l$ is as given in Proposition 2.2. This implies, arguing the same as for the construction of $A_{3}$, that there is $q_{4}||q_{3}$, $q_{4}>(q_{3})^{\frac{1}{2}}$, and $g_{1}\in A_{1}$ such that for any $p^{n}||q_{5}$, $L_{1}\circ T_{1}(g_{1}X_{0}g_{1}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if we take
$\delta<\frac{c_{1}c_{2}\rho_{1}}{48C_{1}}.$ (5.53)
Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then $X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5})$, where $a,b\in\mathbb{Z}$, and $b$ satisfies, for any $p^{n}||q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the span of $(0,1,0)$ and $(0,0,1)$.
Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{0})=(0,b,0)$. Let $L_{2}:\mathbb{Z}^{3}\mapsto\mathbb{Z}$ be the linear form getting the third component. Then $L_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$. Applying Proposition 4.2 again, for any $Q>q_{4}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{2}\circ T_{2}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
which then implies there is $q_{5}||q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{2}\in A^{l}$ such that for any $p^{n}||q_{6}$, $L_{2}\circ T_{2}(g_{2}X_{0}g_{2}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if
${\delta}<\frac{c_{1}c_{2}\rho_{1}}{96C_{1}}.$ (5.54)
Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.52) with its $g_{1}$ and $g_{2}$ conjugates, we obtain
Lemma 5.55.
There are $q_{5}||q$, $q_{5}>q^{\frac{c_{1}}{32C_{1}}}$, $q_{5}^{\prime}|q_{5}^{\prime\prime}|q_{5}$, and some constant $K_{2}>0$ depending only on $c_{1}$, such that
$1+q_{5}^{\prime}\mathbb{Z}V\subset A_{1}^{K_{2}}(\operatorname{mod}q_{5}^{\prime\prime}),$ (5.56)
where for each $p^{n}||q_{5}$, $v_{p}(q_{5}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and
$\frac{c_{1}c_{3}}{384C_{1}}\rho_{2}n<v_{p}(q_{5}^{\prime\prime}/q_{5})<v_{p}(q_{5}^{\prime\prime})/2$
if
$\rho_{1}<\frac{c_{1}c_{3}}{3\cdot 2^{9}C_{1}}\rho_{2}.$ (5.57)
5.6. Proof of Proposition 5.1
Collecting all the restrictions (5.5), (5.10), (5.19), (5.29), (5.30), (5.39), (5.50), (5.53), (5.54), (5.57) on $\rho_{1},\rho_{2},\delta$, we find we can take
$\rho_{1}=\frac{c_{1}^{2}\cdot c_{3}}{2^{8}3^{4}C_{1}}\rho_{2}$ (5.58)
$\delta=\frac{c_{1}c_{2}c_{3}}{96C_{1}}\rho_{1}$ (5.59)
We need the elementary lemma whose proof we leave as an exercise:
Lemma 5.60.
Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime. Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{2}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$. Then
$1+p^{m_{1}+n_{1}}V\subset(H_{1}H_{2})^{4}(\operatorname{mod}p^{m_{2}+n_{2}}).$
Proposition 5.1 is thus proved by iteratively applying Lemma 5.60 to (5.56), with $q_{*}=q_{5}$, $c=\frac{c_{1}}{32C_{1}}$ and $\rho=O_{c_{1}}(\rho_{2})$, $C=O_{S}(\log\frac{1}{\delta})$.
6. A gluing technique for the $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ case
In this section and the next, we prove Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. Suppose all the assumptions about $A$ hold but the conclusion (2.4) does not hold, then by taking $\delta$ sufficiently small, we will find a constant $K$ independent of $\delta$, such that $|\pi_{q}(A^{K})|>|\Lambda_{q}|^{1-\frac{\epsilon}{2}}$. But this will contradict Helfgott’s inequality (5.2) when $\delta$ is sufficiently small.
Proposition 5.1 so far only covers a large subset in one simple factor of the group and for some not too small modulus $q_{*}$. Apply Proposition 5.1 again with $q$ replaced by $q/q_{*}$, we will cover a large subset for some modulus $q_{*}^{\prime}$ coprime to $q_{*}$. This leads to considering the following question: Given a set $B\subset\Gamma_{q_{1}q_{2}},\text{gcd}(q_{1},q_{2})=1$, such that $\pi_{q_{1}}(B),\pi_{q_{2}}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, is a bounded power of $B$ covers a large subset of $\Gamma_{q_{1}q_{2}}$?
The following proposition is the main gluing tool.
Proposition 6.1.
Fix $0<\theta<10^{-12}$ and $10<\eta<$ 1, and suppose
$\delta<\frac{c_{2}\eta\theta^{1/2}}{2},$ (6.2)
where $\delta$ is given as in Proposition 2.2 and $c_{2}$ is the implied $c$ in Proposition 4.2. Take
$L>\frac{12}{\delta}.$ (6.3)
Let $q=\prod_{i\in I}p_{i}^{n_{i}}=q_{s}q_{l}$, where $q_{s}=\prod_{p_{i}|q,v_{p_{i}}\leq L}p_{i}^{n_{i}}$ and $q_{l}=\prod_{p_{i}|q,v_{p_{i}}>L}p_{i}^{n_{i}}$. Let $q_{1},q_{2}||q,q_{3}||q_{l}$, $\text{gcd}(q_{1},q_{3})=1$, and $q_{1}q_{2},q_{3}>q^{\eta}$. Suppose for some set $B\subset\Lambda$, we have
$|\pi_{q_{1},q_{2}}(B)|>(q_{1}q_{2})^{3-\theta},$ (6.4)
$|\pi_{q_{3},1}(B)|>q_{3}^{3-\theta}.$ (6.5)
Then there exists $q_{3}^{*}||q_{3}$, $q_{3}^{*}>q_{3}^{\frac{1}{4}10^{-4}}$, such that
$|\pi_{q_{1}q_{3}^{*},q_{2}}(B\cup A)^{O(\log 1/\theta)}|>(q_{1}q_{2}q_{3}^{*})^{3-O(\theta^{\frac{1}{4}})}.$ (6.6)
Remark 6.7.
The point of Proposition 6.1 is we have an increase of modulus from $q_{1}$ to $q_{1}q_{3}^{*}$. To see the necessity of getting help from $A$, for instance, let $q_{1}=1$, $q_{2}=q_{3}$ and $B=\{(\lambda,\lambda),\lambda\in\Gamma\}$. Then $B\cdot B=B$ so there is no hope to expand only from $B$.
Write
$\Gamma_{q_{1}q_{3}}\times\Gamma_{q_{2}}\cong(\Gamma_{q_{1}}\times\Gamma_{q_{2}})\times(\Gamma_{q_{3}}\times 1).$
Since $B$ satisfies (6.4) and (6.5), by Proposition 3.12, there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{3}^{\prime}|q_{3},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{40\theta},q_{3}^{\prime}<(q_{3})^{40\theta}$, such that
$\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})\subset\pi_{q_{1},q_{2}}(B^{2880}).$ (6.8)
$\Gamma(q_{3}^{\prime})/\Gamma(q_{3})\subset\pi_{q_{3},1}(B^{2880}).$ (6.9)
Write
$G=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})$
From (6.8), we can construct a section map
$\psi:G\rightarrow B^{2880},$
such that $\pi_{q_{1},q_{2}}(\psi(x))=x$.
Since $q_{3}^{\prime}<q_{3}^{40\theta}$, there is $q_{4}||q_{3},q_{4}>q_{3}^{1-\theta^{\frac{1}{2}}}$ such that for every $p^{n}||q_{4}$, we have $p^{[40\theta^{\frac{1}{2}}n]}\nmid q_{3}^{\prime}$.
Write $q_{4}=\prod_{j\in J}p_{j}^{n_{j}}$. For each $p_{j}^{n_{j}}||q_{4}$, we consider $\psi_{j}=\pi_{p_{j}^{[n_{j}\theta^{1/4}]}}\circ\mathbb{P}_{1}\circ\psi$.
Define
$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)\}$
According to Proposition 3.1, there are two scenarios:
  1. (1)
    $|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (6.10)
  2. (2)
    There is a subset $S_{j}\in G$, $|S_{j}|\geq\frac{99}{100}|G|$ such that $\psi_{j}=h_{j}$ over $S_{j}$ where $h_{j}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]})$.
Let $J=J_{1}\sqcup J_{2}$ where $J_{1},J_{2}$ is the collection of indices falling into Case (1) and Case (2), respectively. Write $q_{4}=q_{4}^{\prime}q_{4}^{\prime\prime}$, where
$q_{4}^{\prime}=\prod_{j\in J_{1}}p_{j}^{n_{j}},\hskip 28.45274ptq_{4}^{\prime\prime}=\prod_{j\in J_{2}}p_{j}^{n_{j}}.$
We further write $q_{4}^{\prime\prime}=q_{5}q_{5}^{\prime}$, where
$q_{5}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}=1\end{subarray}}p_{j}^{n_{j}},\hskip 28.45274ptq_{5}^{\prime}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}\neq 1\end{subarray}}p_{j}^{n_{j}}.$
We divide our analysis into three cases.
6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
In this case we have
$\sum_{j\in J_{1}}(\log p_{j}^{n_{j}})|\mathcal{G}_{j}|>\log\left((q_{4}^{\prime})^{10^{-4}}\right)|G|^{2}$ (6.11)
The left side of (6.11) is equal to
$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(\prod_{j\in U}p_{j}^{n_{j}})|\cap_{j\in U}\mathcal{G}_{j}\bigcap\cap_{j\in J_{1}-U}\mathcal{G}_{j}^{c}|$ (6.12)
Since the number of subsets of $J_{1}$ is $<q_{4}^{0+}$, there exists $J_{1}^{\prime}\subset J_{1}$, $\tilde{q}=\prod_{i\in J_{1}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime})^{\frac{1}{2}10^{-4}}$, such that
$|\cap_{i\in J_{1}}\mathcal{G}_{i}|>q^{0-}|G|^{2}.$
Take any $(g_{1},g_{2})\in\cap_{i\in J_{1}^{\prime}}\mathcal{G}_{i}$, and consider $\gamma_{0}=\psi(g_{1})\psi(g_{2})\psi(g_{1}g_{2})^{-1}$. Then $\gamma_{0}$ satisfies,
$\pi_{q_{1},q_{2}}(\gamma_{0})=1,$
and for any $p^{n}||\tilde{q}$,
$\pi_{p^{[n\theta^{{1}/{4}}]}}\circ\mathbb{P}_{1}(\gamma_{0})\neq 1.$
Take a set $B^{\prime}\subset B^{2880}$ which maps bijectively to $\Gamma(q_{3}^{\prime})/\Gamma(q_{3})$ under the projection $\pi_{q_{3},1}$. Take
$B^{\prime\prime}=\{b_{1}\gamma_{0}b_{1}^{-1}\gamma_{0}^{-1}\cdot(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})\gamma_{0}(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})^{-1}\gamma_{0}^{-1},b_{1},b_{2}\in B^{\prime}\}$
It follows from a quantitative open mapping result due to Salehi-Golsefidy (see Propostions 49, Proposition 49’ of [Gol19]) that
$\pi_{q_{3}}\circ\mathbb{P}_{1}\left(B^{\prime\prime}\right)\supset\Gamma(q_{3}^{\{O(\theta^{1/4})\}}).$ (6.13)
by first taking the logarithmic map to work on the Lie algebra level, then taking the exponential map back. The definition for $B^{\prime\prime}$ is motivated by the fact that for any $v\in\mathfrak{sl}_{2}(\mathbb{Q})$, $\mathfrak{sl}_{2}(\mathbb{Q})=[\mathfrak{sl}_{2}(\mathbb{Q}),v]+[[\mathfrak{sl}_{2}(\mathbb{Q}),v],v]$, where $+$ here means set addition.
Obviously, $\pi_{q_{1},q_{2}}(B^{\prime\prime})=1$. Multiplying $B^{\prime\prime}$ with $B^{2880}$ and observing (6.8) we obtain Proposition 6.1 in this case by setting $q_{3}^{*}=\tilde{q}$.
6.2. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
The local homomorphisms $h_{j},j\in J_{2}$ can be lifted to a homomorphism
$h:G\rightarrow\Gamma(\prod_{j\in J_{2}}p_{j})/\Gamma((q_{4}^{\prime\prime})^{\{\theta^{\frac{1}{4}}\}}).$
Following the previous reasoning for obtaining $J_{1}^{\prime}$, there is a set $J_{2}^{\prime}\subset J_{2}$, such that
$\bar{q}=\prod_{i\in J_{2}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime\prime})^{\frac{99}{200}},$
$|\cap_{i\in J_{2}}S_{i}|>(q_{1}q_{2})^{-\theta}|G|,$
and $\psi\equiv h$ on $S=\cap_{i\in J_{2}}S_{i}$.
By Proposition 3.12, we have $S^{2880}\supset G^{\prime}=\Gamma(q_{1}^{\prime\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime\prime})/\Gamma(q_{2})$ for some $q_{1}^{\prime\prime}||q_{1}$, $q_{2}^{\prime\prime}||q_{2}$, $q_{1}^{\prime\prime}q_{2}^{\prime\prime}<(q_{1}q_{2})^{80\theta}$. This implies the existence of a subgroup $G^{\prime\prime}$ of $G^{\prime}$ of the form
$G^{\prime\prime}=\Gamma(\tilde{q}_{1}(q_{1}^{*})^{\{\theta^{\frac{1}{2}}\}})/\Gamma(q_{1})\times\Gamma(\tilde{q}_{2}(q_{2}^{*})^{\{\theta^{{1}/{2}}\}})/\Gamma(q_{2}),$ (6.14)
where
$q_{1}=\tilde{q_{1}}q_{1}^{*},q_{2}=\tilde{q_{2}}q_{2}^{*},\tilde{q}_{1},q_{1}^{*}||q_{1},\tilde{q}_{2},q_{2}^{*}||q_{2},\tilde{q}_{1}\tilde{q}_{2}<(q_{1}q_{2})^{80\theta^{{1}/{2}}}.$ (6.15)
Then we can define a map $\tilde{\psi}:G^{\prime\prime}\rightarrow B^{2880^{2}}$ as, for any $x\in G^{\prime\prime}$, choose a word $x=s_{1}s_{2}\cdots s_{2880}$ and let
$\tilde{\psi}(x)=\psi(s_{1})\psi(s_{2})\cdots\psi(s_{2880}).$ (6.16)
Then we have $\pi_{(q_{4}^{\prime\prime})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=h(x)$, and thus $\pi_{(q_{5})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=1$.
Recall Lemma 5.43. By iteratively taking commutator of $\tilde{\psi}(G^{\prime\prime})$ $O(\log(1/\theta))$ times, we recover a group
$\Gamma((q_{1}^{*})^{\{{2\theta^{{1}/{4}}}\}}q_{5})/\Gamma(q_{1}^{*}q_{5})\times\Gamma((q_{2}^{*})^{\{{2\theta^{{1}/{4}}}\}})/\Gamma(q_{2}^{*})\subset\pi_{q_{1}^{*}q_{5},q_{2}^{*}}(B^{O(\log(1/\theta))}),$ (6.17)
which is identity when projected to $q_{5}$ and is large when projected to $(q_{1},q_{2}^{\prime})$.
Recall $q_{5}||q_{3}$ and $v_{p}(q_{3}^{\prime})<v_{p}(q_{3})$ for every $p^{n}||q_{5}$. It follows from (6.15), (6.17), and (6.9) that
$|\pi_{q_{1}q_{5},q_{2}}(B^{O(\log 1/\theta)})|>(q_{1}q_{2}q_{5})^{3-{O(\theta^{{1}/{4}})}}.$ (6.18)
In this case, we take $q_{3}^{*}=q_{5}$ and the proposition is proved.
6.3. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}^{\prime}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
We begin with a lemma.
Lemma 6.19.
Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$. Then $p_{1}=p_{2}$ and $p^{x}||\xi-1$ for some $x\leq m_{1}+n_{2}-m_{2}$.
Proof.
We first observe that given $0\leq m_{1}<n_{1},1\leq m_{2}<n_{2}$ and two primes $p_{1},p_{2}$, if there is a nontrivial homomorphism $f$ from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{1}^{n_{2}})$, this will force $p_{1}=p_{2}$. To see this, if $m_{1}\neq 0$, then all elements in $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ have orders powers of $p_{1}$, and all elements in $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ have orders powers of $p_{2}$. Since $f$ is nontrivial, we have $p_{1}=p_{2}$.
If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$ and we obtaine a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$. Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor thorough the center $Z$. $f^{\prime}$ can not be injective because there are elements in $\Gamma_{p_{1}}$ of order $p_{1}$, which is coprime to $p_{2}$. Likewise, $Z$ can not be the kernel of $f^{\prime}$ because one can also find an element $x\in\Gamma_{p_{1}}/Z$ such that the order of $x$ is coprime to $p_{2}$. Hence, $f^{\prime}$ as well as $f$ must be trivial and we arrive at a contradiction, so $p_{1}=p_{2}$.
Suppose $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$. Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $x\leq m_{1}+n_{2}-m_{2}$. ∎
Lemma 6.19 implies that $q_{5}^{\prime}||q_{2}^{*}$, since $(q_{5}^{\prime},q_{1}^{*})=1$.
Let $G^{\prime\prime}$ and $\widehat{\psi}$ be the same as in Case 6.2.
The following isomorphism is clear
$G^{\prime\prime}\cong\Big{(}\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\Big{)}\times\Big{(}\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})\Big{)}\times\Big{(}\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\Big{)}$
Since $\pi_{(q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}}\circ\mathbb{P}_{1}\circ\tilde{\psi}$ is a homomorphism and factors through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, it induces a homomorphism
$\widehat{\psi}:\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\rightarrow\Gamma/\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}).$
Since $\widehat{\psi}$ is nontrivial after reduction by each exact divisor of $(q_{5}^{\prime})^{\left\{\frac{1}{2}\theta^{1/4}\right\}}$, we can find $g_{1}\in\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}})/\Gamma(q_{5}^{\prime})$ such that for each $p|q_{5}^{\prime}$,
$v_{p}(g_{1})=O(n\theta^{1/4}),\hskip 14.22636ptv_{p}(\widehat{\psi}(g_{1}))<(\frac{1}{2}n\theta^{1/4}).$
Let $g_{2}=g_{1}^{p^{\{\frac{1}{4}n\theta^{1/4}\}}}$, then
$v_{p}(g_{2})=\Theta(n\theta^{1/4}),\hskip 14.22636pt\frac{1}{4}n\theta^{1/4}<v_{p}(\widehat{\psi}(g_{2}))<\frac{3}{4}n\theta^{1/4}.$
Write $g_{2}\equiv 1+Q_{1}X_{1}(\operatorname{mod}Q_{1}^{2}),\widehat{\psi}(g_{2})\equiv 1+Q_{2}X_{2}(\operatorname{mod}Q_{2}^{2})$ for some primitive $X_{1},X_{2}\in\text{Lie}(\text{SL}_{2}(\mathbb{Z}))$. Collecting all powers of $g_{2}$, pulling them back to $G^{\prime\prime}$ and using $\widehat{\psi}$ is induced from $\pi_{(q_{5}^{\prime})^{\theta^{1/4}}}\circ\mathbb{P}_{1}\circ\psi$ by factoring through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, we obtain
$\pi_{q_{1}^{*}Q_{2}^{\{{4}/{3}\}},(q_{2}^{*}/q_{5}^{\prime})Q_{1}^{2}}(B^{2880^{2}})\cong\left(\pi_{q_{1}^{*},q_{2}^{*}/q_{5}^{\prime}}\times\pi_{Q_{2}^{\{4/3\}},Q_{1}^{2}}\right)(B^{2880^{2}})$
$\supset$ $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma(({q_{2}^{*}}/{q_{5}^{\prime}})^{\{\theta^{1/2}\}})/\Gamma({q_{2}^{*}}/{q_{5}^{\prime}})\times\left(\pi_{Q_{2}^{\{{4}/{3}\}},Q_{1}^{2}}\left((1,1)+\mathbb{Z}(Q_{2}X_{2},Q_{1}X_{1})\right)\right)$ (6.20)
with $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta^{1/4})$.
With (6.3) at hand, which plays the same role as Theorem B creating a one-parameter group, we run the same arguments in Section 5.5, and we give a sketch here. We first apply Proposition 4.2 with $Q=q_{5}^{\prime}$ to find elements from $A$ to conjugate $X=(X_{1},X_{2})$ to other directions. By considering five primitive linear forms, we can produce $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ such that the following holds: There exists $\bar{q_{5}}||q_{5}^{\prime}$ and
$\bar{q_{5}}>(q_{5}^{\prime})^{1-5\theta^{\frac{1}{2}}},$ (6.21)
such that
$(\bar{q_{5}})^{\{5\theta^{\frac{1}{2}}\}}\pi_{\bar{q_{5}}}(V\times V)\subset\pi_{\bar{q_{5}}}\text{Span}_{\mathbb{Z}}\{X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}\}$ (6.22)
as long as $\delta<c_{2}\eta\theta^{1/2}/2$.
Multiplying the RHS of (6.3) with its conjugates by $g_{1},g_{2},g_{3},g_{4},g_{5}$, followed by taking commutators of this product iteratively, and finally multiplying all these commutators together, we obtain
$\Gamma((q_{1}^{*}\bar{q_{5}}))^{\{5\theta^{1/2}\}}/\Gamma(q_{1}^{*}\bar{q_{5}})\times\Gamma(((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})^{\{5\theta^{1/2}\}})/\Gamma((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})$
$\subset$ $\pi_{q_{1}^{*}\bar{q_{5}},(q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}}}\{B\cup A\}^{O(\log 1/\theta)}$
and thus
$|\pi_{q_{1}{q_{5}^{\prime}},q_{2}}\{B\cup A\}^{O(\log 1/\theta)}|>(q_{1}^{*}q_{2}^{*}\bar{q_{5}})^{3(1-5\theta^{1/2})-}>(q_{1}q_{2}q_{5}^{\prime})^{3-O(\theta^{1/2})}$
In this case, we let $q_{3}^{*}={q_{5}^{\prime}}$ and the proposition is proved.
7. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$
Let $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ to be determined at (7.12). Let $A$ be given in Proposition 2.2 but we assume
$|\pi_{q}(A\cdot A\cdot A)|<|A|^{1+\delta}.$ (7.1)
We will arrive at a contradiction when $\delta$ is sufficiently small.
We divide our proof into two cases.
7.1. The case $q_{s}>q^{\frac{\epsilon}{2}}$
Take
$\delta<\frac{\log\frac{1}{c_{0}}}{2}$ (7.2)
for $c_{0}=c_{0}(L)$ the implied constant from Theorem A, so that
$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|<\frac{1}{2}$
for any $x\in\Gamma_{q_{s}}$. Since $\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(A)>q^{-\delta}$, we have
$|\pi_{q_{s}}(A)|>|\Gamma_{q_{s}}|q^{-\delta}>q_{s}^{6-\frac{2\delta}{\epsilon}-}>q_{s}^{6-2\rho}$ (7.3)
if we take
$\delta<\min\{\frac{\log\frac{1}{c_{0}}}{2},{\epsilon\rho}\}.$ (7.4)
for some $\rho\ll\epsilon$ to be determined at (7.9).
Furthermore, if $q_{s}>q^{1-\frac{\epsilon}{2}}$, then (7.3) implies the assumption in (2.3) is void and so Proposition 2.2 automatically holds. So now we assume $q_{s}<q^{1-\epsilon/2}$.
Now we apply Proposition 5.1 with $q$ replaced by $q_{l}$, then we obtain some $q^{\prime}||q_{l},q^{\prime}>q_{l}^{c}>q^{\frac{\epsilon c}{2}}$ for some $c$ only depending on the generating set $S$, and $C^{\prime}=O_{S}(\log\frac{1}{\rho})$ such that
$|\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C^{\prime}})|>(q^{\prime})^{3-\rho}$
if
$\delta<c^{\prime}\epsilon\rho$ (7.5)
for some $c^{\prime}$ depending only on $S$.
Now we apply Proposition 6.1 with $B=A^{C^{\prime}}$, $q_{1}=q_{2}=q_{s}$, $q_{3}=q^{\prime}$ and $\theta=\rho$. We then obtain $q^{\prime\prime}|q^{\prime},q^{\prime\prime}>(q^{\prime})^{\frac{1}{4}\times 10^{-4}}$, and
$|\pi_{q_{s}q^{\prime\prime},q_{s}}(A^{C^{\prime}C^{\prime\prime}})|>(q_{s}^{2}q^{\prime\prime})^{3-O(\theta^{\frac{1}{4}})},$ (7.6)
for some
$C^{\prime\prime}=O(\log{1}/{\rho}).$ (7.7)
By (7.7), we have increased the modulus of the first component from $q_{s}$ to $q_{s}q^{\prime\prime}$, at the cost of a density loss from $3-\rho$ to $3-O(\rho^{{1}/{4}})$. If $q_{s}q^{\prime\prime}<q^{1-\frac{\epsilon}{2}}$, we apply Proposition 5.1 again with $q$ replaced by $\frac{q}{q_{s}q^{\prime\prime}}$ and $\theta=O(\rho^{1/4})$, after that we can apply Proposition 6.1 again to increase the modulus to a larger one. We apply Proposition 5.1 and Proposition 6.1 iteratively until we reach a modulus $q_{1}^{\star}||q,q_{1}^{\star}>q^{1-\frac{\epsilon}{2}}$. Next, we go through the same procedure to increase the modulus of the second component to $q_{2}^{\star}||q,q_{2}^{\star}>q^{1-\frac{\epsilon}{2}}$. The number of such iterations is bounded by by $\frac{K_{0}\log{(1/\epsilon)}}{c}$ for some absolute constant $K_{0}$.
In the end, we obtain
$\left|\pi_{q_{1}^{\star},q_{2}^{\star}}\left(A^{O_{S}(\log\frac{1}{\rho})}\right)\right|>(q_{1}^{\star}q_{2}^{\star})^{3-O(\rho^{(\frac{1}{4})^{\frac{K_{0}\log(1/\epsilon)}{c}}})}.$ (7.8)
Let $K_{1},K_{2}$ be the implied constants on the LHS and RHS of (7.8), respectively. If we take
$\rho=e^{-(\log(3K_{2}/\epsilon)4^{{\frac{K_{0}\log(1/\epsilon)}{c}}})},$ (7.9)
then from (7.8),
$\left|\pi_{q_{1},q_{2}}\left(A^{K_{1}(\log\frac{1}{\rho})}\right)\right|>q^{6-5\epsilon}.$ (7.10)
Recall also $\pi_{q}(A)<q^{6-6\epsilon}$. Then (7.10) and (5.13) imply
$|\pi_{q}(A\cdot A\cdot A)|>|A|^{1+\frac{\epsilon}{6K_{1}\log(1/\rho)}}.$ (7.11)
We take any $\delta_{0}$ sufficiently small satisfying (7.2), (7.4) and (7.5) for $\delta$. Then we take
$L=\left[\frac{13}{\delta_{0}}\right]$ (7.12)
in view of the condition (6.3). Finally, we set
$\delta=\min\{\delta_{0},\frac{\epsilon}{6K_{1}\log(1/\rho)}\}.$
Then (7.11) contradicts (7.1). Proposition 2.2 is thus proved in this case.
7.2. The case $q_{s}<q^{\frac{\epsilon}{2}}$
In this case the modulus $q_{s}$ can be ignored. We can directly apply Proposition 5.1 and start the iterating process. The analysis is identical to the previous case.
8. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$
The proof of Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ is similar to the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. We give a sketch.
We write $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ depending only on the generating set $S$ and $\epsilon$.
We take $\delta$ sufficiently small and assume
$|\pi_{q}(A\cdot A\cdot A)|<|\pi_{q}(A)|^{1+\delta}.$ (8.1)
If $q_{s}>q^{1-\epsilon/2}$, $q_{l}$ can be ignored and we can apply Theorem A to conclude Proposition 2.2.
If $q^{\epsilon/2}<q_{s}<q^{1-\epsilon/2}$, we first apply Theorem A to find some $C_{1}$ so that $A^{C_{1}}(\operatorname{mod}q_{s})$ is large in $\Lambda/\Lambda(q_{s})$. Then we use the method in Section 7 to find some $C_{2}$ so that $\pi_{q,q_{s}}(A^{C_{2}})$ is large in $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q_{s}\mathbb{Z})^{2}$, and so by Proposition 3.20, there exists $q_{1}||q,q_{2}||(q_{1},q_{s}),q_{1}q_{2}>(qq_{s})^{1-\epsilon/2}$ such that
$\pi_{q_{1},q_{2}}(A^{C_{2}})\supset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ (8.2)
for some $\rho=O(\delta)$. Choose $B\subset A^{C_{2}}$ a representative set of RHS of (8.2) and consider the section map
$\psi:(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})\rightarrow B.$
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is not a homomorphism, then one can apply the argument in Section 6.1 find a large divisor $q_{3}||\text{gcd}(q_{l},q_{1})$ and an element $(g_{0},w_{0})$ from a bounded product of $B$ such that $\pi_{q_{1},q_{2}}((g_{0},w_{0}))=(1,0)$, and $v_{p}(w_{0})=O(\rho n^{1/2})$ for any $p^{n}||q_{3}$. Conjugating $(g_{0},w_{0})$ by $B$ using the formula $(g,x)\cdot(1,y)\cdot(g,x)^{-1}=(1,g(y)),$ and multiplying these conjugates by $B$, we obtain a set which is large in $\Lambda/\Lambda(q_{1})\ltimes(\mathbb{Z}/q_{2}q_{3})^{2}$ under the reduction $\pi_{q_{1},q_{2}q_{3}}$.
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is a homomorphism, then we can find a large divisor $q_{4}||\text{gcd}(q_{l},q_{1})$ and a large subset $Z\subset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ such that $\psi|_{Z}$ is a homomorphism. But then Proposition 3.24 implies $\psi(Z)(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})\subset H(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})$ for some proper algebraic subgroup $H\subset\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ defined by two linear equations. Then using Proposition 4.4 and arguing as in Section 5.1 for creating the set $A_{3}$, we can find a large exact divisor $q_{5}||q_{4}$ and $(g_{1},w_{1})$ from a bounded product of $B\cup(A\cdot A)$, such that $\pi_{q_{1},q_{2}}((g_{1},w_{1}))=(1,0)$, and $v_{p}(w_{1})=O(\rho^{1/2}n)$ for any $p^{n}||q_{5}$. With the help of $(g_{1},w_{1})$, we can boundedly generate a set which is large in $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}q_{5}\mathbb{Z})^{2}$.
In either of the above two cases, we increase the modulus $q_{2}$ by a factor of size at least $q^{\kappa}$ where $\kappa>0$ depending only on $S$ and $\epsilon$, Iterating the above procedure by controlled number of times, we can grow $q_{2}$ to a modulus $q_{6}||q_{1}$ and $q_{6}>q_{1}^{1-\epsilon/2}$, which will lead to a contradiction to (8.1) when $\delta$ is sufficiently small.
If $q_{s}<q^{\epsilon/2}$, then $q_{s}$ can be ignored and the expanding procedure is virtually identical to the previous case.

### v4

@@
 $0=\lambda_{0,q}>\lambda_{1,q}\geq\lambda_{2,q}\cdots$
-For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\epsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\epsilon$, $\forall q\in\mathcal{A}$.
+For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\varepsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\varepsilon$, $\forall q\in\mathcal{A}$.
 The existence of a uniform positive spectral gap only depends on $G$ and is independent of the choice of the finite generating set $S$.
 If $\mathcal{A}=\mathbb{Z}_{+}$, we simply say $G$ has super approximation.
-It is well known that a large class of lattices in $\text{GL}_{n}(\mathbb{Z})$ satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for arithmetic lattices and can not deal with a general non-lattice discrete group.
+It is well known that lattices in semisimple Lie groups satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for lattices and can not deal with a general non-lattice discrete group, which is also called a thin group.
 A breakthrough came from Bourgain and Gamburd, who developed the so-called “Bourgain-Gamburd expansion machine”, which is an analytic-combinatorial tool that allows them to prove super approximation property for any Zariski-dense subgroup of $\text{SL}_{2}(\mathbb{Z})$ with respect to prime moduli [BG08b].
 A critical ingredient in the proof is Helfgott’s triple product theorem [Hel08].
-Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [Gol19], [HdS21].
+Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [BGGT15], [PS16], [Gol19], [HdS21].
 Regarding the most general possible linear groups having super approximation, in [GV12] Salehi-Golsefidy and Varjú conjectures:
@@
 Then $G$ has the super approximation property if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e.
-$[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
+$\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
 In the same paper [GV12], Salehi-Golsefidy-Varjú proved that $G$ has super approximation with respect to square free numbers if $\mathbb{G}_{0}$ is perfect.
@@
 (Salehi-Golsefidy)[Gol19] Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated, then $G$ has the super approximation property with respect to bounded powers of square free integers coprime to $q_{0}$ if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e.
-$[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
-In the paper [Gol19], the author also claims super approximation for $G$ over $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$ for a fixed prime $p$.
-There appears an issue unless $\mathbb{G}_{0}$ has simple closure.
-To address this issue, we introduce a gluing technique, which turns out to be applicable to various other settings.
-We view this gluing technique may be the most novel part of this paper.
-More detailed discussion is given in Section 2.
+$\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
 If we require no restriction on moduli, all known results have restrictions on the Zariski closure.
 In [BV12] Bourgain-Varjú proved Conjecture 1.1 when the Zariski closure of $\mathbb{G}_{0}$ is $\text{SL}_{d}$, using an extra ingredient from homogeneous dynamics [BFLM11].
-Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with simple closure [HdS21].
+Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with $\mathbb{Q}$-simple closure [HdS21].
 The full generality of Conjecture 1.1, despite abundant evidence, has still remained a technical challenge.
-In the paper [Gol20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers.
+In the paper [SG20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers.
 In a separate paper [TZ23], we proved this conjecture:
 Theorem B (Tang-Zhang, 2023).
-Suppose $0<\delta\ll 1$, $d$ is a positive integer, and $N_{0}\gg_{d,\delta}1$ is a positive integer.
-Then there are $0<\varepsilon:=\varepsilon(\delta,d)$ and positive integers $C_{1}=C_{1}(\delta,d),C_{2}=C_{2}(\delta,d),C_{3}=C_{3}(\delta,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$.
+Suppose $0<\alpha<1$, $d$ is a positive integer, and $N_{0}\gg_{d,\alpha}1$ is a positive integer.
+Then there are $0<\varepsilon:=\varepsilon(\alpha,d)$ and positive integers $C_{1}=C_{1}(\alpha,d),C_{2}=C_{2}(\alpha,d),C_{3}=C_{3}(\alpha,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$.
 Suppose $\mathfrak{a}$ is an ideal of $\mathcal{O}$ such that $N(\mathfrak{a}):=|\mathcal{O}/\mathfrak{a}|\geq N_{0}$, and suppose $A\subseteq\mathcal{O}$ such that
-$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\delta}.$
+$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\alpha}.$
 Then there are an ideal $\mathfrak{a}^{\prime}$ of $\mathcal{O}$, and $a\in\mathcal{O}$ such that
@@
 }\pi_{\mathfrak{a}^{\prime}})$ is the reduction map $\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}\text{ (resp.
-}\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
+}\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}=\{a_{1}a_{2}\cdots a_{C_{2}}:a_{1},\cdots,a_{C_{2}}\in A\}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}=\{b_{1}+b_{2}+\cdots+b_{C_{3}}:b_{1},\cdots,b_{C_{3}}\in A^{C_{2}}\}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
 Roughly speaking, Conjecture 1.1 is a quantitative statement that given a set $A\subset\mathcal{O}$ and an ideal $\mathfrak{a}\subset\mathcal{O}$, if $|\pi_{\mathfrak{a}}(A)|$ is not too small, then modulo an ideal $\mathfrak{a}^{\prime}$ comparable to $\mathfrak{a}$, a sum-product set of $A$ contains a thick arithmetic progression.
-With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the spirit of [BG09].
-Combining Theorem B with a gluing technique which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
+With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the strategy of [BG09].
+Combining Theorem B with a gluing technique (Proposition 6.1) which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
 Theorem 1.2.
-Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation property with respect to all positive integers.
-Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple clousre case, and the first non-semisimple closure case.
+Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation with respect to all positive integers.
+Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple closure case, and the first non-semisimple closure case.
 One advantage of our approach using sum-product, compared to the method in [BV12] and [HdS21] is that we can allow denominators for the group.
@@
 Theorem 1.3.
-Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\times{\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})$ $({\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\ltimes(\frac{1}{q_{0}}\mathbb{Z})^{2}$, respectively)))), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)))), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
+Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])\times{\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])$ $({\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])\ltimes(\mathbb{Z}[1/q_{0}])^{2}$, respectively)), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
 Acknowledgements
-We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun and Tran Chieu-Minh for many helpful corrections/suggestions on previous versions of this paper.
+We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun, Tran Chieu-Minh and the anonymous referee for many helpful corrections/suggestions on previous versions of this paper.
 2. Notations and methodology
@@
 For given two subsets $A$ and $B$, we denote their product set by $A\cdot B=\{ab|a\in A,b\in B\}$, and their sum set by $A+B=\{a+b|a\in A,b\in B\}$.
-The $k$-fold product of $A$ is denoted by $A^{k}$, and the $k$-fold sum of $A$ is denoted by $\sum_{k}A$.
-If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by $f*g$ their convolution
+We let $A^{k}=\{a_{1}a_{2}\cdots a_{k}:a_{1},a_{2},\cdots,a_{k}\in A\}$, and $\sum_{k}A=\{a_{1}+a_{2}+\cdots+a_{k}:a_{1},a_{2},\cdots,a_{k}\in A\}$.
+If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by fgf*g their convolution
 $f*g(x)=\sum_{y\in G}f(y)g(xy^{-1}).$
 We write $f^{(l)}$ for the $k$-fold convolution of $f$ with itself.
-For a prime $p$, we write $p^{n}||q$ if $p^{n}|q$ but $p^{n+1}\nmid q$, and we define $v_{p}(q)=n$.
-Similarly, for a matrix $M$, $v_{p}(M)=n$ if $p^{n}|M$ but $p^{n+1}\nmid M$.
-For a prime ideal $\mathcal{P}$ in a ring of integers $\mathcal{O}$, we can define $v_{\mathcal{P}}$ in an analogous way.
-For two integers $q_{1}$ and $q_{2}$, we write $q_{1}||q_{2}$ if for every $p^{n}||q_{1}$, we also have $p^{n}||q_{2}$.
-The exact division ``||"``||"` ` | | " can be extended to ideals in a natural way.
-For a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of a ring $\mathcal{O}$, we write $\mathcal{P}^{n}||\mathfrak{a}$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$.
-Similarly, for two ideals $\mathcal{Q}_{1}\supset\mathcal{Q}_{2}$, we write $\mathcal{Q}_{1}||\mathcal{Q}_{2}$ if for all $\mathcal{P}^{n}||\mathcal{Q}_{1}$, we also have $\mathcal{P}^{n}||\mathcal{Q}_{2}$.
-For $q=\prod_{i}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a real positive number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $[n_{i}\alpha]$ is the integer part of $n_{i}\alpha$.
-Let $\pi_{q}:\mathbb{Z}\rightarrow q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
-Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the reduction map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$.
-Sometimes we need to reduce the two factors of $\Lambda$ by two different moduli $q_{1},q_{2}$, and we denote this reduction by $\pi_{q_{1},q_{2}}$.
-Throughout our paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$.
+For a prime $p$ and an integer $q$, we define $v_{p}(q)=n$ if $p^{n}|q$ but $p^{n+1}\nmid q$.
+Similarly, for a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of the ring of integers $\mathcal{O}$ of a number field $K$, we define $v_{\mathcal{P}}(\mathfrak{a})=n$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$.
+For notational convenience by writing $v_{\mathcal{P}}(a)$ for $a\in\mathcal{\mathcal{O}}$ we mean $v_{\mathcal{P}}((a))$.
+The valuation function is naturally extended to matrices: for a matrix $M\in\text{Mat}_{2}(K)$, and a prime ideal $\mathcal{P}$ in $\mathcal{O}$, $v_{\mathcal{P}}(M)=n$ if the minimum of $v_{\mathcal{P}}$-valuations of all coefficients of $M$ is $n$.
+The notion “exact division” is extremely helpful for the presentation in this paper.
+Given a natural number $q$, for each prime $p|q$, we write $p^{n}\|q$ if $v_{p}(q)=n$.
+Similarly, for two integers $q_{1}$ and $q_{2}$, we write $q_{1}\|q_{2}$ if for every $p|q_{1}$, we have $v_{p}(q_{1})=v_{p}(q_{2})$, and we call $q_{1}$ an exact divisor of $q_{2}$.
+For two ideals $\mathcal{Q}_{1},\mathcal{Q}_{2}$ of a ring of algebraic integers $\mathcal{O}$, we write $\mathcal{Q}_{1}\|\mathcal{Q}_{2}$ if for all prime ideals $\mathcal{P}\supset\mathcal{Q}_{1}$, we have $v_{\mathcal{P}}(\mathcal{Q}_{1})=v_{\mathcal{P}}(Q_{2})$.
+For an ideal $\mathcal{Q}\subset\mathcal{O}$, we say $\mathfrak{q}\in\mathcal{O}$ is a uniformizer of $\mathcal{Q}$ if $\mathfrak{q}\mathcal{O}/\mathcal{Q}^{2}=\mathcal{Q}/\mathcal{Q}^{2}$.
+For $q=\prod_{i\in I}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a positive real number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $n_{i}\alpha]$ is the integer part of $n_{i}\alpha$.
+For any non-empty subset $U\subset I$, we write $q^{U}=\prod_{i\in U}p_{i}^{n_{i}}$, which is an exact divisor of $q$.
+Let $\pi_{q}:\mathbb{Z}\rightarrow\mathbb{Z}/q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
+Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the residue map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$.
+Throughout this paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$ and $V=\mathfrak{sl}_{2}(\mathbb{Z})=\{\gamma\in\text{Mat}_{2}(\mathbb{Z}):\text{tr}(\gamma)=0\}$.
+Sometimes we need to reduce the two factors of $\Gamma\times\Gamma$ by two different moduli $q_{1},q_{2}$, and we denote this residue map by $\pi_{q_{1},q_{2}}$.
+We also use $\pi_{q_{1},q_{2}}$ to denote the reduction $\Gamma\ltimes\mathbb{Z}^{2}\rightarrow\Gamma_{q_{1}}\ltimes(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$ for $q_{2}|q_{1}$.
 We use the standard asymptotic notations $O,\Omega,\Theta$ to describe growth of functions.
-For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large.
-Similarly, $f(q)>q^{c-}$ to mean $f(q)>q^{c-\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large.
+For two positive functions $f,g$ and a set of parameters, we write $f=O_{X}(g)$ if there is a positive function $C_{1}(X)$ of the set of parameters such that $f\leq C_{1}(X)g$; we write $f=\Omega_{X}(g)$ if there is a positive function $C_{2}(X)$ such that $f\geq C_{2}(X)g$; we write $f=\Theta_{X}(g)$ if $f=O_{X}(g)$ and $f=\Omega_{X}(g)$.
+For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\varepsilon}$ for arbitrarily small $\varepsilon$ when $q$ large.
+Similarly, $f(q)>q^{c-}$ means $f(q)>q^{c-\varepsilon}$ for arbitrarily small $\varepsilon$ when $q$ large.
 Let $\chi_{S}$ be the normalized uniform counting measure supported on $S$, i.e., for $A\subset G$, $\chi_{S}(A)=\frac{|A\cap S|}{|S|}$.
@@
 Denote the set of eigenvalues of $T_{q}$ on $l_{0}^{2}(G_{q})$ by $E_{q}$.
-It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$.
 Then following the argument in [BV12] (Proof of Theorem 1, Page 156-158), Theorem 1.2 can be derived from the following proposition:
 Proposition 2.2.
-Let $S$ be a symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in ${\text{SL}_{2}\times\text{SL}_{2}}$ or ${\text{ASL}_{2}}$.
-Then for any $\epsilon>0$ there are $10<\delta<$ 1, $C>0$ depending on $\epsilon$ such that the following holds.
-If $q\in\mathbb{Z}_{+}$ sufficiently large, $A\subset\Lambda$ symmetric, and some integer $l$ satisfying
-$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\epsilon},$ (2.3)
+Let $S$ be a finite symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$} or $\text{ASL}_{2}$}.
+Then for any $\varepsilon>0$ there is $0<\delta<1$ depending on $\varepsilon$ such that the following holds.
+If $q\in\mathbb{Z}_{+}$ sufficiently large , $A\subset\Lambda$ symmetric, and some integer $l$ satisfy
+$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\varepsilon},$ (2.3)
 then
 $|\pi_{q}(A\cdot A\cdot A)|>|\pi_{q}(A)|^{1+\delta}.$ (2.4)
-It is noted that the reduction in [BV12] uses bounds for multiplicities of nontrivial irreducible representations appearing in the regular representation of $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$.
-Similar bounds exist for $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ and $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q\mathbb{Z})^{2}$ as well.
-In the rest of the paper we will focus on proving Proposition 2.2. We give a sketch here.
+We recall how to deduce Theorem 1.3 from Proposition 2.2 using the argument from [BV12], Page 156-158.
+It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$.
+Assume $S$ freely generates a group $G$; if not, applying Tits alternative [Tit72] or Proposition 20, [GV12], which generalizes Tits alternative from semisimple groups to perfect groups, we can find a finite symmetric set $S^{\prime}$ which freely generates a subgroup $G^{\prime}<G$, with $G^{\prime}$ Zariski dense in $\Lambda$.
+Theorem 1.3 for $S^{\prime}$ then implies Theorem 1.3 for $S$.
+Let $\lambda$ be an eigenvalue of $T_{q}$ and let $\mu$ be a corresponding eigenfunction.
+Consider the right regular representation of $\Lambda_{q}$ on $L^{2}(\Lambda_{q})$.
+Let $\rho$ be the irreducible representation that contains $\mu$.
+Assume $\rho$ is faithful, i.e., $\rho$ is not induced from a subrepresentation of the right regular representation of $\Lambda_{q^{\prime}}$ for some $q^{\prime}<q$; otherwise, replace $q$ by $q^{\prime}$.
+According to Proposition 19, [Gol19], if $\lambda$ is outside a finite set $\Sigma$, then $\rho$ has multiplicity at least $|\Lambda_{q}|^{\delta_{0}}$ for some positive $\delta_{0}$, and thus $\lambda$ has multiplicity at least $|\Lambda_{q}|^{\delta_{0}}$.
+We can now bound $\lambda^{2l}$ by computing the trace of $T^{2l}$ in the standard basis:
+$\lambda^{2l}\leq|\Lambda_{q}|^{-\delta_{0}}\text{Tr}(T^{2l})=|\Lambda_{q}|^{1-\delta_{0}}\|\pi_{q}[\chi_{S}^{(l)}]\|_{2}^{2}.$ (2.5)
+The Kesten’s bound implies there exists $l_{0}>c_{0}\log q$ for some small $c_{0}>0$, such that
+$\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}<|\Lambda_{q}|^{-\varepsilon_{0}}$
+for some $\varepsilon_{0}>0$.
+If
+$\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}\geq|\Lambda_{q}|^{-\frac{1}{2}+\frac{\delta_{0}}{4}},$
+then Proposition 2.2 implies
+$\|\pi_{q}[\chi_{S}^{(2l_{0})}]\|_{2}<\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}^{1-\delta}$
+for some $\delta>0$, with the help of a non-commutative analog of Balog-Szameredi-Gowers theorem due to Bourgain-Gamburd.
+See [BV12] for details.
+Keep iterating the convolution until we obtain some $l<C_{0}\log q$ such that
+$\|\pi_{q}[\chi_{S}^{(l)}]\|_{2}<|\Lambda_{q}|^{-\frac{1}{2}+\frac{\delta_{0}}{4}}.$ (2.6)
+Applying (2.6) to (2.5) leads to $\lambda<e^{-\frac{\delta_{0}}{C_{0}}}$ as desired.
 2.1. Sketch of proof for Proposition 2.2
-We assume (2.4) does not hold for a sufficiently small $\delta$. We will then show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\epsilon/2}$, which will force (2.4) to hold afterall.
-We assume $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and we let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components.
-For simplicity we assume each prime divisor of $q$ has sufficiently large exponent.
-On the first stage, we follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q^{\prime}$ of $q$ and some constant $C$ such that one simple factor, say $\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q^{\prime}}$ (Proposition 5.1).
+We explain the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
+The case $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ follows in a similar fashion.
+Assuming (2.4) does not hold for a sufficiently small $\delta$, we will show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\varepsilon/2}$, which will force (2.4) to hold after all.
+Let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components.
+If all prime divisors of $q$ have large exponents, we can follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q_{*}$ of $q$ and some constant $C$ such that for one simple factor, say $\pi_{q_{*}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q_{*}}$ (see Proposition 5.1).
 In this process, compared to [BG09], we apply Theorem A to take a shortcut to deduce some results on non-concentration of random walks at Archimedean place.
-However, for the purpose of proving Proposition 2.2, we need to take $q^{\prime}$ a very large divisor of $q$, and this leads to the following question:
-Suppose for two exact divisors $q_{1},q_{2}$ of $q$, we have a set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
-We need to deal with the case $q_{1},q_{2}$ coprime for the above question when growing $q^{\prime}$ to a very large divisor of $q$ for one simple factor, and we also need to consider the case $q_{1},q_{2}$ non co-prime when passing from one simple factor to two simple factors.
-In the paper [Gol19], the author also claims super approximation for a general group with perfect Zariski closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$.
-We are reluctant to cite this result unless the group has simple closure, as there seems to be an issue in the paper [SG17] where this claim is based on, and this issue already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$.
-By first applying sum-product which creates thick one-parameter groups in one simple factor while losing control of the other factor, and then taking commutators, the author essentially arrives at the situation of the aforementioned question for the case $q_{1}=q_{2}=p^{n}$.
-It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ is large implies $B$ is large, for which simple counter examples can be found.
-See the sentence “Lemma 71 and Lemma 34 imply Proposition 37” (Pg.
-7254, [SG17]) and trace from there.
-To address this question, our approach is as follows. For simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma_{q_{2}}$. We consider a section map
+However, for the purpose of proving Proposition 2.2, we need to take $q_{*}$ a very large divisor of $q$.
+This leads to the following question:
+Question 2.7.
+Suppose for two exact divisors $q_{1}=q^{I_{1}},q_{2}=q^{I_{2}}$ of $q=\prod_{i\in I}p_{i}^{n_{i}}$, $I_{1},I_{2}\subset I$, we have a symmetric set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
+Remark 2.8.
+We need to consider the case $q_{1},q_{2}$ coprime for Question 2.7 when growing $q_{*}$ to a very large divisor of $q$ for one simple factor.
+We also need to consider $q_{1},q_{2}$ non-coprime when passing from one simple factor to another.
+In the paper [SG17], the author claims super approximation for a general group with semisimple closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$.
+There is an issue, which already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$.
+In the proof, The author arrives at the situation of Question 2.7 for $q_{1}=q_{2}=p^{n}$.
+It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ being large implies $B$ is large, for which simple counter examples can be found (e.g.
+diagonal groups).
+This issue can be resolved if we also allow some extra help from the set $A$ given by Proposition 2.2.
+This is good enough for proving Proposition 2.2.
+We develop a gluing tool (Proposition 6.1) to address a more general situation of Question 2.7, which in our view is a main technical contribution of this paper.
+We give a (more detailed) sketch of solution to Question 2.7 here, which contains all essential ideas for Proposition 6.1.
+Related to our application, we assume $n_{i}$ is large for each $i\in I_{2}$ (say, $n_{i}>\theta^{-2}$ for a parameter $\theta$ introduced below), and assume $\mathbb{P}_{2}(B)\subset\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$, where $\tilde{q}_{2}$ is the square free part of $q_{2}$, so that $\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$ is a product of $p$-groups, and the loss of size from $\Gamma_{q_{2}}$ to $\Gamma(\tilde{q}_{2})/\Gamma(q_{2})$ is insignificant.
+We do not have any such requirement for $q_{1}$, e.g.
+$q_{1}$ can be square free.
+Without hiding key features, for simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma(\tilde{q}_{2})/\Gamma({q_{2}})$, where $\tilde{q}_{2}$ is the square free part of $q_{2}$.
+Our goal is to show that a bounded product of $B\cup A$ covers a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$.
+Consider a section map
 $\psi:\Gamma_{q_{1}}\rightarrow B$
-such that $\mathbb{P}_{1}\circ\psi$ is the identity map. Take a small parameter $10<\theta<$ 1, e.g. $\theta=\frac{1}{1000}$. According to a dichotomy (Proposition 3.1), there are two scenarios:
+such that $\mathbb{P}_{1}\circ\psi$ is the identity map.
+Take a small parameter $0<\theta<1$, e.g.
+$\theta=10^{-12}$.
+For each $i\in I_{2}$ (i.e.
+$p_{i}^{n_{i}}\|q_{2}$), we consider the map $\psi_{i}=\pi_{p^{[n\theta]}}\circ\mathbb{P}_{2}\circ\psi$.
+According to a dichotomy (Proposition 3.1), there are two scenarios:
   1. (1)
-    There exists a large exact divisor $q_{3}$ of $q_{2}$, and $x,y\in\Gamma_{q_{1}}$ such that for any $p^{n}||q_{3}$, $\psi(xy)\neq\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
+    There exists $\mathcal{G}_{j}\subset\Gamma_{q_{1}}\times\Gamma_{q_{1}},|\mathcal{G}_{j}|>10^{-4}|\Gamma_{q_{1}}|^{2}$, such that $\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)$, $\forall(x,y)\in\mathcal{G}_{j}$.
   2. (2)
-    There exists a large exact divisor ${q_{4}}$ of $q_{2}$, and a large subset $S$ of $\Gamma_{q_{1}}$ such that for any $x,y\in S$, any $p^{n}||q_{4}$, $\psi(xy)=\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
-In the first scenario, one can take $B^{\prime}$ to be the collection of all conjugates of $\psi(xy)\psi(x)^{-1}\psi(y)^{-1}$ by $B$ so that $\mathbb{P}_{1}(B^{\prime})=1$ and $\pi_{q_{3}}\circ\mathbb{P}_{2}(B^{\prime})$ is large, thus $\pi_{q_{1},q_{3}}(BB^{\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{3}}$.
-In the second scenario, we consider two cases.
-If for a large exact divisor $q_{4}^{\prime}$ of ${q_{4}}$, we have $\pi_{q_{3}^{\{\theta/2\}}}\circ\mathbb{P}_{2}\circ\psi=1$, then by taking commutator of $\psi(\Gamma_{q_{1}})$ iteratively one can create a set $B^{\prime\prime}$ such that $\pi_{q_{4}^{\prime}}\circ\mathbb{P}_{2}(B^{\prime\prime})=1$ and $\mathbb{P}_{1}(B^{\prime\prime})$ is large in $\Gamma_{q_{1}}$, so that $\pi_{q_{1},q_{4}^{\prime}}(BB^{\prime\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime}}$.
-If for a large exact divisor $q_{4}^{\prime\prime}$ of ${q_{4}}$ such that for any $p^{n}||q_{4}^{\prime\prime}$, we have $\pi_{q_{3}^{\{{\theta}/{2}\}}}\circ\mathbb{P}_{2}\circ\varphi\neq 1$, then we must have $q_{4}^{\prime\prime}||q_{1}$, and we can find a set $B^{\prime\prime\prime}\subset\text{Im}(\psi)$ such that $B^{\prime\prime\prime}$ contains all the preimage of a set $\mathcal{P}\subset\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$ under the following canonical projection:
-$\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}\cong\Gamma_{q_{1}/q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$
-where $Q_{1},Q_{2}|q_{4}^{\prime\prime}$ such that for each $p^{n}||q_{4}^{\prime\prime}$, $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta)$, and $\mathcal{P}=\{(1,1)+n(Q_{1}X_{1},Q_{2}X_{2})\}_{n\in\mathbb{Z}}$, where $X_{1},X_{2}$ are two traceless primitive matrices in $\text{Mat}(2,\mathbb{Z})$.
-Then one can apply quantitative non-concentration of random walks to find elements from $A$ to conjugates $(X_{1},X_{2})$ to different directions.
-Multiplying these conjugates together and taking commutators, we can recover a large set in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}$.
-This whole scheme is carried out in the proof of Proposition 6.1. One can then apply Proposition 6.1 iteratively to create a large subset of $\Lambda_{q}$, which forces (2.4) to hold afterall.
-The idea of considering section maps comes from Bourgain’s gluing scheme for $\mathbb{Z}_{q}$ (See Lemma 6.71 of [Bou08]).
-If the group is non-semismple, this method can also be applied to connect the semisimple part of the group to the unipotent part.
-See Section 8.
+    There exists $S_{j}\subset\Gamma_{q_{1}},|S_{j}|\geq\frac{99}{100}|\Gamma_{q_{1}}|$, such that over $S_{j}$, $\psi_{j}$ agrees with a homomorphism $h_{j}:\Gamma_{q_{1}}\rightarrow\Gamma_{p^{[n\theta]}}$.
+Split $I_{2}=J_{1}\sqcup J_{2}$, where $J_{1}$ is the collection of indices when Case (1) happens, and $J_{2}$ is the complement.
+We further split $J_{2}=J_{3}\sqcup J_{4}$ where $J_{3}$ is the set of indices such that $h_{j}$ is trivial at the half level, i.e.
+$\pi_{p_{j}^{[\frac{1}{2}{n_{j}\theta}]}}\circ h_{j}=1$, and $J_{4}$ is the complement.
+Case 1: $q^{J_{1}}\geq(q_{2})^{\frac{1}{2}}$.
+A probability argument shows that there is a large subset $\mathcal{G}\subset\Gamma_{q_{1}}\times\Gamma_{q_{1}}$ (in fact $|\mathcal{G}|>q_{2}^{0-}|\Gamma_{q_{1}}|^{2}$), and $U_{0}\subset J_{1}$ with $q^{U_{0}}\geq(q^{J_{1}})^{\frac{1}{2}10^{-4}}$, such that $\psi_{j}(g_{1})\psi_{j}(g_{2})\neq\psi_{j}(g_{1}g_{2}),\forall(g_{1},g_{2})\in\mathcal{G},\forall j\in U_{0}$.
+Take any $(g_{1},g_{2})\in\mathcal{G}$.
+Let $w=\psi(g_{1})\psi(g_{2})(\psi(g_{1}g_{2}))^{-1}$, and $w_{B}=\{[[w,b_{1}],b_{2}]:b_{1},b_{2}\in B\}$.
+From Lemma 3.39 and Lemma 3.33, we have $\mathbb{P}_{1}((w_{B})^{N_{1}})=1$ and $\pi_{q^{U_{0}}}\circ\mathbb{P}_{2}((w_{B})^{N_{1}})$ contains a large congruence subgroup of $\Gamma_{q^{U_{0}}}$, for some $N_{1}=O(\log\frac{1}{\theta})$.
+Therefore, $\pi_{q_{1},q^{U_{0}}}(B\cdot(w_{B})^{N_{1}})$ covers a large subset of $\Gamma_{q_{1}}\times\Gamma_{q^{U_{0}}}$.
+Case 2.1: $q^{J_{2}}\geq(q_{2})^{\frac{1}{4}}$.
+Similar to Case 1, we can find a large subset $W_{1}\subset\Gamma_{q_{1}}$, and $U_{1}\subset J_{2}$ with $q^{U_{1}}\geq(q^{J_{2}})^{\frac{99}{200}}$, such that $\pi_{p_{j}^{[\frac{1}{2}{n_{j}\theta}]}}(g)\circ\psi_{j}=1,\forall g\in W_{1},\forall j\in U_{1}$.
+Then applying Lemma 3.33, for some $N_{2}=O(\log\frac{1}{\theta})$, $\mathbb{P}_{1}([W_{1},W_{1}]^{N_{2}})$ covers a large congruence subgroup of $\Gamma_{q_{1}}$, while $\pi_{p_{j}^{2[\frac{1}{2}n_{j}\theta]}}\circ\mathbb{P}_{2}([W_{1},W_{1}]^{N_{2}})=1,\forall j\in U_{1}$, noticing that we have increased the exponent of $p_{j}$ by a factor of 2 after taking commutator.
+Keep iterating until we obtain a set $B_{1}\subset[W_{1},W_{1}]^{O((\log\frac{1}{\theta})^{2})}$ such that $\mathbb{P}_{1}(B_{2})$ remains large in $\Gamma_{q_{1}}$, but $\mathbb{P}_{2}(B_{1})=1(\operatorname{mod}q^{U_{1}})$.
+Then, $\pi_{q_{1},q^{U_{1}}}(B\cdot B_{1})$ covers a very large subset of $\Gamma_{q_{1}}\times\Gamma_{q^{U_{1}}}$.
+Case 2.2: $q^{J_{3}}\geq(q_{2})^{\frac{1}{4}}$.
+In this case, we can find a large set $W_{2}\subset\Gamma_{q_{1}}$ and $U_{2}\subset J_{3}$, such that over $W_{2}$, $\pi_{p_{j}^{[n_{j}\theta]}}\circ\mathbb{P}_{2}\circ\psi=h_{j},\forall j\in U_{2}$.
+The set $W_{2}$ boundedly generates a large subgroup $G$ of $\Gamma_{q_{1}}$.
+The homomorphisms $\{h_{j}:j\in U_{2}\}$ uniquely determine a homomorphism $h:\Gamma_{q_{1}}\rightarrow\Gamma_{{(q^{U_{2}})^{\{\theta\}}}}$ such that $\pi_{p_{j}^{[n_{j}\theta]}}\circ h=h_{j},\forall j\in U_{2}$.
+Thus we can construct a section map
+$\bar{\psi}:G\rightarrow W_{2}^{O(1)},$
+such that $\mathbb{P}_{1}\circ\bar{\psi}$ is the identity map, and $\pi_{(q^{U_{2}})^{\{\theta\}}}\circ\mathbb{P}_{2}\circ\bar{\psi}=h$.
+The homomorphism $h$ is completely determined by homomorphisms between local factors of $G$ and $\Gamma_{(q^{U_{2}})^{\{\theta\}}}$.
+Also recall $\mathbb{P}_{2}(B)=\Gamma({\tilde{q}_{2}})/\Gamma(q_{2})$, which implies $h_{j}(G)$ is a $p_{j}$-group for each $j\in U_{2}$.
+Thus according to Lemma 3.44, since $h_{j}$ is nontrivial for each $j\in U_{2}$, we must have $j\in I_{1}$, so $U_{2}\subset I_{1}$.
+By working locally first and then applying Chinese Remainder Theorem, we can find an element $g_{0}\in G$ such that $\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(g_{0})=1$ and $\pi_{p_{j}^{[\frac{1}{2}n_{j}\theta]}}\circ\psi_{j}(g_{0})\neq 1,\forall j\in U_{2}$.
+Lemma 3.44 implies $v_{p_{j}}(g_{0}-1)=O(\theta n_{j}),\forall j\in U_{2}$ (More precisely, this estimate holds for almost all $j\in U_{2}$, with exceptions coming from those $j$ where $\pi_{p_{j}^{n_{j}}}(G)$ is not large, which are not many since $G$ is a large subgroup of $\Gamma_{q_{1}}$.
+Let us ignore this minor issue here).
+Let $g=g_{0}^{(q^{U_{2}})^{\{\frac{1}{4}\theta\}}}$. Then $\{\bar{\psi}(g^{n}):n\in\mathbb{N}\}$ satisfies
+$\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(\bar{\psi}(g^{n}))=1$
+$\bar{\psi}(g^{n})\equiv(1+nQ_{1}X_{1},nQ_{2}X_{2})(\operatorname{mod}Q_{1}^{2},Q_{3})$ (2.9)
+where $X_{1},X_{2}\in\text{Mat}_{2}(\mathbb{Z})$ traceless and primitive, $Q_{1},Q_{2}|q^{U_{2}}$ satisfies $v_{p}(Q_{1})=v_{p}(g-1)$, $v_{p}(Q_{2})=v_{p}(\mathbb{P}_{2}\circ\bar{\psi}(g)-1)$, which uniquely determines $Q_{1}$ and $Q_{2}$, and $Q_{3}=\text{gcd}(Q_{2}^{2},(q^{U_{2}})^{\{\theta\}})$, where we need $\pi_{Q_{3}}\circ\mathbb{P}_{2}\circ\bar{\psi}$ is a homomorphism.
+Importantly, we have the following valuation estimates:
+$v_{p_{j}}(Q_{1}),v_{p_{j}}(Q_{2}),v_{p_{j}}(Q_{3}),v_{p_{j}}(Q_{3})-v_{p_{j}}(Q_{2})=\Theta(n_{j}\theta),\hskip 14.22636pt\forall j\in U_{2},$
+where the implied constants for $\Theta(\cdot)$ are absolute.
+This explains our choice of $g$.
+Then (2.1) gives a “thick” one parameter group whose projection to each component is thick.
+We can then use random walks (Proposition 4.2) to find $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ to conjugate $X=(X_{1},X_{2})$ to different directions so that
+$\text{gcd}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}),q^{U_{2}})$
+is small.
+Multiplying $\{\bar{\psi}(g^{n}):n\in\mathbb{N}\}$ with all of its $g_{i}$-conjugates together, $1\leq i\leq 5$, and taking a further product set yield a set $B_{2}\subset\{B\cup A\}^{O(\log\frac{1}{\theta})}$ such that $\pi_{q^{I_{1}-U_{2}}}\circ\mathbb{P}_{1}(B_{2})=1$ and $\pi_{q^{U_{2}},q^{U_{2}}}(B_{2})$ is large in $\Gamma_{q^{U_{2}}}\times\Gamma_{q^{U_{2}}}$.
+Therefore, $\pi_{q_{1},q^{U_{2}}}(B\cdot B_{2})$ is a very large subset in $\Gamma_{q_{1}}\times\Gamma_{q^{U_{2}}}$.
+In each of the above case, we manage to find a set from a bounded (depending on $\theta$) product of $B\cup A$ whose reduction by $(q_{1},q^{U_{i}}),0\leq i\leq 2$ is large, compared to the set $B$ whose reduction is known to be large at $(q_{1},1)$.
+It is noted that the progress of modulus increase, measured by $\frac{\log q^{U_{i}}}{\log q_{2}}$ is bounded below by an absolute positive constant.
+One can then iterate the above gluing process boundedly many times to improve the modulus of the second component to a very large exact divisor of $q_{2}$, thus giving an affirmative answer to Question 2.7 (with extra help from AA!).
+Regarding the proof of Proposition 2.2, write $q=q_{s}q_{l}$ where all exponents of prime factors of $q_{s}$ are small and all exponents of prime factors of $q_{l}$ are large.
+If $q_{s}$ is not too small, we can first apply Theorem A to show $\pi_{q_{s}}(A)$ is big in $\Gamma_{q_{s}}\times\Gamma_{q_{s}}$, and iteratively apply Proposition 5.1 (which helps to create large sets under smaller moduli) and Proposition 6.1 (which helps to glu) to improve the moduli, until we find a bounded product of $A$ covers a large congruence subgroup of $\Lambda_{q}$ of size at least $|\Lambda_{q}|^{1-\frac{\varepsilon}{2}}$.
+This forces (2.4) to hold for a sufficiently small choice of $\delta$.
+If $q_{s}$ is too small, we can just work with $q_{l}$ and directly start the iterative process.
 Remark 2.10.
-Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can skip Section 5 and start from Section 6, and reference back to more or less standard lemmas in Sections 3 and 4 when necessary.
-We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3) and for potential further generalization.
+The discussions of Case 1, Case 2.1, Case 2.2 correspond to the analysis in §6.1, §6.2 and §6.3 for the proof of Proposition 6.1.
+Caution: Some notations appearing above, such as $\mathbb{P}_{1},\mathbb{P}_{2},Q_{1},Q_{2},Q_{3},W_{1},W_{2},G$ have different meanings in §6.
+Remark 2.11.
+One can start reading from Section 5 and reference back to more or less standard lemmas in Sections 3 and 4 when necessary.
+Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can assume Proposition 5.1 and skip the rest part in Section 5.
+We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3).
 3. Preliminaries on combinatorics
@@
 Proposition 3.1.
-Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
+Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ be some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
 $\left|\left\{(x,y)\in G_{1}\times G_{1}\mid\psi(xy)=\psi(x)\psi(y)\right\}\right|<\left(1-\varepsilon\right)\left|G_{1}\right|^{2},$ (3.2)
@@
 $\psi|_{S}=f|_{S}.$ (3.3)
-Remark 3.4.
-If $\text{gcd}\left(\left|G_{1}\right|,\left|G_{2}\right|\right)=1$, then (3.3) implies
-|$\left|\left\{x\in G_{1}\mid\psi(x)\neq 1\right\}\right|<\sqrt{\varepsilon}\left|G_{1}\right|.$
 Proposition 3.1 relies on the following Lemma 3.4 and Theorem 3.5.
 Lemma 3.4.
-Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $40<\varepsilon<1/$ 4, such that
+Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $0<\varepsilon<1/4$, such that
 $|\mathcal{G}|>(1-\varepsilon)|A|^{2}$
@@
 where $A\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}A:=\{ab\mid a,b\in A,(a,b)\in\mathcal{G}\}$.
-Proof of Lemma.
-3.5. The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
+Proof of Lemma 3.4.
+The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
 Theorem 3.5 (Noncommutative Freiman-Kneser theorem for small doubling).
@@
 Denote
-$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2}$
-and let $\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}$.
-Then $\mathcal{G}^{\prime}$ is a graph on $A$ and $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\epsilon)|G_{1}|^{2}=(1-\epsilon)|A|^{2}$.
-Apply Lemma 3.5 with $Z=G_{1}\times G_{2}$.
-We obtain a subset $A^{\prime}\subset A$ satisfying
-$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.8)
+$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2},$
+and
+$\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}.$
+Then $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\varepsilon)|G_{1}|^{2}=(1-\varepsilon)|A|^{2}$.
+Applying Lemma 3.4 with $Z=G_{1}\times G_{2}$ to the set $A$, we obtain a subset $A^{\prime}\subset A$ satisfying
+$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|A\right|=(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.7)
 and
@@
 $|H|$ $<\frac{5}{3}\left|A^{\prime}\right|.$ (3.10)
-[…]
-Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
+Let $\mathbb{P}_{1}$ be the projection map $G_{1}\times G_{2}\rightarrow G_{1}$. Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
 […]
@@
 Since $H$ is a subgroup, we get $(1,1)\in H$ so $f(1)=1$.
-Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$, so $f$ is a group homomorphism.
+Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$ by the uniqueness of choice for $f(y_{1}y_{2}^{-1})$, so $f$ is a group homomorphism.
+Therefore, over $A^{\prime}$ we have
+$(x,\psi(x))=(x_{1},x_{2})\cdot(x_{1}^{-1}x,f(x_{1}^{-1}x))\Rightarrow\psi(x)=x_{2}f(x_{1}^{-1}x).$ (3.11)
 Since
@@
 $\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|\geq|\mathcal{G}|+|\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)|-|G_{1}|^{2}>\left(1-2\sqrt{\varepsilon}\right)\left|G_{1}\right|^{2}$
-by inclusion-exclusion; and hence
+by inclusion-exclusion. Hence
 $\left|\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right|\geq\frac{\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|}{|G_{1}|}>(1-2\sqrt{\varepsilon})\left|G_{1}\right|.$
-Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$, and for $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
-$x_{2}f(x_{1}^{-1}y_{1}y_{2})=\psi\left(y_{1}y_{2}\right)=\psi\left(y_{1}\right)\psi\left(y_{2}\right)=x_{2}f(x_{1}^{-1}y_{1})x_{2}f(x_{1}^{-1}y_{2}).$
+Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$.
+For $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
+[…]
 $\Rightarrow f(x_{1})=x_{2}\Rightarrow(x_{1},x_{2})\in H\Rightarrow A^{\prime}\subset H.$
 We finish the proof of Proposition 3.1 by taking $S=\mathbb{P}_{1}\left(A^{\prime}\right)$. ∎
-We also need the following bounded generation result over $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
+We also need the following bounded generation result over congruence quotients of $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
 Proposition 3.12.
@@
 In [Hel08] Helfgott proved that given $A\subset\text{SL}_{2}(\mathbb{F}_{p})$ with $|A|>p^{\frac{8}{3}}$, then $A\cdot A\cdot A={\text{SL}}_{2}(\mathbb{F}_{p})$ for a sufficiently large prime $p$.
-Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to adapt well for a general composite modulus.
+Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to have an easy generalization for a general composite modulus.
 We follow Helfgott’s approach for the proof of Proposition 3.12.
@@
 Proof of Proposition 3.12.
-.
 Since $|A|>(q_{1}q_{2})^{3-\delta}$, by the pigeon hole principle, there exists $\vec{v}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{2},\vec{w}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$, such that the cardinality of the set
-$\{(\gamma_{1},\gamma_{2}):\gamma_{1}\text{ has lower row }\vec{v},\gamma_{2}\text{ has lower row }\vec{w}\}$
+[…]
 exceeds $(q_{1}q_{2})^{1-\delta}$. This implies the cardinality of the set
@@
 Then, $|A_{2}|>(q_{1}q_{2})^{1-\delta}$.
-Define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as:
-$(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$
-if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$.
-There are at most $q_{1}q_{2}$ many such classes.
-By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta}$ many elements from $A$.
+Next, we define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as: $(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$ if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler multiple in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler multiple in $(\mathbb{Z}/q_{2}\mathbb{Z})^{*}$.
+There are at most $(q_{1}q_{2})^{1+}$ many such classes.
+By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta-}$ many elements from $A$.
 This implies the set
 $H_{0}=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},x\in\mathbb{Z}/q_{1}\mathbb{Z},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*},y\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}$
-has cardinality $>(q_{1}q_{2})^{2-\delta}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
+has cardinality exceeding $(q_{1}q_{2})^{2-\delta-}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
 $H=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*}\right\}$
-has cardinality $>(q_{1}q_{2})^{1-\delta}$.
+has cardinality $(q_{1}q_{2})^{1-\delta}$.
 We have the following elementary computation
-$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
-Using (3.18) and applying Lemma 3.14 to the set
+$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
+Using (3.18) and applying Lemma 3.14 to the sets
 $\left\{(\lambda_{1}^{2},\lambda_{2}^{2}):\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right)\in H\right\}$
@@
 can be written as a $a_{1}b_{1}a_{2}b_{2}a_{3}$, where $a_{1},a_{2},a_{3}\in\left(\begin{matrix}1&p^{m}\mathbb{Z}/p^{n}\mathbb{Z}\\ 0&1\end{matrix}\right)$ and $b_{1},b_{2}\in\left(\begin{matrix}1&0\\ p^{m}\mathbb{Z}/p^{n}\mathbb{Z}&1\end{matrix}\right)$.
-From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{*})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{*})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
+From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{\prime})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{\prime})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
 $(A\cdot A^{-1})^{1440}\supset\mathrm{SL}_{2}(q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}).$
@@
 Then there exists $q^{\prime}|q,q^{\prime}<q^{10\delta}$ such that
-$A^{720}\supset\Gamma(q^{\prime})/\Gamma(q).$
+$A^{2880}\supset\Gamma(q^{\prime})/\Gamma(q).$
 We need Lemma 3.19 to prove the following bounded generation result for $\text{ASL}_{2}$.
@@
 Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}<q_{1}^{50\delta},q_{2}^{\prime}<q_{1}^{55\delta}$, such that
-$A^{7204}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
+$A^{28804}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
 Proof.
@@
 $(\gamma_{1},v_{1})\sim(\gamma_{2},v_{2})\text{ if and only if }\gamma_{1}^{-1}v_{1}=\gamma_{2}^{-1}v_{2}.$
-Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is bounded by $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$.
-Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 00.
+Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$.
+Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 0.
 Applying Lemma 3.19, we have
-$A^{1440}\supset(Bb^{-1})^{720}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
+$A^{5760}\supset(Bb^{-1})^{2880}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
 for some $q_{1}^{\prime}<q_{1}^{50\delta}$.
-By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$.
+By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$ elements.
 Take any $b^{\prime}\in B^{\prime}$.
@@
 $\Gamma(p^{m})/\Gamma(p^{n})v-\Gamma(p^{m})/\Gamma(p^{n})v\supset(p^{m}\mathbb{Z}/p^{n}\mathbb{Z})^{2}.$ (3.22)
-Let $L\subset A^{2882}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
-$A^{5764}\supset LL^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
+Let $R\subset A^{11522}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
+$A^{23044}\supset RR^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
 Multiplying the RHS of (3.21) and the RHS of (3.23) we obtain Proposition 3.20 with $q_{2}^{\prime}=\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})$.
-Let $u,v\in\mathbb{Z}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
+Let $u,v\in\mathbb{Q}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
 $f_{u,v}(A,B,C,D,E,F)=\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)-1\right)\left(\begin{matrix}u\\ v\end{matrix}\right)-\left(\begin{matrix}E\\ F\end{matrix}\right).$
-It is an elementary exercise to check
-$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg{|}f_{u,v}=0\right\}.$
-is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
-The following lemma shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$.
+It is straightforward that
+$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg|f_{u,v}(A,B,C,D,E,F)=0\right\}.$
+is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ as $H_{u,v}$ is the stabilizer of $(u,v)^{t}$.
+The following proposition shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$ for some $u,v$.
 Proposition 3.24.
-Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\Gamma(p^{m})/\Gamma(p^{n})|$.
+Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\mathbb{P}_{0}(H)|$.
 Then there exist some $u,v\in\frac{1}{p^{m}}\mathbb{Z}$ such that
-$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})\subset H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})$
+$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{[\frac{n}{4}]})\subset H_{u,v}(\operatorname{mod}p^{[\frac{n}{4}]})$
 Proof.
-Lemma 3.24 follows from proofs of a series of claims.
+Proposition 3.24 follows from proofs of a series of claims.
 Claim 1: Suppose $(g,w)\in H$. If $p^{k}|g-1$, then $p^{k-m}|w$.
@@
 with $g^{p}\in\Gamma(p^{k+1})/\Gamma(p^{n})$ and
-$(1+g+g^{2}+\cdots+g^{p-1})w\equiv(g^{p},p(1+\frac{(p-1)}{2}p^{k}M)w)(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
+$(1+g+g^{2}+\cdots+g^{p-1})w\equiv p(1+\frac{(p-1)}{2}p^{k}M)w(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
 By induction hypothesis,
-$(1+\frac{p-1}{2}p^{k}M)pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
+$p(1+\frac{p-1}{2}p^{k}M)w\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
 Since $1+\frac{p-1}{2}p^{k}M$ is non-singular $\operatorname{mod}p$, we get
@@
 Applying Claim 1 to $A_{4}A_{5}^{-1}$, we obtain $3b_{1}\equiv 0(\operatorname{mod}p^{k})$. Similarly, $3a_{2}\equiv 0(\operatorname{mod}p^{k})$. Claim 2 thus follows.
-Claim 3: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
+Claim 3: If $1\leq m\leq n/4$, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{[\frac{n}{4}]}}$.
 Proof of Claim 3. consider
 $A_{6}=A_{1}^{p^{t}}=\left(\left(\begin{matrix}1&p^{m+t}\\ 0&1\end{matrix}\right),\left(\begin{matrix}\cdots\\ p^{t}v_{1}\end{matrix}\right)\right)\in H.$
-If taking $t=\lfloor\frac{n}{4}\rfloor-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}.$ This finishes the proof of Claim 3.
+If taking $t=[\frac{n}{4}]-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{[\frac{n}{4}]}}$. This finishes the proof of Claim 3.
 By the same consideration, we obtain
-Claim 4: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
-Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$ since
-$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor}),$
+Claim 4: If $1\leq m\leq n/4$, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{[\frac{n}{4}]}}$.
+Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$, since
+$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{[\frac{n}{4}]}),$
 and that $\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right)$ and $\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right)$ generate $\Gamma(p^{2m})/\Gamma(p^{n})$. ∎
+Finally, we record some elementary combinatorics on $\text{SL}_{2}$.
+Lemma 3.25.
+Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime.
+Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{1}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$.
+Then
+$1+2p^{m_{1}+n_{1}}V\subset[H_{1},H_{2}]^{3}(\operatorname{mod}p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}).$ (3.26)
+To prove Lemma 3.25, we first need
+Lemma 3.27.
+Let $p$ be a prime. Let $A,B\in\mathbb{Z},m_{1},m_{2}\in\mathbb{N}$ such that
+$\pi_{p^{m_{1}}}(A)\supset\mathbb{Z}/p^{m_{1}}\mathbb{Z}$
+$\pi_{p^{m_{2}}}(B)\supset\mathbb{Z}/p^{m_{2}}\mathbb{Z}.$
+Then
+$\pi_{p^{m_{1}+m_{2}-1}}\left(\sum_{2}AB\right)\supset\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}.$
+Proof of Lemma 3.27.
+Take $a\in A$ such that $v_{p}(a)=m_{1}-1$, then
+$aB(\operatorname{mod}p^{m_{1}+m_{2}-1}\mathbb{Z})=p^{m_{1}-1}\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}.$ (3.28)
+Take $b\in B$ such that $v_{p}(b)=0$, then
+$Ab(\operatorname{mod}p^{m_{1}-1}\mathbb{Z})=\mathbb{Z}/p^{m_{1}-1}\mathbb{Z}.$ (3.29)
+(3.28) and (3.29) then implies $aB+Bb(\operatorname{mod}p^{m_{1}+m_{2}-1})=\mathbb{Z}/p^{m_{1}+m_{2}-1}\mathbb{Z}$. ∎
+Proof of Lemma 3.25.
+Take $g_{1}\in H_{1}$ and $g_{2}\in H_{2}$. Write
+$g_{1}=1+p^{m_{1}}\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&-a_{1}\end{pmatrix}+p^{2m_{1}}M_{1},$
+$g_{2}=1+p^{m_{2}}\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&-a_{2}\end{pmatrix}+p^{2m_{2}}M_{2},$
+with $a_{1},b_{1},c_{1}\in[0,p^{m_{1}}-1]$, $a_{2},b_{2},c_{2}\in[0,p^{m_{2}}-1]$, and $M_{1},M_{2}\in\text{Mat}_{2}(\mathbb{Z})$.
+By Lemma 5.66, we get
+$g_{1}g_{2}g_{1}^{-1}g_{2}^{-1}$ $\equiv 1+[g_{1},g_{2}]\equiv 1+p^{m_{1}+n_{1}}\bigg[\begin{pmatrix}a_{1}&b_{1}\\ c_{1}&-a_{1}\end{pmatrix},\begin{pmatrix}a_{2}&b_{2}\\ c_{2}&-a_{2}\end{pmatrix}\bigg$]
+$=1+p^{m_{1}+n_{1}}\begin{pmatrix}b_{1}c_{2}-b_{2}c_{1}&2(a_{1}b_{2}-b_{1}a_{2})\\ 2(c_{1}a_{2}-a_{1}c_{2})&-(b_{1}c_{2}-b_{2}c_{1})\\ \end{pmatrix}\pmod{p^{m_{1}+n_{1}+\min\{m_{1},n_{2}\}}}$.
+If we take $a_{1},a_{2}=0$ and $b_{1},c_{1},b_{2},c_{2}$ arbitrary, then we can apply Lemma 3.27 to obtain
+$[H_{1},H_{2}]\supset 1+p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}}$. (3.30)
+Similarly, by setting $b_{1}=0$ and $b_{2}=0$, we can get
+$[H_{1},H_{2}]$ $\supset 1+2p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}0&0\\ 1&0\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}}$, (3.31)
+and by setting $c_{1},c_{2}=0$, we get
+$[H_{1},H_{2}]$ $\supset 1+2p^{m_{1}+n_{1}}\mathbb{Z}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}\pmod{p^{\min\{m_{2}+n_{2}-1,2m_{1}+n_{1},m_{1}+2n_{1}\}}})$ (3.32)
+Then (3.26) follows from (3.30), (3.31) and (3.32). ∎
+Lemma 3.33.
+Let $p$ be a prime, $a,b\in\mathbb{N}$ and $b\geq 2a\geq 0$. Suppose $H\subset\text{SL}_{2}(\mathbb{Z})$ and $\Gamma(p^{a})(\operatorname{mod}p^{b})=H(\operatorname{mod}p^{b})$.
+  1. (1)
+    If $a\neq 0$, then if $p\neq 2$,
+    $[H,H]^{O(\log\frac{b}{a})}(\operatorname{mod}p^{b})=\Gamma(p^{2a})/\Gamma(p^{b}).$ (3.34)
+    If $p=2$,
+    $[H,H]^{O(\log\frac{b}{a})}(\operatorname{mod}2^{b})=\Gamma(2^{2a+1})/\Gamma(2^{b}).$ (3.35)
+  2. (2)
+    If $a=0$, then
+    $[H,H]^{O(\log b)}(\operatorname{mod}p^{b})=\Gamma_{p^{b}}.$ (3.36)
+Proof.
+Assume $p\neq 2$.
+If $a\neq 0$, let $s_{0}$ be the largest integer such that $1.5^{s_{0}}<\frac{b}{a}$.
+Apply Lemma 3.25 to $H_{1}=H_{2}=H$ with $m_{1}=n_{1}=[a\cdot 1.5^{j}]$, $m_{2}=n_{2}=\min\{2m_{1},b\}$, $0\leq j\leq s_{0}$.
+Multiplying the implied sets together yields (3.34).
+If $a=0$, Then we have
+$[H,H]^{3}(\operatorname{mod}p)=\Gamma_{p}.$ (3.37)
+See Theorem 1.1 of [Sha09].
+Let $H_{0}\subset[H,H]^{3}$ be the implied set of representatives of $\Gamma_{p}$.
+It follows from Lemma 4 of [BV12] that $H_{0}(\operatorname{mod}p^{2})$ is not a group, so there is $h_{1},h_{2},h_{3}\in H_{0}$ such that $h_{1}h_{2}h_{3}^{-1}\equiv 1(\operatorname{mod}p)$ but $h_{1}h_{2}h_{3}^{-1}\not\equiv 1(\operatorname{mod}p^{2})$.
+From here it is easy to deduce that $H_{0},H_{0}]^{O(1)}(\operatorname{mod}p^{2})=\Gamma_{p^{2}}$.
+We also know
+$[H,H]^{O(\log b)}(\operatorname{mod}p^{b})\supset\Gamma_{p^{b}}$ (3.38)
+by the $a=1$ case of (3.34). Combining (3.37) and (3.38) gives (3.36).
+The proof for the $p=2$ case is analogous.
+Lemma 3.39.
+Let $p$ be a prime and $b\geq 2a\geq 1$.
+Suppose $\gamma_{0}\in\text{SL}_{2}(\mathbb{Z})$ with $v_{p}(\gamma_{0}-1)=a$, and $H\subset SL_{2}(\mathbb{Z})$ with $H\supset 1+p^{b}V(\operatorname{mod}p^{2b})$.
+Then
+$\gamma_{0},H],H]^{2}\supset 1+4p^{a+2b}V(\operatorname{mod}p^{a+3b}).$
+Proof.
+Assume $p$ is odd. Write $\gamma_{0}\equiv 1+p^{a}X_{0}(\operatorname{mod}p^{2a})$ for some traceless $X_{0}\in\text{Mat}_{2}(\mathbb{Z})$ that is also primitive mod $p$.
+It is an elementary exercise to check that it is always possible to choose $h_{1},h_{2}\in H$, $h_{1}\equiv 1+p^{b}Y_{1}(\operatorname{mod}p^{2b}),h_{2}\equiv 1+p^{b}Y_{2}(\operatorname{mod}p^{2b})$, $Y_{1},Y_{2}$ traceless and primitive mod $p$, so that $X_{1}=X_{0}Y_{1}-Y_{1}X_{0}$ and $X_{2}=X_{0}Y_{2}-Y_{2}X_{0}$ are primitive and are not colinear mod $p$.
+Write $\gamma_{1}=\gamma_{0}h_{1}\gamma_{0}^{-1}h_{1}^{-1}$ and $\gamma_{2}=\gamma_{0}h_{2}\gamma_{0}^{-1}h_{2}^{-1}$. From Lemma 5.66,
+$\gamma_{1}\equiv 1+\gamma_{0}h_{1}-h_{1}\gamma_{0}\equiv 1+p^{a+b}X_{1}(\operatorname{mod}p^{2a+b})$ (3.40)
+$\gamma_{2}\equiv 1+\gamma_{0}h_{2}-h_{2}\gamma_{0}\equiv 1+p^{a+b}X_{2}(\operatorname{mod}p^{2a+b})$ (3.41)
+Take $h_{3},h_{4}\in H$ and write $h_{3}\equiv 1+p^{b}Y_{3}(\operatorname{mod}p^{2b})$ and $h_{4}\equiv 1+p^{b}Y_{4}(\operatorname{mod}p^{2b})$.
+Applying Lemma 5.66 again,
+$\gamma_{1}h_{3}\gamma_{1}^{-1}h_{3}^{-1}\cdot\gamma_{2}h_{4}\gamma_{2}^{-1}h_{4}^{-1}\equiv 1+p^{a+2b}([X_{1},Y_{3}]+[X_{2},Y_{4}])(\operatorname{mod}p^{a+3b}).$ (3.42)
+Since $X_{1},X_{2}$ are primitive and not colinear mod $p$, it is another elementary exercise to check that
+$\{[X_{1},Y_{3}]+[X_{2},Y_{4}]:Y_{3},Y_{4}\in V\}(\operatorname{mod}p^{b})=V(\operatorname{mod}p^{b}),$ (3.43)
+(Hint: it suffices to assume $X_{1}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right)$ and $X_{2}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right)$, by identifying $V$ with $\mathbb{Z}^{3}$ and use a matrix in $\text{SL}_{3}(\mathbb{Z})$ to change variables.) (3.42) and (3.43) then imply Lemma 3.39.
+The $p=2$ case follows from the same argument. The extra factor 4 is due to the fact that $2V\subset[V,V]\neq V$.
+Lemma 3.44.
+Let $p_{1},p_{2}$ be two primes.
+Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$.
+Then $p_{1}=p_{2}$ and $v_{p_{1}}(\xi-1)\leq m_{1}+n_{2}-m_{2}$.
+Proof.
+We first observe that given $0\leq m_{1}<n_{1},1\leq m_{2}<n_{2}$ and two primes $p_{1},p_{2}$, if there is a nontrivial homomorphism $f$ from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{1}^{n_{2}})$, this will force $p_{1}=p_{2}$.
+To see this, if $m_{1}\neq 0$, then all elements in $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ have orders powers of $p_{1}$, and all elements in $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ have orders powers of $p_{2}$.
+Since $f$ is nontrivial, we have $p_{1}=p_{2}$.
+If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$, which induces a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$.
+Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor through the center $Z$.
+$f^{\prime}$ can not be injective or factor through the center because $\left(\begin{matrix}1&1\\ 0&1\end{matrix}\right)\in\Gamma_{p_{1}}$ has order $p_{1}$, which forces $p_{1}|p_{2}$ if $f$ is injective or factor through $Z$, a contradiction.
+Hence, $f^{\prime}$ as well as $f$ must be trivial.
+Therefore, if $f$ is nontrivial, we must have $p_{1}=p_{2}$.
+Suppose $p_{1}=p_{2}=p$ and $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$.
+Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $v_{p}(\xi)\leq m_{1}+n_{2}-m_{2}$.
+∎
 4. Preliminaries on Random Walks
-Recall that $S$ be a finitely symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$.
-The following lemmas are quantitative statements about non-concentration of self-convolutions of $\chi_{S}$ in proper subvarieties.
-For the discussion in the following in this section, we prefix a primitive linear form
-$L\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right),\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)=X_{1}x_{1}+Y_{1}y_{1}+Z_{1}z_{1}+W_{1}w_{1}+X_{2}x_{2}+Y_{2}y_{2}+Z_{2}z_{2}+W_{2}w_{2},$
-i.e., $\text{gcd}(X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2})=1$.
+Recall that $S$ is a finite symmetric set in ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$.
+In this section we record some quantitative statements on non-concentration of iterated convolutions of $\chi_{S}$ in certain subvarieties.
+Proposition 4.1 and Proposition 4.3 concern linear subvarieties, and Proposition 4.2 is on certain quadratic subvarieties.
 Proposition 4.1.
@@
 Let $\chi_{S}$ be the uniform probability measure supported on $S$.
-There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
-$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c}.$
+There are constants $0<c_{S},c<1$ depending only on $S$ such that for any sufficiently large $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$, and any $n\in\mathbb{Z}$, we have
+$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c},$
+where $L$ is a general primitive linear form on $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, i.e.,
+$L\left(\left(\begin{matrix}x_{0}&x_{1}\\ x_{2}&x_{3}\end{matrix}\right),\left(\begin{matrix}x_{4}&x_{5}\\ x_{6}&x_{7}\end{matrix}\right)\right)=X_{0}x_{0}+X_{1}x_{1}+X_{2}x_{2}+X_{3}x_{3}+X_{4}x_{4}+X_{5}x_{5}+X_{6}x_{6}+X_{7}x_{7},$
+with $\text{gcd}(X_{0},X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{7})=1$.
 Proposition 4.2.
-There are constants $10<c_{S},c<$ 1 depending only on $S$ such that the following holds.
-Let $Q\in\mathbb{Z}_{+}$ large enough and $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
-$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0$
-and either for every $p|Q$, $\pi_{p}(\xi_{1}),\pi_{p}(\xi_{2})\neq 0$, or for every $p|Q$, $\pi_{p}(\eta_{1}),\pi_{p}(\eta_{2})\neq 0$. Then for $l>c_{S}\log Q$ and any $n\in\mathbb{Z}$,
+Let $S$ be given as in Proposition 4.1.
+There are constants $0<c_{S},c<1$ depending only on $S$ such that the following holds.
+Suppose $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
+$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0,$
+Then for any sufficiently large $Q\in\mathbb{Z}_{+}$ with either $\text{gcd}(Q,\text{entries of }\xi_{1}\text{ and }\eta_{1})=1$ or $\text{gcd}(Q,\text{entries of }\xi_{2}\text{ and }\eta_{2})=1$, for any $l>c_{S}\log Q$ and for any $n\in\mathbb{Z}$,
 $\chi_{S}^{(l)}(\{(g_{1},g_{2})\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|\text{Tr}(g_{1}\xi_{1}g_{1}^{-1}\eta_{1})+\text{Tr}(g_{2}\xi_{2}g_{2}^{-1}\eta_{2})\equiv n(\operatorname{mod}{Q})\})<Q^{-c}.$
-In the $\mathrm{ASL}_{2}$ case, we also have the following analogue about non-concentration in proper subvarieties.
+In the $\mathrm{ASL}_{2}$ case, we need
 Proposition 4.3.
-Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense in $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
+Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense.
 Let $\chi_{S}$ be the uniform probability measure supported on $S$.
-There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
+There are constants $0<c_{S},c<1$ depending only on $S$ such that for any sufficiently large $Q\in\mathbb{Z}_{+}$, any $l>{c_{S}}{\log Q}$ and any $n\in\mathbb{Z}$, we have
 $\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c},$
 where $L$ is a primitive linear form on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
-We focus on proving Proposition 4.1; Proposition 4.2 and Proposition 4.4 follow in a similar way (see also Proposition 4.1 in [BG09]). We first show that Proposition 4.1 follow from
+Remark 4.4.
+In Proposition 4.1 and Proposition 4.3, it is important that the implied constant $c$ is independent of $L$ and $n$.
+We focus on proving Proposition 4.1. Proposition 4.2 and Proposition 4.3 follow in a similar way. Our method is similar to the proof of Proposition 4.1 in [BG09].
+We first show that Proposition 4.1 follows from
 Lemma 4.5.
-There are constants $10<c_{S},c_{0}<$ 1 depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for $l\in\mathbb{Z}_{+}$ with $1\ll_{S}l<c_{S}\log Q$, we have
-$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n\pmod{{Q}}\})<e^{-c_{0}l}.$ (4.6)
+There are constants $0<c_{S},c_{0}<1$ depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for any $1\ll_{S}l\leq c_{S}\log Q$, we have
+$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\})<e^{-c_{0}l}.$ (4.6)
 Proof of Proposition 4.1 assuming Lemma 4.5.
@@
 $\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)$
-$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{(g^{\prime})^{-1}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
+$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{g^{\prime}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
 $<$ $e^{-c_{0}l_{0}}<Q^{-c_{0}c_{S}}$
-Proposition 4.1 is thus proved with $c=c_{0}c_{S}$
-Now we focus on proving Lemma 4.5.
+Proposition 4.1 is thus proved with $c=c_{0}c_{S}$. ∎
+Now we focus on proving Lemma 4.5. We begin with
 Lemma 4.7 (Non-concentration at Archimedean place).
@@
 Proof.
-By Theorem A , there exists an absolute constant $10<\lambda<$ 1 which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1).
+By Theorem A, there exists an absolute constant $0<\lambda<1$ which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1).
 So
-$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}$ (4.8)
-for all prime $p$. Then since $|\pi_{p}(\Gamma)|\approx p^{6}$, if $l>\frac{6\log p}{(\log 1/\lambda)}$,
+$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}\|\chi_{S}\|_{2}={\lambda^{l}}{|S|^{-\frac{1}{2}}}$ (4.8)
+for all prime $p$. Since $|\pi_{p}(\Gamma)|\sim p^{6}$, if $l>\frac{6\log p}{\log(1/\lambda)}$,
 $\chi_{S}^{(l)}(g)<\dfrac{2}{|\pi_{p}(\Gamma)|}$ (4.9)
@@
 $<\frac{4}{p},$
-by counting points in the subvariety of codimension $1$ determined by a primitive linear equation.
+by counting points in subvarieties of codimension $1$ determined by primitive linear equations.
 So we are done by picking any prime $p\in[e^{\frac{l}{8}\log\frac{1}{\lambda}},e^{\frac{l}{7}\log\frac{1}{\lambda}}]$, which exists by the Prime Number Theorem when $l$ is sufficiently large.
@@
 Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ without common zeros in $\mathbb{C}^{n}$ with $n\geq 2,\operatorname{deg}\mathcal{P}_{j}\leqslant d,d\geqslant 3$, $h\left(\mathcal{P}_{j}\right)\leqslant h$.
-Here $h(\mathcal{P})=\log\max\limits_{i}\left|a_{i}\right|$ for $\mathcal{P}(x)=a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{n}x^{n}$ is the logarithm height function of polynomials.
+Here $h(\mathcal{P})$ is the logarithmic of the modulus of the coefficient of $\mathcal{P}$ of largest absolute value.
 Then there is an integer $D\in\mathbb{Z}_{+}$, polynomials $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ such that
@@
 Proof of Lemma 4.5.
-Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$. We first deal with the case $n=0$.
-By the primitivity of $L$, for each $p|Q$, at least one of
-$t\in\mathcal{A}:=\{X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2}\}$
+Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$.
+Case 1: $n=0$. By the primitivity of $L$, for each $p|Q$, at least one element in
+$\mathcal{A}:=\{X_{0},X_{1},X_{2},X_{3},X_{4},X_{5},X_{6},X_{7}\}$
 must be invertible $\operatorname{mod}\ p$. For each $t\in\mathcal{A}$, let
-$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}||Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}$
+$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}\|Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}.$
 Since $\prod_{t\in\mathcal{A}}Q_{t}\geq Q$, there exists $t\in\mathcal{A}$ such that
 $Q^{\prime}:=Q_{t}\geq Q^{\frac{1}{8}}.$
-Assume $t=X_{1}$ without loss of generality. Then $Q^{\prime}||Q$ and $(Q^{\prime},X_{1})=1$.
-Now let
-$\|(g_{1},g_{2})\|=\max\{\text{absolute values of coefficients of }g_{1}\text{ and }g_{2}\}.$
-Let $C_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$. Define
+Assume $t=X_{0}$ without loss of generality. Then $Q^{\prime}\|Q$ and $(Q^{\prime},X_{0})=1$.
+[…]
+$\|(g_{1},g_{2})\|=\max\{\|g_{1}\|,\|g_{2}\|\}.$
+Let $N_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$.
+Define
 $\mathcal{G}=\left\{g\in\text{supp }[\chi_{S}^{(l)}]\mid L(g)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\right\}.$
 […]
-$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl}.$
+$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl},$ (4.10)
 for some $c>0$.
-For each $\gamma=\left(\begin{pmatrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{pmatrix},\begin{pmatrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{pmatrix}\right)\in\mathcal{G}$ we introduce a linear polynomial $f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Q}[\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}]$ as follows:
-$f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+{\tilde{Y}}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}$ (4.10)
+For each $\gamma=\left(\begin{pmatrix}\gamma_{0}&\gamma_{1}\\ \gamma_{2}&\gamma_{3}\end{pmatrix},\begin{pmatrix}\gamma_{4}&\gamma_{5}\\ \gamma_{6}&\gamma_{7}\end{pmatrix}\right)\in\mathcal{G}$, we introduce a linear polynomial
+$f_{\gamma}(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})\in\mathbb{Q}[x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7}]$
+as follows:
+$f_{\gamma}(x_{1},\cdots,x_{7})=\gamma_{0}+\sum_{i=1}^{7}\gamma_{i}x_{i}$ (4.11)
 Then we get
-$X_{1}f_{\gamma}\left(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\text{ for all }\gamma\in\mathcal{G}.$ (4.11)
-Here $\bar{X_{1}}$ is the multiplicative inverse of $X_{1}$ mod $Q^{\prime}$.
-Also, by definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely entries of $\gamma$, are bounded by $(2C_{1})^{l}$. Hence,
-$h\left(f_{g}\right)<2l\log C_{1},$
-We claim that there is a common zero $(\tilde{Y},\tilde{Z},\tilde{W},\tilde{A},\tilde{B},\tilde{C},\tilde{D})\in\mathbb{C}^{7}$ to the following system of equations:
-$f_{\gamma}(Y,Z,W,A,B,C,D)=0\quad\text{ for all }\gamma\in\mathcal{G},$ (4.12)
+$f_{\gamma}\left(X_{1}\overline{X_{0}},X_{2}\overline{X_{0}},X_{3}\overline{X_{0}},X_{4}\overline{X_{0}},X_{5}\overline{X_{0}},X_{6}\overline{X_{0}},X_{7}\overline{X_{0}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})$ (4.12)
+for all $\gamma\in\mathcal{G}.$ Here $\overline{X_{0}}$ is the multiplicative inverse of $X_{0}$ mod $Q^{\prime}$.
+Also, by the definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely the entries of $\gamma$, are bounded by $(2N_{1})^{l}$. Hence,
+$h\left(f_{g}\right)<2l\log N_{1},$
+Claim: There is a common zero $(\tilde{x}_{1},\cdots,\tilde{x}_{7})\in\mathbb{C}^{7}-\{\vec{\bf 0}\}$ to the following system of equations:
+$f_{\gamma}({x}_{1},\cdots,{x}_{7})=0,\quad\gamma\in\mathcal{G}.$ (4.13)
+Hence $\mathcal{G}$ is contained in some proper subvariety of $\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})$. Note that the size of $\mathcal{G}$ is bounded by $|S|^{l}$.
+Assume the claim fails to hold.
+We invoke Theorem C with $n=7,d=3,h=2l\log N_{1}$ and $N\leq|S|^{l}$.
+It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{\gamma}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
+$\sum_{\gamma\in\mathcal{G}}f_{\gamma}\varphi_{\gamma}=M$ (4.14)
+with
+$0<\log M,h\left(\varphi_{\gamma}\right)<\mathfrak{X}(7)3^{59}((\log N_{1}+\log(2|S|))l+3\log 3)<N_{2}l$ (4.15)
+where $N_{2}=3^{61}\mathfrak{X}(7)\log(2N_{1}|S|)$.
+Now we take $(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6},y_{7})\in\mathbb{Z}^{7}$ such that
+$(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6},y_{7})\equiv(X_{1}\overline{X_{0}},X_{2}\overline{X_{0}},X_{3}\overline{X_{0}},X_{4}\overline{X_{0}},X_{5}\overline{X_{0}},X_{6}\overline{X_{0}},X_{7}\overline{X_{0}})(\operatorname{mod}Q^{\prime}).$
+It follows that
+$\sum_{\gamma\in\mathcal{G}}f_{\gamma}(y_{1},\cdots,y_{7})\varphi_{\gamma}(y_{1},\cdots,y_{7})-M\equiv 0(\operatorname{mod}Q^{\prime}).$
 […]
-Assume the claim fails to hold.
-We invoke Theorem C with $n=7,d=3,h=2l\log C_{1}$.
-It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{l}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
-$M=\sum_{l=1}^{N}F_{l}\varphi_{l}$ (4.13)
+Since the linear system (4.13) admits a solution and the coefficients of $f_{\gamma}$ are all integral, it must admit a rational solution $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\tilde{x}_{4},\tilde{x}_{5},\tilde{x}_{6},\tilde{x}_{7})$.
+In other words,
+$\gamma_{0}+\sum_{i=1}^{7}\gamma_{i}\tilde{x}_{i}=0,\forall\gamma\in\mathcal{G}$ (4.16)
+If at least one element $\gamma\in\mathcal{G}$ has $\gamma_{0}\neq 0$, then $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\tilde{x}_{4},\tilde{x}_{5},\tilde{x}_{6},\tilde{x}_{7})\neq\vec{\bf{0}}$.
+Rescaling to get rid of the common denominator in (4.16), we obtain $(v_{0},v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7})\in\mathbb{Z}^{8}-\vec{\bf 0}$ such that $\text{gcd}(v_{0},v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7})=1$, and for all $\gamma\in\mathcal{G}$, we have
+$\sum_{i=0}^{7}\gamma_{i}v_{i}=0.$
+Apply Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $v_{1},v_{2},v_{0},v_{3},v_{4},v_{5},v_{6},v_{7}$: For $1\ll_{S}l<c_{S}\log Q$,
+$\chi_{S}^{(l)}(\mathcal{G})\leq\$$\chi_{S}^{(l)}\left(\left\{\gamma\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid\tilde{L}(\gamma)=0\right\}\right)$
+[…]
+where $c_{0}$ is the constant given in Lemma 4.7.
+If all $\gamma\in\mathcal{G}$ have $\gamma_{0}=0$, then $\mathcal{G}$ lies in a linear subvariety, and (5.30) is given by Lemma 4.7. We have thus fully proved the case of Lemma 4.5 when $n=0$.
+Case 2: $n\neq 0$. Write $Q_{1}=(Q,n)$, $n=Q_{1}n_{1}$ and $Q=Q_{1}Q_{2}$. We clearly have $(n_{1},Q_{2})=1$. We analyse the cases $Q_{1}$ small and $Q_{1}$ large separately.
+Case 2.1: $Q_{1}<Q_{2}^{\frac{1}{3N_{3}}}$, where the constant $C_{3}$ is given at (4.19). Consider the following set:
+$\mathcal{G}^{\prime}=\left\{g\in\text{supp }[\chi_{S}^{(l)}]\mid L(g)-Q_{1}n_{1}\equiv 0\ (\operatorname{mod}{Q_{2}})\right\}.$
+For each $\gamma=\left(\left(\begin{matrix}\gamma_{0}&\gamma_{1}\\ \gamma_{2}&\gamma_{3}\end{matrix}\right),\left(\begin{matrix}\gamma_{4}&\gamma_{5}\\ \gamma_{6}&\gamma_{7}\end{matrix}\right)\right)\in\mathcal{G}^{\prime}$, define
+$f_{\gamma}(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})=\sum_{i=0}^{7}\gamma_{i}x_{i}-Q_{1}.$
+If $\{f_{\gamma},\gamma\in\mathcal{G}\}$ has no common zero, then there exist integral polynomials $\varphi_{\gamma}$ and $M\in\mathbb{Z}_{+}$ such that
+$\sum_{\gamma\in\Gamma}f_{\gamma}\varphi_{\gamma}=M.$ (4.17)
 with
-$0<\log M,h\left(\varphi_{l}\right)<\mathfrak{X}(7)3^{59}((\log C_{1}+\log(2|S|))l+3\log 3)<C_{2}l$ (4.14)
-where $C_{2}=3^{61}\mathfrak{X}(7)\log 2C_{1}|S|$.
-Now we take $(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\in\mathbb{Z}^{7}$ such that
-$(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\equiv(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}})(\operatorname{mod}Q^{\prime}).$
-It follows that
-$M\equiv\sum_{l=1}^{N}F_{l}\varphi_{l}\Bigg{|}_{(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})}\equiv 0\ (\operatorname{mod}{Q^{\prime}}),$
-[…]
-$x_{1}+\tilde{Y}y_{1}+\tilde{Z}z_{1}+\tilde{W}w_{1}+\tilde{A}x_{2}+\tilde{B}y_{2}+\tilde{C}z_{2}+\tilde{D}w_{2}=0$
-Rescaling to get rid of the common denominator, we obtain $(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Z}^{8}$ such that $\text{gcd}(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=1$, and for all $\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right)\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)\in\mathcal{G}$, we have
-$\tilde{X}_{1}x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+\tilde{Y}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}=0.$
-We finish the proof of Lemma 4.5 of the case $n=0$ by applying Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}.$ For $1\ll_{S}l<c_{S}\log Q$,
-$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid L(g)\equiv 0(\operatorname{mod}{Q})\right\}\right)$
-[…]
-=\ = $\chi_{S}^{(l)}(\mathcal{G})$
-[…]
-<\ < $e^{-c_{0}l}.$
-where $c_{0}$ is the constant given in Lemma 4.7.
-Now let us deal with the case $n\neq 0$. Let $C_{3}=3^{70}\mathfrak{X}(8)$, where $\mathfrak{X}(8)$ is the constant given by Theorem C, and take $c_{S}=\frac{1}{C_{3}}$. Write $Q_{1}=(Q,n)$.
-Suppose $Q_{1}<Q^{\frac{1}{C_{3}}}$.
-First we can find $n^{\prime}=Q_{1}m\equiv 0(\operatorname{mod}Q)$ with $(m,Q)=1$.
-The equation $L(\gamma)=n(\operatorname{mod}Q)$ is the same as $L(\gamma)=Q_{1}m(\operatorname{mod}Q)$.
-Since $m$ is invertible mod $Q$, we can define
-$f_{\gamma}(X,Y,Z,W,A,B,C,D)=xX+yY+zZ+wW+aA+bB+cD+dD-Q_{1}$
-for $\gamma=\left(\left(\begin{matrix}x&y\\ z&w\end{matrix}\right),\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)$, and proceed as the $n=0$ case.
-The reason in the $n=0$ case we do not define $f_{\gamma}$ in this way is that we need to ensure a solution to $f_{\gamma}=0$ is nonzero.
-But if $n\neq 0$, $f_{\gamma}=0$ never admits a zero solution.
-If $Q_{1}\geq Q^{\frac{1}{C_{3}}}$, then we can reduce this case to the $n=0$ case by considering the equation $L(g)\equiv 0(\operatorname{mod}Q_{1})$ instead of $L(g)\equiv n(\operatorname{mod}Q)$, and letting $c_{S}=\frac{1}{8C_{2}C_{3}}$.
+$0<\log M,h\left(\varphi_{\gamma}\right)<\mathfrak{X}(8)3^{67}((\log N_{1}+\log(2|S|))l+\log Q_{1}+3\log 3)<N_{3}(l+\log Q_{1})$ (4.18)
+where
+$N_{3}=3^{69}\mathfrak{X}(8)\log(2|S|N_{1})$ (4.19)
+Let $(z_{0},\cdots,z_{7})\in\mathbb{Z}^{8}$ such that $(z_{0},\cdots,z_{7})\equiv(\gamma_{0}\overline{n_{1}},\cdots,\gamma_{7}\overline{n_{1}})(\operatorname{mod}Q_{2})$, where $\overline{n_{1}}$ is the multiplicative inverse of $n_{1}(\operatorname{mod}Q_{2})$, then (4.17) implies
+$\sum_{\gamma\in\Gamma}f_{\gamma}(z_{0},\cdots,z_{8})\varphi_{\gamma}(z_{0},\cdots,z_{8})\equiv M(\operatorname{mod}Q_{2}),$
+so that
+$M\geq Q_{2},$ (4.20)
+which contradicts (4.18) if we take $c_{S}\leq\frac{1}{3N_{3}}\log Q$.
+Hence $\{f_{\gamma},\gamma\in\mathcal{G}\}$ have a common zero $(v_{1},\cdots,v_{8})\in\mathbb{Z}^{8}-\{\vec{\bf 0}\}$.
+Applying Lemma 4.7 then yields
+$\chi_{S}^{(l)}(\mathcal{G}^{\prime})<e^{-cl},$ (4.21)
+for some $c>0$.
+Case 2.2: $Q_{1}\geq Q_{2}^{\frac{1}{N_{3}}}$.
+In this case the equation $L(\gamma)-n\equiv 0(\operatorname{mod}Q)$ implies $L(\gamma)\equiv 0(\operatorname{mod}Q_{1})$.
+Thus we reduce this case to Case 1.
 ∎
@@
 Let $\mathbb{P}_{0}$ be the projection from $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ to $\text{SL}_{2}(\mathbb{Z})$, and let $\mathbb{P}_{1}$ ($\mathbb{P}_{2}$, resp.) be the projection map from $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ to its first factor (second factor, resp.).
-The goal of this section is to prove, the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
+The goal of this section is to prove the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
 Proposition 5.1.
-Let $S,\epsilon$ be as given in Proposition 2.2.
-There are constants $c>0$ depending only on the generating set $S$ and $\epsilon$, and a constant $\rho=\rho(\delta)>0$, $C=C(\delta)\in\mathbb{Z}_{+}$, with $\rho=O_{S}(\delta)$, such that for any $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}>\frac{12}{\delta}$, if the conclusion of Proposition 2.2 fails, i.e.,
+Let $S$ be a finite symmetric set in $\Lambda$ which generates a Zariski dense subgroup.
+There exist positive functions $\rho=\rho(\delta)=O_{S}(\delta),C=C(\delta)\in\mathbb{Z}_{+}$ defined for any sufficiently small $\delta>0$, such that the following holds: For any sufficiently large $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}\geq\frac{3}{\delta}$, and for any $l\in\mathbb{Z}_{+}$ and $A\subset\Lambda$ that satisfy
+$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q$
+and
 $|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta},$ (5.2)
-then we have
-$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &\mathbb{(}P_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
-where $q_{*}||q$ and $q_{*}\geq q^{c}$.
-We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$. The proofs for the other two cases are identical. Our proof mostly follows the steps of [BG09].
-5.1. A large set of commuting elements
-We use Helfgott’s argument [Hel08] to create a large set of commuting elements under an appropriate modulus.
-Let $c_{1},c_{2}$ be the implied constants for $c$ from Proposition 4.1 and Proposition 4.2 respectively. Let $A_{0}=A\cdot A\cap\Lambda(q_{0})$, where $q_{0}=\prod_{p|q_{l}}p$. Then
-$\chi_{S}^{2l}(A_{0})>q^{-2\delta}q_{0}^{-12}>q^{-3\delta},$ (5.4)
-where we have used the exponent for each prime divisor of $q$ exceeds $\frac{12}{\delta}$.
-Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$ and let
-$A_{1}=\mathbb{P}_{1}(A_{0}).$
+we have
+$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &(\mathbb{P}_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
+where $q_{*}\|q$ and $q_{*}\geq q^{c}$ for some $c>0$ depending only on $S$.
+We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$.
+The proofs for the remaining two cases are virtually identical.
+Our proof follows the framework in [BG09], which are outlined below.
+These statements are understood to hold under the reduction of appropriate moduli.
+  1. (a)
+    Use non-concentration of random walks (Proposition 4.1) to find a large set of matrices with distinct traces from $\mathbb{P}_{1}(A^{2})$ (Lemma 5.7).
+  2. (b)
+    Assuming (5.2) holds, use the matrices obtained from (a) to create a large set of commutative matrices (Lemma 5.18).
+  3. (c)
+    Diagonalize the matrices from (b) over a quadratic field to generate a large set of diagonal matrices that satisfy multiple density and proximity-to-identity constraints (Lemma 5.32).
+  4. (d)
+    Apply the sum-product theorem (Theorem B) to create a line in $\mathfrak{sl}_{2}$ (Lemma 5.50).
+  5. (e)
+    Use random walks (Proposition 4.2) to move the line from (d) to different directions by conjugation, thus to create a thick segment in $\mathfrak{sl}_{2}$ (Lemma 5.92).
+  6. (f)
+    Iteratively take commutators of the segment obtained from (e) to finish the proof of Proposition 5.1.
+Remark 5.4.
+The method for finding commutative matrices can be traced back to [Hel08].
+The presentation in this section involves several constants. For readers’ convenience we record them here.
+Let $c_{1},c_{2}$ be the implied constants for $c$ in Propositions 4.1, 4.2. We take $c_{1},c_{2}<1$.
+Let $C_{1},C_{2},C_{3},\varepsilon=\varepsilon_{0}$ be the implied constants in Theorem B for $\alpha=\frac{c_{1}}{6}$.
+Let $\rho_{1},\rho_{2}$ be two small quantities linearly depending on the (only) variable $\delta$:
+$\rho_{1}=L_{1}\delta,\hskip 28.45274pt\rho_{2}=L_{2}\rho_{1}=L_{1}L_{2}\delta.$ (5.5)
+Here $L_{1}$ and $L_{2}$ are two big constants up to several adjustments in the proof process, and we assume at each adjustment, they satisfy not only the current requirement but also all preceding ones.
+While it is important $L_{1}$ and $L_{2}$ only depend on $S$, we take $L_{1}$ and $L_{2}$ so large that any positive linear combination of $\delta$ and $\rho_{1}$ appearing in our presentation is dominated by the $\rho_{1}$ term, and any positive linear combination of $\rho_{1}$ and $\rho_{2}$ is dominated by the $\rho_{2}$ term.
+So $\delta$ should be considered significantly smaller than $\rho_{1}$, which in turn is significantly smaller than $\rho_{2}$.
+The main parameter at (5.5) is $\rho_{2}$; the parameter $\rho_{1}$ plays a correction role.
+Unless otherwise specified, the Landau notations $\Theta,\Omega,O$ in this section describe functions of $\delta$ and the implied constants depend only on $S$.
+We also need to introduce two small quantities $w_{1},w_{2}=\Theta(\rho_{2})$ given in Lemma 5.32. So by our convention above, $w_{1}$, $w_{2}$ is much larger than $\rho_{1}$.
+5.1. A large set of traces
+Let
+$A_{0}=(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))\cap\Gamma(\tilde{q})\subset\mathbb{P}_{1}(A^{2})\subset\text{SL}_{2}(\mathbb{Z}),$
+where $\tilde{q}=\prod_{p|q}p$ is the square-free part of $q$.
+Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$.
+For each $x\in\Gamma_{\tilde{q}}$, let
+[…]
+Then,
+$\nu^{(2l)}(A_{0})$ $=\sum_{x\in\Gamma_{\tilde{q}}}\nu^{(l)}(B_{x})\nu^{(l)}(B_{x^{-1}})\stackrel{{\scriptstyle\text{symmetry of $A$}}}{{=}}\sum_{x\in\Gamma_{\tilde{q}}}\nu^{(l)}(B_{x})^{2}$
+$\stackrel{{\scriptstyle\text{Cauchy-Schwarz}}}{{\geq}}\frac{\nu^{(l)}(\mathbb{P}_{1}(A))^{2}}{|\Gamma_{\tilde{q}}|}>q^{-2\delta}(\tilde{q})^{-3}\geq q^{-3\delta},$ (5.6)
+where in the last inequality we have used that the exponent for each prime divisor of $q$ exceeds $\frac{3}{\delta}$.
+The goal of this section is to prove
+Lemma 5.7.
+There is a set $A_{1}\subset A_{0}$, an exact divisor $q_{1}\|q$ with $q_{1}>q^{\frac{1}{2}}$, such that
+$|A_{1}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{2}},$ (5.8)
+$\forall\gamma_{1},\gamma_{2}\in A_{1},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\not\equiv\text{tr}(\gamma_{2})(\operatorname{mod}q_{1}^{\{\rho_{2}\}})$ (5.9)
+$\forall\gamma=\left(\begin{matrix}\gamma_{11}&\gamma_{12}\\ \gamma_{21}&\gamma_{22}\end{matrix}\right)\in A_{1},\text{tr}(\gamma)\not\equiv\pm 2,\gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4n\rho_{1}]}),\forall p^{n}\|q_{1},$ (5.10)
+where $c_{1}$ is the implied constant $c$ from Proposition 4.1.
+Remark 5.11.
+Lemma 5.7 gives us a large set of matrices with distinct traces not close to $\pm 2$ at all non-Archimedean places $p$ where $v_{p}(q_{1})>0$ .
+We also need conditions (5.9) and (5.10) because they ensure that we can diagonalize any element from $A_{1}$ by a matrix with controlled denominator.
+Proof of Lemma 5.7.
 For each $Q|q$, let
-$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)^{2}-4\equiv 0(\operatorname{mod}Q))\},$
+$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)\equiv\pm 2\equiv 0(\operatorname{mod}Q))\},$
 $\mathcal{E}_{2}(Q)=\left\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\gamma_{21}\equiv 0(\operatorname{mod}Q))\right\}.$
-Then Proposition 4.1 gives
-$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}}.$
-Take $\rho_{1}>0$, and let
-$A_{2}=A_{1}-\cup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\{\rho_{1}\}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)),$
-We should think of $\rho_{1}$ as a relative level depending only on $S$, so as $\rho_{2}$ in the following; the exact value of $\rho_{1}$ is determined at (5.58).
+Proposition 4.1 gives
+$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}},$
+for sufficiently large $Q$. Let
+$A_{1}^{\prime}=A_{0}-\bigcup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\rho_{1}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)).$
 Since the number of divisors of $q$ is bounded by $q^{0+}$, it follows that
-$\nu^{(2l)}(A_{2})>q^{-3\delta-},$
+$\nu^{(2l)}(A_{1}^{\prime})>q^{-3\delta-},$
 if we let
-$\delta<\frac{\rho_{1}c_{1}}{4}.$ (5.5)
-A simple argument by contradiction shows that for each $\gamma\in A_{2}$, there exists $q_{\gamma}||q$, $q_{\gamma}>q^{\frac{1}{2}}$, such that for any $p^{n}||q_{\gamma}$,
-$\text{tr}(\gamma)^{2}-4\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.6)
-$\gamma_{21}\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.7)
-Since again the number of divisors of $q$ is bounded by $q^{0+}$, there exists $q_{1}||q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{3}\subset A_{2}$, with
-$\nu^{(2l)}(A_{3})>q^{-3\delta-},$ (5.8)
+$L_{1}>\frac{1}{c_{1}}.$ (5.12)
+For each $\gamma\in A_{1}^{\prime}$, let
+$q_{\gamma}^{{}^{\prime}}=\prod_{\begin{subarray}{c}p^{n}\|q\\ \text{tr}(\gamma)\not\equiv\pm 2(\operatorname{mod}p^{[4n\rho_{1}]})\end{subarray}}p^{n},$
+and
+$q_{\gamma}^{{}^{\prime\prime}}=\prod_{\begin{subarray}{c}p^{n}\|q\\ \gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4n\rho_{1}]})\end{subarray}}p^{n}.$
+Since $\text{gcd}(\text{tr}(\gamma)\pm 2,q)<q^{\rho_{1}}$ and $\text{gcd}(\gamma_{21},q)<q^{\rho_{1}}$, we must have $q_{\gamma}^{\prime},q_{\gamma}^{\prime\prime}>q^{\frac{3}{4}}$.
+Let $q_{\gamma}=\text{gcd}(q_{\gamma}^{\prime},q_{\gamma}^{\prime\prime})$. Then $q_{\gamma}>q^{\frac{1}{2}}$, and for any $p^{n}\|q_{\gamma}$,
+$\text{tr}(\gamma)\not\equiv\pm 2(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.13)
+$\gamma_{21}\not\equiv 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.14)
+Since again the number of divisors of $q$ (which bounds the number of choices for $q_{\gamma}$) is bounded by $q^{0+}$, there exists $q_{1}\|q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{1}^{\prime\prime}\subset A_{1}^{\prime}$, with
+$\nu^{(2l)}(A_{1}^{\prime\prime})>q^{-3\delta-}>q_{1}^{-6\delta-},$ (5.15)
+such that for any $\gamma\in A_{1}^{\prime\prime}$, any $p^{n}\|q_{1}$, (5.13) and (5.14) are satisfied.
+Let
+$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}q_{1}^{\{\rho_{2}\}})\}.$
+By Proposition 4.1, for any $n\in\mathbb{Z}$,
+$\nu^{(2l)}(W(n))<q_{1}^{-\rho_{2}c_{1}}.$ (5.16)
+If we take
+$L_{2}>\frac{12}{c_{1}},$ (5.17)
+then (5.15) implies
+$\nu^{(2l)}(A_{1}^{\prime\prime})>q_{1}^{-\frac{\rho_{2}c_{1}}{2}},$
+which together with (5.16) implies the existence a set $A_{1}\subset A_{1}^{\prime\prime}$ satisfying all requirements in Lemma 5.7.
+5.2. A large set of commutative matrices
+Our goal in this section is to prove
+Lemma 5.18.
+Let $A_{1}$ be as given in Lemma 5.7. There is an element $\alpha_{0}\in A_{1}$ and a set $A_{2}\subset\mathbb{P}_{1}(A)^{2}$ such that
+  1. (1)
+    $\forall\gamma_{1},\gamma_{2}\in A_{2}$, $\gamma_{1}\neq\gamma_{2}\Rightarrow\gamma_{1}\neq\gamma_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})$,
+  2. (2)
+    $|A_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{3}},$ (5.19)
+  3. (3)
+    $A_{2}$ commutes with $\alpha_{0}$ mod $q_{1}^{\{\rho_{2}\}}$.
+Proof.
+Denote
+$Q_{0}=q_{1}^{\{\rho_{2}\}}.$
+First, it follows from Proposition 4.1 that for any $\gamma_{0}\in\Gamma_{q}$,
+$\nu^{(2l)}\{\gamma\in\Gamma:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{0})\}<Q_{0}^{-c_{1}}.$ (5.20)
+Therefore,
+$|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|>Q_{0}^{c_{1}}q^{-3\delta}>Q_{0}^{\frac{c_{1}}{2}},$ (5.21)
+recalling (5.17).
+Next, we use our assumption (5.2) to get a control over the growth of $\pi_{Q_{0}}(\mathbb{P}_{1}(A))$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
+$|\pi_{q}(A^{l})|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.22)
+$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.23)
+This implies
+$|\pi_{Q_{0}}(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{36}}.$ (5.24)
+To see this, suppose not, then
+$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A)\cdot\mathbb{P}_{1}(A))|\cdot\max_{\gamma_{0}\in\Gamma_{Q_{0}}}|(\gamma_{1},\gamma_{2})\in A:\gamma_{1}\equiv\gamma_{0}(\operatorname{mod}Q_{0})|$
+$\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{36}}\cdot\max_{\gamma_{0}\in\Gamma_{Q_{0}}}|(\gamma_{1},\gamma_{2})\in A:\gamma_{1}\equiv\gamma_{0}(\operatorname{mod}Q_{0})|$
+$\geq$ $|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{\frac{c_{1}}{36}}\cdot|\pi_{q}(A)|$
 […]
-Let
-$Q_{1}=q_{1}^{\{\rho_{2}\}},$
-where we will take $\rho_{2}\gg\rho_{1}$.
-Let
-$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}Q_{1})\}.$
-By Proposition 4.1, for any $n\in\mathbb{Z}$,
-$\nu^{2l}(W(n))<Q_{1}^{-c_{1}}.$ (5.9)
-If we take
-${\delta}<{\frac{\rho_{2}c_{1}}{12}},$ (5.10)
-then $\nu^{(2l)}(A_{3})>Q_{1}^{-\frac{c_{1}}{2}}$, and together with (5.9), it implies that there exists a set $A_{4}\subset A_{3}$, such that
-$|A_{4}(\operatorname{mod}Q_{1})|>Q_{1}^{c_{1}/2}$ (5.11)
-$\forall\gamma_{1},\gamma_{2}\in A_{4},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\neq\text{tr}(\gamma_{2})(\text{mod }Q_{1})$ (5.12)
-Next we need to use our assumption (5.2) to get a control over the growth of $\pi_{Q_{1}}(A_{1})$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
-$|\pi_{q}(\prod_{l}A)|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.13)
-$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.14)
-It follows from Proposition 4.1 that for any $\gamma_{0}\in\Lambda_{q}$,
-$\nu^{(2l)}\{\gamma\in\Lambda:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{1})\}<Q_{1}^{-c_{1}}.$ (5.15)
-From (5.15) and (5.10), we conclude that
-$|\pi_{Q_{1}}(A_{1})|>Q_{1}^{c_{1}}q^{-3\delta}>Q_{1}^{\frac{c_{1}}{2}}.$
-Next we show small expansion for $A$ implies small expansion for $A_{1}$:
-$|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$ (5.16)
-To see this, suppose not, then
-$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
-$\geq$ $|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
-$\geq$ $|\pi_{Q_{1}}(A_{1})|^{\frac{4\delta}{\rho_{2}c_{1}}}\cdot|\pi_{q}(A)|$
-$>$ $|\pi_{q}(A)|^{1+\delta}$ (5.17)
-which contradicts (5.2).
-For each $\alpha\in A_{4}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in A_{1}\}$.
-Since distinct elements in $A_{4}$ have distinct eigenvalues mod $Q_{1}$, the sets $\{\pi_{Q_{1}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint.
+which contradicts (5.22) for $l=4$, if we take $L_{2}$ sufficiently large.
+For each $\alpha\in A_{1}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in\mathbb{P}_{1}(A)\}$.
+Since distinct elements in $A_{1}$ have distinct eigenvalues (mod $Q_{0}$), the sets $\{\pi_{Q_{0}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint.
 Therefore,
-$\sum_{\alpha\in A_{4}}|\pi_{Q_{1}}(C_{\alpha})|=|\cup_{\alpha\in A_{4}}\pi_{Q_{1}}(C_{\alpha})|<|\pi_{Q_{1}}(A_{1}^{3})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$
-Thus, for some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{4}$,
-$|\pi_{Q_{1}}(C_{\alpha_{0}})|\leq|\pi_{Q_{1}}(A)|^{1+\frac{4\delta}{\rho_{2}c_{1}}}Q_{1}^{-c_{1}/2}<|\pi_{Q_{1}}(A)|Q_{1}^{-\frac{c_{1}}{4}}.$ (5.18)
-provided that
-$\delta<\frac{\rho_{2}c_{1}^{2}}{48}.$ (5.19)
-To proceed, we first diagonalise $\alpha_{0}$ in an extended number field, i.e., let
+$\sum_{\alpha\in A_{1}}|\pi_{Q_{0}}(C_{\alpha})|=|\cup_{\alpha\in A_{1}}\pi_{Q_{0}}(C_{\alpha})|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A)^{4})|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{18}}.$
+Thus, recalling (5.8), there exists some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{1}$,
+$|\pi_{Q_{0}}(C_{\alpha_{0}})|\leq|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{18}}Q_{0}^{-c_{1}/2}<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|Q_{0}^{-\frac{c_{1}}{3}}.$ (5.26)
+It follows from (5.26) that there exists $x_{0}\in\mathbb{P}_{1}(A)$ and
+$A_{2}^{\prime}:=\{x\in\mathbb{P}_{1}(A):x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{0})\},$ (5.27)
+such that
+$|\pi_{Q_{0}}(A_{2}^{\prime})|>Q_{0}^{\frac{c_{1}}{3}}.$ (5.28)
+Therefore, $A_{2}^{\prime\prime}=A_{2}^{\prime}\cdot x_{0}^{-1}$ commutes with $\alpha_{0}$ mod $Q_{0}$ and
+$\displaystyle|\pi_{Q_{0}}(A_{2}^{\prime\prime})|>Q_{0}^{\frac{c_{1}}{3}},$ (5.29)
+Choose $A_{2}\subset A_{2}^{\prime\prime}$ to be a set of representatives for $\pi_{Q_{0}}(A_{2}^{\prime\prime})$, which gives Lemma 5.18. ∎
+5.3. Applying sum-product
+To proceed, we first diagonalize $\alpha_{0}$ in an extended number field. Let
 $\lambda_{1}=\frac{\text{tr}(\alpha_{0})+\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2},\hskip 14.22636pt\lambda_{2}=\frac{\text{tr}(\alpha_{0})-\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2}$
 be the two eigenvalues of $\alpha_{0}$.
-Let $K=\mathbb{Q}[\lambda_{1}]$, and so $\mathcal{O}=\mathcal{O}_{K}=\mathbb{Z}[\lambda_{1}]$.
+Let $K=\mathbb{Q}[\lambda_{1}]$, and $\mathcal{O}=\mathcal{O}_{K}$ be the ring of integers of $K$.
 Let $M=\left(\begin{matrix}\lambda_{1}-\alpha_{22}&\lambda_{2}-\alpha_{22}\\ \alpha_{21}&\alpha_{21}\end{matrix}\right)$.
 Then
-$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$
+$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$ (5.30)
 Since $\alpha_{0}$ satisfies the condition (5.10) for $\gamma$,
-$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]})$ (5.20)
-for any $p^{n}||q_{1}$.
-From (5.18), there exists $x_{0}\in A_{4}$ and
-$A_{5}:=\{x\in A:x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{1})\},$ (5.21)
-such that
-$|\pi_{Q_{1}}(A_{5})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.22)
-Therefore, $A_{5}^{-1}\cdot x_{0}$ commute with $\alpha_{0}(\operatorname{mod}Q_{1})$ and
-$|\pi_{Q_{1}}(A_{5}^{-1}\cdot x_{0})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.23)
-Let $C_{1},C_{2},C_{3}$ be the implied constants of Theorem B for $\delta=\frac{c_{1}}{17}$.
-The condition (5.23) implies that there exists $\upsilon\in[\frac{c_{1}}{24},\frac{1}{2}]$, such that
-$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{5})|>|\pi_{Q_{1}^{\{\upsilon\}}}(A_{5})|Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.24)
-If we let $A_{6}=\left((A_{5}^{-1}\cdot x_{0})\cdot(A_{5}^{-1}\cdot x_{0})^{-1}\right)\cap\Gamma(Q_{1}^{\{\upsilon\}})=A_{5}^{-1}A_{5}\cap\Gamma(Q_{1}^{\{\upsilon\}})$, then
-$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{6})|>Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.25)
+$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]}),\hskip 28.45274pt\forall p^{n}\|q_{1}.$ (5.31)
+Let $C_{1},C_{2},C_{3},\varepsilon=\varepsilon_{0}$ be the implied constants of Theorem B for $\alpha=\frac{c_{1}}{6}$.
+The first goal of this section is to prove
+Lemma 5.32.
+There are scales $0<w_{1},w_{2}<1$ with $w_{1},w_{2}=\Theta(\rho_{2})$, $w_{2}\leq w_{1}$, and a set $H\subset M^{-1}\cdot A_{2}^{2}\cdot M$, and an ideal $\mathcal{H}\supset\left(q_{1}^{\{\frac{w_{2}}{C_{1}}\}}\right)$ in $\mathcal{O}$, such that
+  1. (i)
+    $H\equiv 1(\operatorname{mod}q_{1}^{\{w_{1}\}})$,
+  2. (ii)
+    $H$ is diagonal mod $q_{1}^{\{w_{1}+w_{2}\}}$,
+  3. (iii)
+    (lower bound for dimension) Elements in $H$ are distinct mod $q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}$ and
+    $\left|\pi_{q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}}(H)\right|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{5}-}.$
+  4. (iv)
+    (common divisor) For each $h\in H$, writing $h\equiv\left(\begin{matrix}a_{h}&0\\ 0&a_{h}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}})$ for some $a_{h}\in\mathcal{O}$, then $\langle{a_{h}-1},q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}\rangle=q_{1}^{\{w_{1}\}}\mathcal{H}$.
+Remark 5.33.
+The reason that we choose the scales $w_{1},w_{2}$ in the above way is that subsequently we will construct a set $\mathscr{H}\subset\mathcal{O}$ from $H$ and apply Theorem B to $A=\mathscr{H}$, $\mathfrak{a}=(q_{1}^{\{\frac{w_{2}}{C_{1}}\}})$ and density $\alpha=\frac{c_{1}}{6}$.
+The conclusion of Theorem B then ensures that the obtained ideal $\mathfrak{a}^{\prime}\supset(q_{1}^{\{w_{2}\}})$, so that we can use Property (ii) from Lemma 5.32 in the computation of some matrix products for the proof of Lemma 5.53.
+Proof of Lemma 5.32.
+The denominator of $M^{-1}$ is controlled $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from which together with (5.10), we have
+$p^{[8\rho_{1}n]}\nmid\text{Det}(M),\hskip 28.45274pt\forall p^{n}\|q_{1}.$ (5.34)
+Therefore, from Lemma 5.18, since $A_{2}$ commutes with $\alpha_{0}$ mod $q_{1}^{\{\rho_{2}\}}$, we have
+$H_{0}:=M^{-1}A_{2}M$
+commutes with $M^{-1}\alpha_{0}M$ mod $q_{1}^{\{\rho_{2}-8\rho_{1}\}}$. Moreover, because of the condition (5.31), we have, for any $\gamma\in A_{2}$, we can find $a_{\gamma}\in\mathcal{O}$ such that
+$M^{-1}\gamma M\equiv\left(\begin{matrix}a_{\gamma}&0\\ 0&a_{\gamma}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}}).$ (5.35)
+We first give an estimate for $|H_{0}|$. The density condition (5.19) implies that
+$|A_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}-20\rho_{1}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{4}},$ (5.36)
+by taking $L_{2}$ sufficiently large.
+If $\gamma_{1}\not\equiv\gamma_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}-20\rho_{1}\}})$, then $M^{-1}\gamma_{1}M\not\equiv M^{-1}\gamma_{2}M(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})$.
+(5.36) thus implies:
+$|H_{0}(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{4}}.$ (5.37)
+We should think of the $\rho_{1}$ terms in (5.35), (5.36), (5.37) negligible compared to the $\rho_{2}$ terms, and this applies to any similar expressions.
+The density condition (5.37) implies that at east one of the following two events occurs: Event 1: There is a scale $w_{1}\in[\frac{c_{1}}{80}\rho_{2},\frac{1}{2}(\rho_{2}-12\rho_{1})]$, and $h_{0}\in H_{0}$ such that the set
+$H_{1}:\{h\in H_{0}:h\equiv h_{0}(\operatorname{mod}q_{1}^{\{w_{1}\}})\}$ (5.38)
+satisfies
+$|H_{1}(\operatorname{mod}q_{1}^{\{w_{1}(1+C_{1}^{-1})\}})|>q_{1}^{\frac{w_{1}}{C_{1}}\cdot\frac{c_{1}}{5}},$ (5.39)
+In this case, we set $w_{2}=w_{1}$.
+Event 2: There is a scale $w_{1}\in[\frac{1}{2}(\rho_{2}-12\rho_{1}),(1-\frac{c_{1}}{80})\rho_{2}]$ and $h_{0}\in H_{0}$ such that the set
+$H_{1}:\{h\in H_{0}:h\equiv h_{0}(\operatorname{mod}q_{1}^{\{w_{1}\}})\}$ (5.40)
+satisfies
+$|H_{1}(\operatorname{mod}q_{1}^{\{w_{1}+\frac{\rho_{2}-12\rho_{1}-w_{1}}{C_{1}}\}})|>q_{1}^{\frac{\rho_{2}-12\rho_{1}-w_{1}}{C_{1}}\cdot\frac{c_{1}}{5}},$ (5.41)
+In this case, we set $w_{2}=\rho_{2}-12\rho_{1}-w_{1}$.
+Indeed, suppose no such $w_{1}$ and $h_{0}$ exist.
+Set $t_{0}=\frac{c_{1}}{80}\rho_{2},$ and $t_{s}=t_{0}(1+C_{1}^{-1})^{s}$ for $1\leq s\leq s_{1}$, where $s_{1}$ is the largest integer such that $t_{s_{1}}\leq\frac{1}{2}(\rho_{2}-12\rho_{1})$.
+We continue to iteratively define $t_{s_{1}+1}=t_{s_{1}}+\frac{\rho_{2}-12\rho_{1}-t_{s_{1}}}{C_{1}}$, $t_{s_{1}+2}=t_{s_{1}+1}+\frac{\rho_{2}-12\rho_{1}-t_{s_{1}+1}}{C_{1}}\cdots$, until we reach a first integer $s_{2}$ such that $t_{s_{2}}>(1-\frac{c_{1}}{80})\rho_{2}$.
+Then
+$|H_{0}(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})|$ $\leq|\pi_{q_{1}^{\{t_{s_{2}}\}}}(H_{0})|\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
+$\leq|\pi_{q_{1}^{\{t_{s_{2}-1}\}}}(H_{0})|\cdot q_{1}^{({t_{s_{2}}-t_{s_{2}-1}})\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
+$\leq|\pi_{q_{1}^{\{t_{s_{2}-2}\}}}(H_{0})|\cdot q_{1}^{({t_{s_{2}-1}-t_{s_{2}-2}})\cdot\frac{c_{1}}{5}}\cdot q_{1}^{({t_{s_{2}}-t_{s_{2}-1})}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
+$\cdots$
+$\leq|\pi_{q_{1}^{\{t_{0}\}}}(H_{0})|\cdot q_{1}^{{t_{s_{2}}}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
+$\leq q_{1}^{\frac{\rho_{2}c_{1}}{40}}\cdot q_{1}^{{t_{s_{2}}}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}\leq q_{1}^{\frac{\rho_{2}c_{1}}{4}},$
+contradicting (5.37). So such $h_{0}$ and $w_{1}$ have to exist.
+Let $H_{2}\subset h_{0}^{-1}H_{1}$ be a set of representatives of $h_{0}^{-1}H_{1}(\operatorname{mod}q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}})$.
+Then $H_{2}$ satisfies the conditions (i), (ii), (iii) for Lemma 5.32.
+For each $h\in h_{0}^{-1}H_{2}$, take some $a_{h}\in\mathcal{O}$ so that
+$h\equiv\left(\begin{matrix}a_{h}&0\\ 0&a_{h}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}}).$
+From (5.40), for each $h\in H_{2}$, we have $a_{h}\equiv 1(\operatorname{mod}q_{1}^{\{w_{1}\}})$.
+Let $\mathcal{H}_{h}$ be an ideal of $\mathcal{O}$ such that $q_{1}^{\{w_{1}\}}\mathcal{H}_{h}=\langle a_{h}-1,q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}\rangle$.
+Since the number of divisors of $\left(q_{1}^{\left\{\frac{w_{2}}{C_{1}}\right\}}\right)$ is $O(q_{1}^{0+})$, there is $\mathcal{H}$ such that
+$H:=\{h\in H_{2}:\mathcal{H}_{h}=\mathcal{H}\}$
+satisfies
+$|H(\operatorname{mod}q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}})|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{5}-}.$ (5.42)
+Then $H$ satisfies all requirements in Lemma 5.32.
+For each $h\in H$, there is $\lambda_{h}\in\mathcal{O}$ such that
+$a_{h}-a_{h}^{-1}\equiv q_{1}^{\{w_{1}\}}\lambda_{h}(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}}).$ (5.43)
 Let
-$Q_{2}:=q_{1}^{\{\rho_{2}\upsilon-9\rho_{1}\}},$
-where we will take $\rho_{2}\upsilon\gg 9\rho_{1}$. Since $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from (5.7), (5.6), we have
-$p^{[6\rho_{1}n]}\nmid\text{Det}(M),\forall p^{n}||q_{1}.$ (5.26)
-Combining (5.26) with (5.20), we conclude that the matrix $M$ conjugate every matrix $\sigma$ in $A_{6}$ to a matrix in $\text{SL}_{2}(\mathcal{O})$ which is congruent to identity mod $Q_{2}$ and diagonal mod $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$.
-In other words, there exists $a_{\sigma}\in\mathcal{O}_{K}$ and $a_{\sigma}\equiv 1(\operatorname{mod}Q_{2})$ such that
-$M^{-1}\sigma M\equiv\left(\begin{matrix}a_{\sigma}&0\\ 0&a_{\sigma}^{-1}\end{matrix}\right)(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}).$ (5.27)
-It also follows from (5.25) that
-$\Big{|}\pi_{Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}(A_{6})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}}.$ (5.28)
-if
-$\rho_{1}<\frac{\upsilon c_{1}}{2^{5}3^{3}}\rho_{2},$ (5.29)
-which is guaranteed if
-$\rho_{1}<\frac{c_{1}^{2}}{2^{8}3^{4}}\rho_{2}.$ (5.30)
-For each $\gamma\in A_{6}$, let ${q}_{\gamma}$ be the largest divisor of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ such that $\gamma\equiv 1(\operatorname{mod}q_{\gamma})$.
-Obviously, $Q_{2}|q_{\gamma}$.
-Then we can write
-$A_{6}=\sqcup_{Q_{2}|q^{\prime}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}B_{q^{\prime}},$
-where $B_{q^{\prime}}=\{\gamma\in A_{6}:q_{\gamma}=q^{\prime}\}$.
-Since the number of divisors of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ is bounded by $q^{0+}$, there exists $\tilde{q}$ such that $Q_{2}|\tilde{q}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ and
-$\Big{|}B_{\tilde{q}}(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.31)
-5.2. Applying sum-product
-Let $H=\{h_{\sigma}=(a_{\sigma}-a_{\sigma}^{-1})/\tilde{q}:\sigma\in B_{\tilde{q}}\}\subset\mathcal{O}$, where $a_{\sigma}$ is as given in (5.27).
-Since the map $a_{\sigma}\mapsto a_{\sigma}-a_{\sigma}^{-1}$ is bounded to one, we have
-$\Big{|}H(\operatorname{mod}{Q_{2}^{\{\frac{1}{C_{1}}\}}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.32)
-Recall $\mathcal{O}$ is a degree 2 extension of $\mathbb{Z}$.
-We apply Theorem B to the set $H$ with $\mathfrak{a}=(Q_{2}^{\{\frac{1}{C_{1}}\}})$ and $\delta=\frac{c_{1}}{17}$.
-Theorem B then produces some $\xi\in\mathcal{O}$, $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$, $c_{3}=c_{3}(c_{1})>0$ such that
-$\sum_{C_{3}}H^{C_{2}}\supset\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime}),$ (5.33)
-and
-$|\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime})|>N(\mathfrak{a})^{c_{3}}>q_{1}^{\frac{\rho_{2}\upsilon c_{3}}{C_{1}}}>q_{1}^{\frac{\rho_{2}c_{1}c_{3}}{24C_{1}}}.$ (5.34)
-We pass (5.34) to a local property.
-From (5.34), it follows that there exists $q_{2}||q_{1}$, $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}$, such that for any $p^{n}||q_{2}$, letting $\overline{\mathfrak{a}_{p}}=\langle\mathfrak{a}^{\prime},p^{n}\rangle$, the ideal of $\mathcal{O}$ generated by $\mathfrak{a}^{\prime}$ and $p^{n}$, we have
-$|\mathbb{Z}\xi(\operatorname{mod}\overline{\mathfrak{a}_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{48C_{1}}}.$
-Since there are at most two prime ideals in $\mathcal{O}$ lying above $p$, there exists a prime ideal $\mathcal{P}_{p}\supset(p)$ in $\mathcal{O}$, $\mathcal{P}_{p}^{m_{p}}||\bar{\mathfrak{a}_{p}}$, such that
-$|\mathbb{Z}\xi(\operatorname{mod}\mathcal{P}_{p}^{m_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{96C_{1}}}.$ (5.35)
-For simplicity, we assume for each $p|q_{3}$, the arising prime ideal $\mathcal{P}_{p}$ is unramified, so $v_{p}(\tilde{q})=v_{\mathcal{P}}(\tilde{q})$.
-The general case adds no additional technical challenge.
-We obtain from (5.35), $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$ and the definition of $\mathfrak{a}$ that
-$\frac{\rho_{2}c_{1}c_{3}}{96C_{1}}n<m_{p}<\rho_{2}n.$ (5.36)
-It also follows from (5.35) that for any $p|q_{2}$,
-$v_{p}(Q_{2})\leq v_{p}(\tilde{q})<v_{p}(Q_{2}^{\{1+\frac{1}{C_{1}}\}}).$ (5.37)
-5.3. Constructing an appropriate element $\zeta$
-We also need to find an element $\zeta=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ from $A_{1}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$.
-This leads us to work with the linear form
-$\mathcal{L}_{1}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$
-which gives the 2-1 entry of $\alpha_{0}\zeta-\zeta\alpha_{0}$. If $\mathcal{L}_{1}(\zeta)$ is non-zero, this certainly implies $\zeta$ and $\alpha_{0}$ are not commutative.
+$\mathscr{H}:=\{\lambda_{h}:h\in H_{2}\}.$
+From (5.42) and the fact that for an arbitrary ideal $\mathfrak{a}\subset\mathcal{O}$, the map
+$(\mathcal{O}/\mathfrak{a})^{*}\rightarrow\mathcal{O}/\mathfrak{a}:x\rightarrow x-x^{-1}$
+is $O(|\mathcal{O}/\mathfrak{a}|^{0+})$ to one, we have
+$|\mathscr{H}(\operatorname{mod}q_{1}^{\{\frac{w_{2}}{C_{1}}\}})|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{6}}.$ (5.44)
+For a construction in the next section, we apply Theorem B to the set $\mathscr{H}$, the density $\alpha=\frac{c_{1}}{6}$ and the ideal $\mathfrak{a}=(q_{1}^{\{\frac{w_{2}}{C_{1}}\}})$, which gives us an ideal $\mathcal{Q}\subset\mathfrak{a}^{C_{1}}=(q_{1}^{\{w_{2}\}})$, $\xi\in\mathcal{O}$, and some $\varepsilon_{0}>0$ depending only on $c_{1}$, such that
+$\pi_{\mathcal{Q}}(\mathbb{Z}\xi)\subset\pi_{\mathcal{Q}}\left(\sum_{C_{3}}\mathscr{H}^{C_{2}}-\sum_{C_{3}}\mathscr{H}^{C_{2}}\right),$ (5.45)
+$|\pi_{\mathcal{Q}}(\mathbb{Z}\xi)|>q_{1}^{{w_{2}\varepsilon_{0}}}.$ (5.46)
+Remark 5.47.
+In fact, following the proof of Theorem B, the ideal $\mathcal{Q}$ satisfies the property that different prime ideals $\mathcal{P}|\mathcal{Q}$ live above different natural primes $p$.
+We will assume this property of $\mathcal{Q}$ in the following discussion.
+5.4. Constructing a line
+Let $q_{2}$ be the product of all $p^{n}\|q_{1}$ where the projection of the arithmetic progression (5.46) localized at $p$ is large, i.e.
+write $\mathcal{Q}_{p}=\langle\mathcal{Q},p^{n}\rangle$, and let
+$q_{2}:=\prod_{\begin{subarray}{c}p^{n}\|q_{1}\\ |\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|>p^{nw_{2}\cdot\frac{\varepsilon_{0}}{2}}\end{subarray}}p^{n}.$ (5.48)
+The density condition (5.46) implies
+$q_{2}>q_{1}^{{\varepsilon_{0}}/{4}}.$ (5.49)
+Indeed, suppose not, then
+$|\pi_{\mathcal{Q}}(\mathbb{Z}\xi)|=\prod_{p^{n}\|q_{2}}|\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|\cdot\prod_{p^{n}\|\frac{q_{1}}{q_{2}}}|\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|\leq q_{2}^{2}(q_{1}/q_{2})^{\frac{w_{2}\varepsilon_{0}}{2}}\leq q_{1}^{w_{2}\varepsilon_{0}},$
+contradicting (5.46).
+For each $p|q_{2}$, let $\mathcal{P}_{p}|\mathcal{Q}$ be the unique prime ideal that lives above $p$ (see Remark 5.47), and $\mathcal{Q}_{p}$ be the $\mathcal{P}_{p}$ power that exactly divides $\mathcal{Q}$.
+In our case, the extension degree is at most 2, so $v_{\mathcal{P}_{p}}(p)=1$ or 2.
+The following is our goal in this section:
+Lemma 5.50.
+There is a positive integer $K_{1}$ depending only on $S$, an exact divisor $q_{4}\|q_{2}$, $q_{4}\geq q_{2}^{1/4}$, rational integers $Q_{1}|Q_{2}|q_{4}$, a traceless matrix $X=\left(\begin{matrix}a&b\\ c&-a\end{matrix}\right)\in{\text{Mat}}_{2}(\mathbb{Z})$ coprime to $q_{4}$, such that
+$1+\mathbb{Z}Q_{1}X\subset A_{0}^{K_{1}}(\operatorname{mod}Q_{2}),$
+where for each $p^{n}\|q_{4}$, we have
+$v_{p}(Q_{2})=\Theta(\rho_{2}n)$ (5.51)
+$v_{p}(Q_{2})-v_{p}(Q_{1})=\Theta(\rho_{2}n).$ (5.52)
+We first use random walk to find an element from a product of $A_{0}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$ mod $q_{2}$, where $\alpha_{0}$ is given as in Lemma 5.18, and then conjugate it by $M$ so as to obtain a non-diagonal matrix.
+The following is our goal:
+Lemma 5.53.
+There is an exact divisor $q_{4}\|q_{2}$ with $q_{4}\geq q_{2}^{\frac{1}{4}}$, an ideal $\mathscr{Q}_{1}\supset(q_{4})$, an element $\zeta\in M^{-1}A_{0}HA_{0}^{-1}H^{-1}M$, such that
+$\zeta\equiv 1+\mathfrak{q}_{1}\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}^{2})$
+where $\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)$ is a primitive matrix in $\text{Mat}_{2}(\mathcal{O})$, and $\mathfrak{q}_{1}$ is an element in $\mathcal{O}$ that uniformizes $\mathscr{Q}_{1}$.
+For each $p^{n}\|q_{4}$, we have the following estimates for valuations:
+$0.99v_{\mathcal{P}_{p}}(p^{[nw_{1}]})\leq v_{\mathcal{P}_{p}}(\mathscr{Q}_{1})\leq 1.01v_{\mathcal{P}_{p}}(p^{[nw_{1}]})$ (5.54)
+where $w_{1}$ is given as in Lemma 5.32 and
+$v_{\mathcal{P}_{p}}(Z)=O(n\rho_{1}).$ (5.55)
+Proof.
+We work with the linear form
+$\mathcal{L}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$ (5.56)
+which gives the 2-1 entry of $\alpha_{0}\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)-\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\alpha_{0}$.
+If (5.56) is non-zero, this certainly implies $\zeta$ and $\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ are not commutative.
 Let
-$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}_{1}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
-By Proposition 4.1, for any $Q|q_{2}$ ,
+$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
+By Proposition 4.1, for any sufficiently large $Q|q_{2}$ ,
 $\nu^{(2l)}(\mathcal{E}_{3}(Q))<Q^{-c_{1}},$ (5.57)
 Define
-$A_{7}=A_{1}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
-By (5.4), $\nu^{(2l)}(A_{1})>q^{-3\delta}$. We also have $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}>q^{\frac{c_{3}}{4C_{1}}}$. Then
-$\nu^{(2l)}(A_{7})>q^{-4\delta}$
-if $q_{2}^{-\rho_{1}c_{1}}<q^{-4\delta}$, or
-${\delta}<\frac{c_{1}c_{3}\rho_{1}}{16C_{1}}.$ (5.39)
-Then we produce some $q_{3}||q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{1}$ such that for any $p^{n}||q_{3}$,
-$L({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}).$ (5.40)
-From (5.40) and (5.20), we have $M\zeta_{1}M^{-1}=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$ with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}p^{[6n\rho_{1}]})$.
-Take any element $\zeta_{2}\in B_{\tilde{q}}$.
-The second inequality of (5.37) ensures that $v_{p}(\zeta_{2})=v_{p}(\tilde{q})$.
-If we let $\zeta=\zeta_{1}\zeta_{2}\zeta_{1}^{-1}$ and write $M{\zeta}M^{-1}=\left(\begin{matrix}\zeta_{11},\zeta_{12}\\ \zeta_{21},\zeta_{22}\end{matrix}\right)$. Then
-$v_{p}(M\zeta M^{-1}-1)\geq v_{p}(\tilde{q})-[6\rho_{1}n]$ (5.41)
+$A_{3}=A_{0}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
+By (5.1), (5.49) and (5.57),
+$\nu^{(2l)}(A_{3})>q^{-3\delta}-q_{2}^{-\rho_{1}c_{1}-}>q^{-3\delta}-q_{1}^{-\frac{\varepsilon_{0}\rho_{1}c_{1}}{4}-}>q^{-3\delta}-q^{-\frac{\varepsilon_{0}\rho_{1}c_{1}}{8}-}>q^{-4\delta},$
+if taking $L_{1}$ sufficiently large.
+Then using the same argument for constructing $A_{1}$ in Lemma 5.7, by taking $L_{1}$ sufficiently large, we produce some $q_{3}\|q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{3}$ such that for any $p^{n}\|q_{3}$,
+$\mathcal{L}({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}),$
+which of course implies for any $p^{n}\|q_{3}$,
+$\mathcal{L}({\zeta_{1}})\neq 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}}).$ (5.58)
+The matrix $\zeta_{1}$ itself is very non-commutative with $\alpha_{0}$ in the sense of (5.58).
+However for our purpose, we need to find a non-commutative element that is also sufficiently close to identity (at the scale $w_{1}$) as required by lemma 5.53, and then conjugate it by $M$.
+The modification is as follows: First recall the diagonalization of $\alpha_{0}$ at (5.30) and that $\lambda_{1}-\lambda_{2}=\sqrt{\text{tr}(\alpha_{0})^{2}-4}$.
+For any $p^{n}\|q_{3}$, since $v_{\mathcal{P}_{p}}(\text{tr}(\alpha_{0})^{2}-4)=O(\rho_{1}n)$, we have
+$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}}).$ (5.59)
+From (5.58) and (5.59), we have
+$\zeta_{2}=M^{-1}\zeta_{1}M=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$
+with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(\rho_{1}n)\}})$.
+Therefore, there exists an exact divisor $q_{4}\|q_{3},q_{4}\geq q_{3}^{\frac{1}{2}}$ such that one of $a_{12}$ and $a_{21}$, let us say $a_{21}$ without loss of generality, satisfies: $\forall p^{n}\|q_{4}$,
+$a_{21}\not\equiv 0(\operatorname{mod}\mathcal{P}_{p}^{\{O(n\rho_{1})\}}).$ (5.60)
+Take any element $\bar{h}\in H$ as given by Lemma 5.32 with $\bar{h}\equiv\left(\begin{matrix}a_{\bar{h}}&0\\ 0&(a_{\bar{h}})^{-1}\end{matrix}\right)(\operatorname{mod}q_{4}^{\{w_{1}+w_{2}\}})$.
+A direct computation gives
+$\zeta=[\zeta_{2},\bar{h}]=\zeta_{2}\bar{h}\zeta_{2}^{-1}{\bar{h}}^{-1}\equiv 1+\left(\begin{matrix}a_{12}a_{21}(1-(a_{\bar{h}})^{-2})&a_{11}a_{12}(1-(a_{\bar{h}})^{2})\\ a_{21}a_{22}(1-(a_{\bar{h}})^{-2})&a_{12}a_{21}(1-(a_{\bar{h}})^{2})\end{matrix}\right)(\operatorname{mod}q_{4}^{\{w_{1}+w_{2}-O(\rho_{1})\}}),$ (5.61)
+where again we have a loss of an insignificant $O(\rho_{1})$ term in the exponent of $q_{4}$ because of the denominator of $M$.
+We have the following estimate for $v_{\mathcal{P}_{p}}(\bar{h}-1)$:
+$v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}\}})\leq v_{\mathcal{P}_{p}}(\bar{h}-1)<v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}).$ (5.62)
+The second inequality of (5.62) in particular holds, because otherwise, for any $h\in H$,
+$v_{\mathcal{P}_{p}}(\lambda_{h})=v_{\mathcal{P}_{p}}(\lambda_{\bar{h}})\geq v_{\mathcal{P}_{p}}(q_{4}^{\{\frac{w_{2}}{C_{1}}\}})=v_{\mathcal{P}_{p}}(q_{1}^{\{\frac{w_{2}}{C_{1}}\}}),$ (5.63)
+recalling (5.43) and Property (iv) in Lemma 5.32.
+Following the proof of Theorem B in [TZ23], (5.63) would imply the implied arithmetic progression $\mathbb{Z}\xi$ has no contribution at the modulus $\mathcal{Q}_{p}$, i.e., $\mathbb{Z}\xi(\operatorname{mod}\mathcal{Q}_{p})=\{0\}$.
+This contradicts the thickness bound of the projection of the arithmetic progression $\mathbb{Z}\xi$ at $\mathcal{Q}_{p}$ in the definition of $q_{2}$ at (5.48).
+So (5.62) holds.
+Let $\mathscr{Q}_{1}$ be the ideal of $\mathcal{O}$ generated by $q_{4}$ and all entries of $\xi-1$.
+Write
+$\zeta\equiv 1+\mathfrak{q}_{1}\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}^{2}),$
+where $\mathfrak{q}_{1}$ is an element in $\mathcal{O}$ that uniformizes $\mathscr{Q}_{1}$, and $\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)$ is a primitive matrix in $\text{Mat}_{2}(\mathcal{O})$.
+From (5.61) and (5.60), we have for any $p^{n}\|q_{4}$,
+$v_{\mathcal{P}_{p}}(\mathscr{Q}_{1})=v_{\mathcal{P}_{p}}(\mathfrak{q}_{1})=v_{\mathcal{P}_{p}}(\bar{h}-1)\pm O(\rho_{1}n)$ (5.64)
 and
-$\min\{v_{p}(\zeta_{12}),v_{p}(\zeta_{21})\}\leq v_{p}(\tilde{q})+[10\rho_{1}n].$ (5.42)
-5.4. Creating a line
-Let
-$\left\{\vec{e}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right),\vec{f}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\vec{h}=\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right)\right\}$
-[…]
-To proceed, we recall the following well known identity.
+$v_{\mathcal{P}_{p}}(Z)=O(\rho_{1}n).$ (5.65)
+Requiring $C_{1}>100$ and taking $L_{2}$ sufficiently large then give (5.54) in Lemma 5.53.
+To proceed, we recall the following well known identity (see Lemma 6.1, [BG09]).
 Lemma 5.66.
@@
 $xyx^{-1}y^{-1}\equiv 1+xy-yx(\operatorname{mod}\mathcal{P}^{m+m^{\prime}+\min\{m,m^{\prime}\}})$
-Lemma 5.43 has an obvious generalization to a general ideal by multiplicativity.
-To apply Lemma 5.43, we take $x=M^{-1}\sigma M$ given at (5.27) and $y=M^{-1}\zeta M$, with corresponding moduli $\mathfrak{q}_{1}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})}$ and $\mathfrak{q}_{2}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})-[6\rho_{1}n]}$.
+Lemma 5.66 has an obvious generalization to general ideals by multiplicativity:
+Lemma 5.67.
+Let $\mathcal{I}_{1},\mathcal{I}_{2}$ be two ideals of $\mathcal{O}$, and let $x,y\in\text{SL}_{2}(\mathcal{O})$, $x\equiv 1(\operatorname{mod}\mathcal{I}_{1}),y\equiv 1(\operatorname{mod}\mathcal{I}_{2})$.
 Then
-$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2})$ (5.44)
+$xyx^{-1}y^{-1}\equiv 1(\operatorname{mod}\mathcal{I}_{1}\mathcal{I}_{2}),$
 and
-$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1+\tilde{q}h_{\sigma}\left(\zeta_{21}\vec{e}-\zeta_{12}\vec{f}\right)(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2}^{2})$ (5.45)
-Applying Lemma 5.43 iteratively, for any $\sigma_{1},\sigma_{2},\cdots,\sigma_{C_{3}}\in A_{6}$, we have
-$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}-1}\mathfrak{q}_{2}^{2})$ (5.46)
+$xyx^{-1}y^{-1}\equiv 1+xy-yx(\operatorname{mod}\mathcal{I}_{1}\mathcal{I}_{2}\langle\mathcal{I}_{1},\mathcal{I}_{2}\rangle).$
+To apply Lemma 5.67, we take $x=\zeta$ given by Lemma 5.53 with $\mathcal{I}_{1}=\mathscr{Q}_{1}$, and $y=h_{1}\in H$ given in Lemma 5.32 with $\mathcal{I}_{2}=\mathscr{Q}_{2}:=(q_{4}^{\{w_{1}\}})$.
+Write $\mathscr{Q}_{3}=(q_{4}^{\{w_{2}-O(\rho_{1})\}})$, where the implied constant for $O(\rho_{1})$ is taken so that $\mathscr{Q}_{3}$ divides $\langle\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle$, noticing that for each $p^{n}\|q_{4}$, $v_{\mathcal{P}_{p}}(\langle\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)=v_{\mathcal{P}_{p}}(q_{4}^{\{w_{1}\}})\pm O(\rho_{1}n)$.
+Recall $h_{1}\equiv\left(\begin{matrix}a_{h_{1}}&0\\ 0&a_{h_{1}}^{-1}\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{2}\mathscr{Q}_{3})$ and $a_{h_{1}}-a_{h_{1}}^{-1}=q_{1}^{\{w_{1}\}}\lambda_{h_{1}}$.
+Also write $\mathfrak{q}_{2}=q_{1}^{\{w_{1}\}}$, which is a uniformizer for $(q_{4}^{\{w_{1}\}})$.
+Then from Lemma 5.67,
+$\zeta^{\prime}:=\zeta h_{1}\zeta^{-1}h_{1}^{-1}\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2})$ (5.68)
 and
-$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1+\tilde{q}^{C_{3}}\left(\prod_{i=1}^{C_{3}}h_{\sigma_{i}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.47)
-where $X=\left(\begin{matrix}0&\zeta_{12}\\ (-1)^{C_{3}}\zeta_{21}&0\end{matrix}\right)$.
-The sum structure is clear. For any $\sigma_{1,1},\cdots,\sigma_{C_{2},C_{3}}\in A_{6}$, we have
-$(M^{-1}[\sigma_{1,C_{3}},\cdots[\sigma_{1,2},[\sigma_{1,1},\zeta]]M)\cdots(M^{-1}[\sigma_{C_{2},C_{3}},\cdots[\sigma_{C_{2,2}},[\sigma_{C_{{2},1}},\zeta]]\cdots]M)$
-$\equiv$ $1+\tilde{q}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.48)
-From (5.33) and (5.4), we conclude there is a constant $K_{1}$ depending only on $c_{1}$, such that
-$1+\tilde{q}^{C_{3}}\mathbb{Z}\xi X\subset M^{-1}A_{1}^{K_{1}}M(\operatorname{mod}\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}\langle\mathfrak{a}^{\prime},\mathfrak{q}_{2}\rangle)$ (5.49)
-Conjugating (5.49) back by $M$, noticing that $A_{1}^{K_{1}}$ is real and also having (5.35), (5.42), and (5.49) in mind, if we take
-$\rho_{1}<\frac{c_{1}c_{3}}{28*192C_{1}}\rho_{2},$ (5.50)
-then we obtain
+$\zeta^{\prime}$ $\equiv 1+\zeta h_{1}-h_{1}\zeta\hskip 5.69054pt(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathcal{\langle}\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)$
+$\equiv 1+(\zeta-1)h_{1}-h_{1}(\zeta-1)\hskip 5.69054pt(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathcal{\langle}\mathscr{Q}_{1},\mathscr{Q}_{2}\rangle)$
+$\Rightarrow\hskip 8.5359pt$ $\zeta^{\prime}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}\left(\begin{matrix}0&-\lambda_{h_{1}}Y\\ \lambda_{h_{1}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}\mathscr{Q}_{3}),$ (5.69)
+Next, applying Lemma 5.67 with $x=\zeta^{\prime}$, $\mathcal{I}_{1}=\mathscr{Q}_{1}\mathscr{Q}_{2}$, $y=h_{2}\in H$, $\mathcal{I}_{2}=\mathscr{Q}_{2}$, we obtain
+$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2})$
+and
+$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1+\zeta^{\prime}h_{2}-h_{2}\zeta^{\prime}\equiv 1+(\zeta^{\prime}-1)h_{2}-h_{2}(\zeta^{\prime}-1)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}),$ (5.70)
+where we notice that $\mathscr{Q}_{3}|\mathscr{Q}_{2}$.
+As $\zeta^{\prime}-1\equiv 0(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2})$ and $h_{2}\equiv\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{2}\mathscr{Q}_{3})$, we have
+$(\zeta^{\prime}-1)h_{2}-h_{2}(\zeta^{\prime}-1)$ $\equiv(\zeta^{\prime}-1)\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)-\left(\begin{matrix}a_{h_{2}}&0\\ 0&a_{h_{2}}^{-1}\end{matrix}\right)(\zeta^{\prime}-1)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3})$ (5.71)
+$\equiv\mathfrak{q}_{1}\mathfrak{q}_{2}^{2}\left(\begin{matrix}0&\lambda_{h_{1}}\lambda_{h_{2}}Y\\ \lambda_{h_{1}}\lambda_{h_{2}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}),$ (5.72)
+(5.70) and (5.71) then gives
+$\zeta^{\prime}h_{2}\zeta^{\prime-1}h_{2}^{-1}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{2}\left(\begin{matrix}0&\lambda_{h_{1}}\lambda_{h_{2}}Y\\ \lambda_{h_{1}}\lambda_{h_{2}}Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{2}\mathscr{Q}_{3}).$
+Applying Lemma 5.67 iteratively, we have for any $h_{1},h_{2},\cdots,h_{C_{3}}\in\mathscr{H}$,
+$[\cdots[[\zeta,h_{1}],h_{2}],\cdots h_{C_{3}}]\equiv 1(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}),$ (5.73)
+$[\cdots[[\zeta,h_{1}],h_{2}],\cdots h_{C_{3}}]\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\left(\begin{matrix}0&(\prod_{i=1}^{C_{3}}\lambda_{h_{i}})(-1)^{C_{3}}\cdot Y\\ (\prod_{i=1}^{C_{3}}\lambda_{h_{i}})Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathscr{Q}_{3})$ (5.74)
+The sum structure is clear. For any $h_{1,1},\cdots,h_{C_{2},C_{3}}\in\mathscr{H}$, write
+$u_{j}=[\cdots[[\zeta,h_{j,1}],h_{j,2}],\cdots h_{j,C_{3}}],\hskip 14.22636pt1\leq j\leq C_{2}.$
+We have
+$u_{1}u_{2}\cdots u_{C_{2}}\equiv 1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)\left(\begin{matrix}0&(-1)^{C_{3}}Y\\ Z&0\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathscr{Q}_{3})$ (5.75)
+From (5.46) and (5.75), we conclude there is a constant $K_{1}$ depending only on $S$ such that
+$1+\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\mathbb{Z}\xi\left(\begin{matrix}0&(-1)^{C_{3}}Y\\ Z&0\end{matrix}\right)\subset M^{-1}A_{0}^{K_{1}}M(\operatorname{mod}\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathcal{Q}^{\prime})$ (5.76)
+where $\mathcal{Q}^{\prime}=\langle\mathcal{Q},\mathscr{Q}_{3}\rangle$ with $\mathcal{Q}$ given in (5.46).
+Now for each $p^{n}\|q_{4}$, we measure the thickness of the above progression localized at $\mathcal{P}_{p}$ in terms of the valuation $v_{\mathcal{P}_{p}}$. We have
+[…]
+where for the second equality we have used that $\mathcal{Q}$ divides $(q_{4}^{\{w_{2}\}})$ and $\mathscr{Q}_{3}=(q_{4}^{\{w_{2}-O(\rho_{1})\}})$. From the definition of $q_{2}$ (5.48), we have
+$v_{\mathcal{P}_{p}}(\mathcal{Q})-v_{\mathcal{P}_{p}}(\xi)>\frac{nw_{2}\varepsilon_{0}}{4},$
+and so
+$v_{\mathcal{P}_{p}}(\mathscr{Q}_{1}\mathscr{Q}_{2}^{C_{3}}\mathcal{Q}^{\prime})-v_{\mathcal{P}_{p}}(\mathfrak{q}_{1}\mathfrak{q}_{2}^{C_{3}}\xi Z)>\frac{nw_{2}\varepsilon_{0}}{5}$ (5.77)
+by taking $L_{2}$ large.
+Collecting (5.76) and (5.77), and conjugating (5.76) back by $M$, we obtain
 Lemma 5.78.
-There exists $q_{3}||q_{1}$, $q_{3}>q^{\frac{c_{1}}{8C_{1}}}$, $q_{3}^{\prime}|q_{3}^{\prime\prime}|q_{3}$, some $K_{1}=K_{1}(c_{1})>0$, a matrix $X_{0}\in{{\text{Lie(SL}}}_{2})(\mathbb{Z})$ coprime to $q_{3}$, such that
-$1+q_{3}^{\prime}\mathbb{Z}X_{0}\subset A_{1}^{K_{1}}(\operatorname{mod}q_{3}^{\prime\prime})$ (5.52)
-where for each $p^{n}||q_{3}$, $v_{p}(q_{3}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and $\frac{c_{1}c_{3}}{192C_{1}}\rho_{2}n<v_{p}(q_{3}^{\prime\prime}/q_{3}^{\prime})<{v_{p}(q_{3}^{\prime\prime})}/2$.
-5.5. Creating a segment
-Next, we want to find $g_{1},g_{2}$ from a bounded product of $A_{1}$ so that $g_{1},g_{2}$ conjugates $X_{0}$ to different directions.
+There is a constant $K_{1}$ depending only on $S$ and $\varepsilon$, an exact divisor $q_{4}\|q_{2}$, $q_{4}\geq q^{\Omega(1)}$, an ideal $\mathcal{U}\subset\mathcal{O}$ dividing $(q_{4})$, an element $\mathfrak{u}\in\mathcal{O}$, and a primitive traceless matrix $W\in\text{ Mat}_{2}(\mathcal{O})$, such that
+$1+\mathbb{Z}\mathfrak{u}W\subset A_{0}^{K_{1}}(\operatorname{mod}\mathcal{U}),$ (5.79)
+where for each $p^{n}\|q_{4}$, we have
+$v_{\mathcal{P}_{p}}(\mathcal{U})=\Theta(\rho_{2}n),$ (5.80)
+$\frac{v_{\mathcal{P}_{p}}(\mathcal{U})}{2}\geq v_{\mathcal{P}_{p}}(\mathcal{U})-v_{\mathcal{P}_{p}}(\mathfrak{u})\geq\Omega(\rho_{2}n).$ (5.81)
+Proof of Lemma 5.50.
+Let $r_{m}\in A_{0}^{K_{1}},m\in\mathbb{Z}$ such that
+$r_{m}\equiv 1+m\mathfrak{u}W(\operatorname{mod}\mathcal{U})$ (5.82)
+as given in Lemma 5.78. Write
+$r_{1}\equiv 1+Q_{1}X(\operatorname{mod}Q_{1}^{2}).$
+where $Q_{1}|q_{4}$ and $X\in\text{Mat}_{2}(\mathbb{Z})$ is traceless and coprime to $q_{4}$. The choice for $Q_{1}$ is unique.
+For each $p^{n}\|q_{4}$, we have
+$v_{\mathcal{P}_{p}}(Q_{1})=v_{\mathcal{P}_{p}}(\mathfrak{u}).$ (5.83)
+and so
+$v_{p}(Q_{1})=\frac{v_{\mathcal{P}_{p}}(\mathfrak{u})}{v_{\mathcal{P}_{p}}(p)}.$ (5.84)
+It follows from (5.82) that
+$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}\mathcal{U}).$ (5.85)
+Taking Galois conjugate of (5.85), we obtain
+$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}\bar{\mathcal{U}})$ (5.86)
+Let $\mathcal{U}\cap\bar{\mathcal{U}}=(Q_{2})$ for some $Q_{2}\in\mathbb{N}$, $Q_{2}|q_{4}$. Then (5.85) and (5.86) gives
+$\forall m\in\mathbb{Z},\hskip 14.22636ptr_{1}^{m}r_{m}^{-1}\equiv 1(\operatorname{mod}Q_{2})$ (5.87)
+by Chinese Remainder Theorem, which implies
+$r_{m}\equiv 1+mQ_{1}X(\operatorname{mod}Q_{2}).$ (5.88)
+We also have $v_{\mathcal{P}_{p}}(Q_{2})=v_{\mathcal{P}_{p}}(\mathcal{U})$, which can be seen from the prime factorization of $\mathcal{U}$, so
+$v_{p}(Q_{2})=\frac{v_{\mathcal{P}_{p}}(Q_{2})}{v_{\mathcal{P}_{p}}(p)}=\frac{v_{\mathcal{P}_{p}}(\mathcal{U})}{v_{\mathcal{P}_{p}}(p)}.$ (5.89)
+(5.84) and (5.89) give
+$v_{p}(Q_{2})-v_{p}(Q_{1})=\frac{v_{\mathcal{P}_{p}}(\mathcal{U})-v_{\mathcal{P}_{p}}(\mathfrak{u})}{v_{\mathcal{P}_{p}}(p)}.$ (5.90)
+(5.90), (5.81), (5.84), (5.89) and (5.90) then give the desired estimates for valuations in Lemma 5.50.
+5.5. Constructing a segment
+We want to find $g_{1},g_{2}$ from a bounded product of $\mathbb{P}_{1}(A)$ so that $g_{1},g_{2}$ conjugates the arithmetic progressions given in Lemma (5.50) to different directions.
+Multiplying them together then creates a thick segment.
+Lemma 5.92 is our goal.
+Identify $\text{Lie(SL}_{2})(\mathbb{Z})$ with $\mathbb{Z}^{3}$ by $\left(\begin{matrix}a&b\\ c&-a\end{matrix}\right)\mapsto(a,b,c)^{t}$. Under this identification, write $X_{0}=(x_{1},x_{2},x_{3})$.
+Choose an element $T_{1}\in\text{SL}_{3}(\mathbb{Z})$ so that $T_{1}(x_{1},x_{2},x_{3})^{t}=(1,0,0)^{t}$.
+Define $\mathcal{L}_{1}:\mathbb{Z}^{3}\rightarrow\mathbb{Z}$ by $\mathcal{L}_{1}(a,b,c)=b$.
+Then $\mathcal{L}_{1}\circ T_{1}$ is a primitive linear form on $\mathbb{Z}^{3}$.
+We apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that
+$\text{Tr}(gX_{0}g^{-1}\eta_{1})=\mathcal{L}_{1}\circ T_{1}(gX_{0}g^{-1}).$
+Recall $c_{2}$ be the implied constant from Propositions 4.2. Then we obtain, for any sufficiently large $Q>q_{4}^{\rho_{1}}$,
+$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|\mathcal{L}_{1}\circ T_{1}(gX_{0}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
+where $l$ is as given in Proposition 2.2.
+This implies, arguing in the same way as for the construction of $A_{1}$ in §5, that there is $q_{5}\|q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{1}\in\mathbb{P}_{1}(A)$ such that for any $p^{n}\|q_{5}$,
+$\mathcal{L}_{1}\circ T_{1}(g_{1}Xg_{1}^{-1})\neq 0(p^{[2\rho_{1}n]}),$
+by taking $L_{1}$ at (5.5) sufficiently large.
+Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then we can write
+$X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5}),$
+where $a,b\in\mathbb{Z}$ with $b$ satisfying, for any $p^{n}\|q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the $\mathbb{Z}$-span of $(0,1,0)$ and $(0,0,1)$.
+Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{1})=(0,b,0)$.
+Let $\mathcal{L}_{2}:\mathbb{Z}^{3}\rightarrow\mathbb{Z}$ be the linear form getting the third component.
+Then $\mathcal{L}_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$.
+Applying Proposition 4.2 again, for any $Q>q_{5}^{\rho_{1}}$, we have
+$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|\mathcal{L}_{2}\circ T_{2}(gX_{0}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
+which then implies there is $q_{6}\|q_{5}$, $q_{6}>(q_{5})^{\frac{1}{2}}$, and $g_{2}\in\mathbb{P}_{1}(A)$ such that
+$c:=\mathcal{L}_{2}\circ T_{2}(g_{2}Xg_{2}^{-1})\not\equiv 0(p^{[2\rho_{1}n]})$
+for any $p^{n}\|q_{6}$. Therefore, for any $p^{n}\|q_{6}$,
+$\text{Det}(X_{0},X_{1},X_{2})=\text{Det}\left(\begin{matrix}1&0&0\\ 0&b&0\\ *&*&c\end{matrix}\right)\not\equiv 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.91)
+Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.50) with its $g_{1}$ and $g_{2}$ conjugates, with (5.91) in mind, we obtain
+Lemma 5.92.
+There is an exact divisor $q_{6}\|q_{4}$, $q_{6}\geq q_{4}^{1/4}$, $Q_{3}|Q_{4}|q_{6}$, and some constant $K_{2}>0$ depending only on $S$ and $\varepsilon$, such that
+$1+Q_{3}V(\operatorname{mod}Q_{4})\subset A_{0}^{K_{2}}(\operatorname{mod}Q_{4}),$ (5.93)
+where for each $p^{n}\|q_{6}$,
+$v_{p}(Q_{4})=\Theta(\rho_{2}n),$ (5.94)
+$v_{p}(Q_{4})-v_{p}(Q_{3})=\Theta(\rho_{2}n).$ (5.95)
+5.6. Proof of Proposition 5.1
+We iteratively apply Lemma 3.25 to sets starting with $E\subset A_{0}^{K_{2}}$ implied by (5.93), until we create a set $E_{0}$ such that
+$1+Q_{5}V(\operatorname{mod}Q_{6})\subset E_{0}(\operatorname{mod}Q_{6}),$
+where for each $p^{n}\|q_{6}$, $v_{p}({Q_{6}})=\Theta(\rho_{2}n)$ and
+$v_{p}{(Q_{6})}-v_{p}{(Q_{5})}\geq v_{p}(Q_{3}).$ (5.96)
+Then we iteratively taking commutators and obtain
+$E_{1}=[E_{0},E],1+Q_{3}Q_{5}V(\operatorname{mod}Q_{3}Q_{6})\subset E_{1}(\operatorname{mod}Q_{3}Q_{6});$
+$\cdots$
+$E_{k}=[E_{k-1},E],1+(Q_{3})^{k}Q_{5}V(\operatorname{mod}(Q_{3})^{k}Q_{6})\subset E_{k}(\operatorname{mod}(Q_{3})^{k}Q_{6});$
+$\cdots$
+until we reach a set with all elements congruent to identity mod $q_{6}$, which takes $O(\frac{1}{\rho_{2}})$ steps. Since
 […]
-If $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that $\text{Tr}(g\xi_{1}g^{-1}\eta_{1})=L_{1}(g)$.
-If $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, apply Proposition 4.4 with $L=L_{1}$.
-Let $c_{2}$ be the implied constant from Propositions 4.2 and 4.4.
-Then we obtain, for any $Q>q_{3}^{\rho_{1}}$, we have
-$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{1}\circ T_{1}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
-where $l$ is as given in Proposition 2.2.
-This implies, arguing the same as for the construction of $A_{3}$, that there is $q_{4}||q_{3}$, $q_{4}>(q_{3})^{\frac{1}{2}}$, and $g_{1}\in A_{1}$ such that for any $p^{n}||q_{5}$, $L_{1}\circ T_{1}(g_{1}X_{0}g_{1}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if we take
-$\delta<\frac{c_{1}c_{2}\rho_{1}}{48C_{1}}.$ (5.53)
-Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then $X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5})$, where $a,b\in\mathbb{Z}$, and $b$ satisfies, for any $p^{n}||q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the span of $(0,1,0)$ and $(0,0,1)$.
-Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{0})=(0,b,0)$.
-Let $L_{2}:\mathbb{Z}^{3}\mapsto\mathbb{Z}$ be the linear form getting the third component.
-Then $L_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$.
-Applying Proposition 4.2 again, for any $Q>q_{4}^{\rho_{1}}$, we have
-$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{2}\circ T_{2}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
-which then implies there is $q_{5}||q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{2}\in A^{l}$ such that for any $p^{n}||q_{6}$, $L_{2}\circ T_{2}(g_{2}X_{0}g_{2}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if
-${\delta}<\frac{c_{1}c_{2}\rho_{1}}{96C_{1}}.$ (5.54)
-Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.52) with its $g_{1}$ and $g_{2}$ conjugates, we obtain
-Lemma 5.55.
-There are $q_{5}||q$, $q_{5}>q^{\frac{c_{1}}{32C_{1}}}$, $q_{5}^{\prime}|q_{5}^{\prime\prime}|q_{5}$, and some constant $K_{2}>0$ depending only on $c_{1}$, such that
-$1+q_{5}^{\prime}\mathbb{Z}V\subset A_{1}^{K_{2}}(\operatorname{mod}q_{5}^{\prime\prime}),$ (5.56)
-where for each $p^{n}||q_{5}$, $v_{p}(q_{5}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and
-$\frac{c_{1}c_{3}}{384C_{1}}\rho_{2}n<v_{p}(q_{5}^{\prime\prime}/q_{5})<v_{p}(q_{5}^{\prime\prime})/2$
-if
-$\rho_{1}<\frac{c_{1}c_{3}}{3\cdot 2^{9}C_{1}}\rho_{2}.$ (5.57)
-5.6. Proof of Proposition 5.1
-Collecting all the restrictions (5.5), (5.10), (5.19), (5.29), (5.30), (5.39), (5.50), (5.53), (5.54), (5.57) on $\rho_{1},\rho_{2},\delta$, we find we can take
-$\rho_{1}=\frac{c_{1}^{2}\cdot c_{3}}{2^{8}3^{4}C_{1}}\rho_{2}$ (5.58)
-$\delta=\frac{c_{1}c_{2}c_{3}}{96C_{1}}\rho_{1}$ (5.59)
-We need the elementary lemma whose proof we leave as an exercise:
-Lemma 5.60.
-Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime.
-Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{2}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$.
-Then
-$1+p^{m_{1}+n_{1}}V\subset(H_{1}H_{2})^{4}(\operatorname{mod}p^{m_{2}+n_{2}}).$
-Proposition 5.1 is thus proved by iteratively applying Lemma 5.60 to (5.56), with $q_{*}=q_{5}$, $c=\frac{c_{1}}{32C_{1}}$ and $\rho=O_{c_{1}}(\rho_{2})$, $C=O_{S}(\log\frac{1}{\delta})$.
-6. A gluing technique for the $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ case
-In this section and the next, we prove Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
-Suppose all the assumptions about $A$ hold but the conclusion (2.4) does not hold, then by taking $\delta$ sufficiently small, we will find a constant $K$ independent of $\delta$, such that $|\pi_{q}(A^{K})|>|\Lambda_{q}|^{1-\frac{\epsilon}{2}}$.
-But this will contradict Helfgott’s inequality (5.2) when $\delta$ is sufficiently small.
-Proposition 5.1 so far only covers a large subset in one simple factor of the group and for some not too small modulus $q_{*}$.
-Apply Proposition 5.1 again with $q$ replaced by $q/q_{*}$, we will cover a large subset for some modulus $q_{*}^{\prime}$ coprime to $q_{*}$.
-This leads to considering the following question: Given a set $B\subset\Gamma_{q_{1}q_{2}},\text{gcd}(q_{1},q_{2})=1$, such that $\pi_{q_{1}}(B),\pi_{q_{2}}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, is a bounded power of $B$ covers a large subset of $\Gamma_{q_{1}q_{2}}$?
-The following proposition is the main gluing tool.
+we deduce
+$E_{0}E_{1}E_{2}\cdots(\operatorname{mod}q_{6})\supset\Gamma(Q_{5})/\Gamma(q_{6}).$
+Proposition 5.1 is thus proved with $q_{*}=q_{6}$, $c=O(1)$, $\rho=\Theta(\rho_{2})$ and $C=O(\frac{1}{\rho_{2}})$.
+6. A gluing technique for $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$
+In this section, we continue our progress towards proving Proposition 2.2.
+The set $S$ and the constant $\varepsilon>0$ in Proposition 2.2 are fixed.
+We still let $c_{1}$ and $c_{2}$ be the implied constants for $c$ in Propositions 4.1 and 4.2 respectively.
+We let $c_{0}$ be the implied constant for $c$ in Proposition 5.1.
+We free all other notations, such as $\mathbb{P}_{1},\mathbb{P}_{2},q_{1},q_{2}\cdots$ and $Q_{1},Q_{2}\cdots$ newly introduced in Section 5.
+The Landau notations $O,\Omega,\Theta$ describe asymptotic behaviours of functions with implied constants depending only on $S$.
+Proposition 5.1 so far only covers a very large subset for some not too small modulus $q_{*}$ in one simple factor of the group.
+For the purpose of proving Proposition 2.2, we need to grow $q_{*}$ to a very large divisor of $q$, and also cross from one simple factor to another simple factor.
+Our main goal in this section is to prove the following gluing tool:
 Proposition 6.1.
-Fix $0<\theta<10^{-12}$ and $10<\eta<$ 1, and suppose
-$\delta<\frac{c_{2}\eta\theta^{1/2}}{2},$ (6.2)
-where $\delta$ is given as in Proposition 2.2 and $c_{2}$ is the implied $c$ in Proposition 4.2. Take
-$L>\frac{12}{\delta}.$ (6.3)
-Let $q=\prod_{i\in I}p_{i}^{n_{i}}=q_{s}q_{l}$, where $q_{s}=\prod_{p_{i}|q,v_{p_{i}}\leq L}p_{i}^{n_{i}}$ and $q_{l}=\prod_{p_{i}|q,v_{p_{i}}>L}p_{i}^{n_{i}}$.
-Let $q_{1},q_{2}||q,q_{3}||q_{l}$, $\text{gcd}(q_{1},q_{3})=1$, and $q_{1}q_{2},q_{3}>q^{\eta}$.
-Suppose for some set $B\subset\Lambda$, we have
+Fix $0<\theta<\min\{10^{-12},(c_{0}\varepsilon)^{10}\}$. Suppose $A$ satisfies all the assumptions of Proposition 2.2 but fails (2.4) for some $\delta\leq c_{0}\varepsilon\theta$.
+Write
+$q=\prod_{i\in I}p_{i}^{n_{i}}=q_{s}q_{l},$
+where
+$q_{s}=\prod_{\begin{subarray}{c}p_{i}|q\\ n_{i}\leq\frac{3}{\delta}\end{subarray}}p_{i}^{n_{i}},\hskip 28.45274ptq_{l}=\prod_{\begin{subarray}{c}p_{i}|q\\ n_{i}>\frac{3}{\delta}\end{subarray}}p_{i}^{n_{i}}.$
+Let $q_{1},q_{2}\|q$ and $q_{3}\|q_{l}$, with $\text{gcd}(q_{1},q_{3})=1$, and $q_{1}q_{2},q_{3}>q^{\frac{\varepsilon c_{0}}{2}}$.
+Let $\tilde{q}_{3}$ be the square free part of $q_{3}$.
+Suppose for some set $B\subset\Gamma(\tilde{q}_{3})\times\Gamma\subset\Lambda$, we have
 $|\pi_{q_{1},q_{2}}(B)|>(q_{1}q_{2})^{3-\theta},$ (6.2)
 $|\pi_{q_{3},1}(B)|>q_{3}^{3-\theta}.$ (6.3)
-Then there exists $q_{3}^{*}||q_{3}$, $q_{3}^{*}>q_{3}^{\frac{1}{4}10^{-4}}$, such that
-$|\pi_{q_{1}q_{3}^{*},q_{2}}(B\cup A)^{O(\log 1/\theta)}|>(q_{1}q_{2}q_{3}^{*})^{3-O(\theta^{\frac{1}{4}})}.$ (6.6)
+Then there exists $q_{3}^{*}\|q_{3}$, $q_{3}^{*}>q_{3}^{\frac{1}{4}10^{-4}}$, such that
+$|\pi_{q_{1}q_{3}^{*},q_{2}}(B\cup B^{-1}\cup A)^{O((\log\frac{1}{\theta})^{2})}|>(q_{1}q_{2}q_{3}^{*})^{3-O(\theta^{\frac{1}{4}})}.$ (6.4)
 Remark 6.5.
-The point of Proposition 6.1 is we have an increase of modulus from $q_{1}$ to $q_{1}q_{3}^{*}$.
-To see the necessity of getting help from $A$, for instance, let $q_{1}=1$, $q_{2}=q_{3}$ and $B=\{(\lambda,\lambda),\lambda\in\Gamma\}$.
-Then $B\cdot B=B$ so there is no hope to expand only from $B$.
+The point of Proposition 6.1 is that we have an increase of modulus from $(q_{1},q_{2})$ to $(q_{1}q_{3}^{*},q_{2})$, over which the projection of $B$ is large, and the increasing speed, described by $\frac{\log q_{3}^{*}}{\log q}$, is lowered bounded by $\frac{10^{-4}}{8}\varepsilon c_{0}$}, which only depends on $S$ and $\varepsilon$.
+To see the necessity of the set $A$, for instance, we can take $q_{1}=1$, $q_{2}=q_{3}$ and $B=\{(\gamma,\gamma)\in\Lambda:\lambda\in\Gamma\}$.
+Then $B\cdot B=B$ and there is no hope to expand only from $B$.
+Remark 6.6.
+We require $q_{3}\|q_{l}$, so that the loss of size from $\Gamma_{q_{3}}$ to $\Gamma(\tilde{q_{3}})/\Gamma(q_{3})$ is not significant.
+The reason we work with $\Gamma(\tilde{q_{3}})/\Gamma(q_{3})$ instead of $\Gamma_{q_{3}}$ is that the former is a product of $p$-groups, from which we can deduce some simple properties for a general homomorphism $f:\Gamma_{q_{1}}\times\Gamma_{q_{2}}\rightarrow\Gamma(\tilde{q_{3}})/\Gamma(q_{3})$.
+Proof of Proposition 6.1.
+We have a natural isomorphism
+$\Gamma_{q_{1}q_{3}}\times\Gamma_{q_{2}}\cong\Gamma_{q_{1}}\times\Gamma_{q_{2}}\times\Gamma_{q_{3}}.$ (6.7)
+We denote the projections of RHS of (6.7) to the first, the second, and the third factor by $\mathbb{P}_{1}$, $\mathbb{P}_{2}$ and $\mathbb{P}_{3}$.
+We also denote the projection to $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $\mathbb{P}_{1,2}$.
+(Caution: to avoid confusion, these notations have different meanings from the $\mathbb{P}_{i}$’s in the statement of Proposition 5.1)
+For every $U\subset I$, write $q^{U}=\prod_{j\in U}p_{j}^{n_{j}}$.
+In this way, we identify an exact divisor of $q$ with a subset of $I$.
+Write $q_{1}=q^{I_{1}},q_{2}=q^{I_{2}},q_{3}=q^{I_{3}}$ for some $I_{1},I_{2},I_{3}\subset I$.
+Since the set $B$ satisfies (6.2) and (6.3), by Proposition 3.12, there exists $Q_{1}|q_{1},Q_{2}|q_{2},Q_{3}|q_{3}$, such that
+$Q_{1}Q_{2}<(q_{1}q_{2})^{40\theta},Q_{3}<(q_{3})^{40\theta},$ (6.8)
+and
+$\Gamma(Q_{1})/\Gamma(q_{1})\times\Gamma(Q_{2})/\Gamma(q_{2})\subset\mathbb{P}_{1,2}(B^{2880}).$ (6.9)
+$\Gamma(Q_{3})/\Gamma(q_{3})\subset\mathbb{P}_{3}(B^{2880}).$ (6.10)
+Let
+$I_{3}^{\prime}:=\{j\in I_{3}:v_{p_{j}}(Q_{3})\leq\theta^{\frac{1}{2}}v_{p_{j}}(q_{3})\}$
+The density condition (6.8) implies
+$q_{3}^{\prime}:=q^{I_{3}^{\prime}}\geq q_{3}^{1-40\theta^{\frac{1}{2}}}.$
+Let $B_{1}\subset B^{2880}$ be a set of representatives of
+$G:=\Gamma(Q_{1})/\Gamma(q_{1})\times\Gamma(Q_{2})/\Gamma(q_{2})$ (6.11)
+implied by (6.9) and $B_{2}\subset B^{2880}$ be a set of representatives of $\Gamma(Q_{3})/\Gamma(q_{3})$ implied by (6.10).
+Let $\psi:G\rightarrow B_{1}$ be inverse the map of $\mathbb{P}_{1,2}:B_{1}\rightarrow G$.
+For each $j\in I_{3}^{\prime}$, we consider the map
+$\psi_{j}:G\rightarrow\Gamma(p_{j})/\Gamma({p_{j}^{[{n_{j}\theta^{\frac{1}{4}}}]}}),\hskip 14.22636pt\psi_{j}(x)=\pi_{p_{j}^{[n_{j}\theta^{\frac{1}{4}}]}}\circ\mathbb{P}_{3}(\psi(x)).$
+According to Proposition 3.1, there are two scenarios:
+Event 1: Define
+$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)\}$
+We have,
+$|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (6.12)
+Event 2: There is a subset $S_{j}\in G$, $|S_{j}|\geq\frac{99}{100}|G|$ such that $\psi_{j}\equiv h_{j}$ over $S_{j}$}, where $h_{j}$ is a homomorphism from $G$ to $\Gamma(p_{j})/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]})$.
+In case Event 2 happens, We consider further whether $h_{j}$ is trivial at the half level:
+Event 2.1:
+$h_{j}\equiv 1(\operatorname{mod}{p_{j}^{[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]}}).$
+Event 2.2:
+$h_{j}\not\equiv 1(\operatorname{mod}{p_{j}^{[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]}}).$
+Let $I_{3}^{\prime}=J_{1}\sqcup J_{2}$ where $J_{1}$ is the collection of indices in $I_{3}^{\prime}$ where Event 1 holds, and $J_{2}$ is the complement of $J_{1}$ in $I_{3}^{\prime}$.
+In particular, for any $j\in J_{2}$, Event 2 holds.
+We further split $J_{2}=J_{3}\sqcup J_{4}$, where $J_{3}$ is the collection of indices $j$ such that Event 2.1 occurs, and $J_{4}$ is the collection of indices $j$ such that Event 2.2 occurs.
+In §6.1, §6.2 and §6.3, we divide our analysis according to which of Event 1, Event 2.1, and Event 2.2 is more popular. We will establish Proposition 6.1 for each case.
+6.1. The case $q^{J_{1}}\geq(q_{3}^{\prime})^{\frac{1}{2}}$
+By the definition of $J_{1}$, we have
+$\sum_{j\in J_{1}}(\log p_{j}^{n_{j}})|\mathcal{G}_{j}|>\log\left((q^{J_{1}})^{10^{-4}}\right)|G|^{2}.$ (6.13)
+The left hand side of (6.13) is equal to
+$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(q^{U})|\cap_{j\in U}\mathcal{G}_{j}\bigcap\cap_{j\in J_{1}-U}\mathcal{G}_{j}^{c}|.$ (6.14)
+so
+$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(q^{U})|\cap_{j\in U}\mathcal{G}_{j}|>10^{-4}\log\left(q^{J_{1}}\right)|G|^{2}.$ (6.15)
+Split the $U$-sum on LHS of (6.15) into two sums $\Sigma_{1}+\Sigma_{2}$ according to whether $\log(q^{U})>\frac{1}{2}\cdot 10^{-4}\log(q^{J_{1}})$ or not.
+Bound $\sum_{2}$ by $\frac{1}{2}\cdot 10^{-4}\log(q^{J_{1}})|G|^{2}$.
+We then obtain
+$\sum_{\begin{subarray}{c}U\subset J_{1}\\ \log q^{U}>\frac{1}{2}\cdot 10^{-4}\log(q^{J_{1}})\end{subarray}}\log(q^{U})|\cap_{j\in U}\mathcal{G}_{j}|>\frac{1}{2}\cdot 10^{-4}\log\left(q^{J_{1}}\right)|G|^{2}.$ (6.16)
+Since the number of subsets of $J_{1}$ is $q^{0+}$, (6.16) implies there exists $U_{0}\subset J_{1}$ such that
+$q^{U_{0}}>(q^{J_{1}})^{\frac{1}{2}\cdot 10^{-4}},$
+and that
+$|\cap_{i\in U_{0}}\mathcal{G}_{i}|>q^{0-}|G|^{2}.$
+Take any $(g_{1},g_{2})\in\cap_{i\in U_{0}}\mathcal{G}_{i}$, and consider $w=\psi(g_{1})\psi(g_{2})\psi(g_{1}g_{2})^{-1}$. Then $w$ satisfies,
+$\mathbb{P}_{1,2}(w)=1,$
+$\pi_{p_{j}^{[n_{j}\theta^{\frac{1}{4}}]}}\circ\mathbb{P}_{3}(w)\neq 1,\hskip 14.22636pt\forall j\in U_{0}.$
+Now we claim Lemma 3.39 implies there is a set $B_{3}\subset(B_{2}\cup\{w\})^{O(\log\frac{1}{\theta})}$, and an integer $Q_{3}^{\prime}|q_{3}^{\prime}$, such that
+$\mathbb{P}_{1,2}(B_{3})=1,$
+$\mathbb{P}_{3}(B_{3})(\operatorname{mod}q^{U_{0}})\supset\Gamma(Q_{3}^{\prime})/\Gamma({q}^{U_{0}}),$ (6.17)
+where $v_{p}(Q_{3}^{\prime})=O(\theta^{\frac{1}{4}}n)$ for each $p^{n}\|q^{U_{0}}$. Then
+$\pi_{{q}_{1}q^{U_{0}},q_{2}}(B_{1}B_{3})=|\mathbb{P}_{1,2}(B_{1})|\cdot|\mathbb{P}_{3}(B_{3})|>({q}_{1}q^{U_{0}}{q}_{2})^{3-O(\theta^{\frac{1}{4}})},$
+giving the proof of Proposition 6.1 in this case with $q_{3}^{*}=q^{U_{0}}$.
+Indeed, let $b_{i}=1.5^{i}\theta^{\frac{1}{2}},i\in\mathbb{N}$, and let $i_{0}$} be the smallest integer such that $b_{i_{0}}\geq\frac{1}{2}$.
+We have $i_{0}=O(\log\frac{1}{\theta})$.
+For each $j\in U_{0}$, we apply Lemma 3.39 iteratively with $\gamma_{0}=w$, $H=B_{2}$, $a=v_{p_{j}}(\mathbb{P}_{3}(w)-1)=O(n_{j}\theta^{\frac{1}{4}})$ fixed, and $b=b_{0}n_{j},b_{1}n_{j},\cdots,b_{i_{0}}n_{j}$.
+Multiplying the implied sets together, we obtain
+$\mathbb{P}_{1,2}([[w,B_{2}],B_{2}]^{O(\log\frac{1}{\theta})})=1$
+$\pi_{p_{j}^{n_{j}}}\circ\mathbb{P}_{3}([[w,B_{2}],B_{2}]^{O(\log\frac{1}{\theta})})\supset\Gamma(p_{j}^{O(n_{j}\theta^{\frac{1}{4}})})/\Gamma(p_{j}^{n_{j}}),\hskip 14.22636pt\forall j\in U_{0}$ (6.18)
+Using multiplicativity of $\mathbb{P}_{3}(B_{2})$, we can derive (6.1) from (6.1).
+6.2. The case $q^{J_{2}}\geq(q_{3}^{\prime})^{\frac{1}{2}},q^{J_{3}}>(q^{J_{2}})^{\frac{1}{2}}$.
+Recall for each $j\in J_{2}$, we obtain a set $S_{j}\subset G$, $|S_{j}|\geq 0.99|G|$ such that $\psi_{j}$ agrees with a homomorphism $h_{j}$ on $S_{j}$.
+Following the reasoning for obtaining $U_{0}$ in §6.1, there is a set $U_{1}\subset J_{3}$, $W_{1}:=\cap_{i\in U_{1}}S_{i}$, such that
+$q^{U_{1}}>(q^{J_{2}})^{\frac{99}{200}},$
+$|W_{1}|>(q_{1}q_{2})^{0-}|G|>(q_{1}q_{2})^{1-40\theta-},$ (6.19)
+and
+$\psi_{j}(W_{1})\equiv h_{j}(W_{1})\equiv 1(\operatorname{mod}p_{j}^{[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]}),\hskip 14.22636pt\forall j\in U_{1}.$ (6.20)
+From (6.2) and Proposition 3.12, we have
+$W_{1}^{2880}\supset\Gamma(Q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(Q_{2}^{\prime})/\Gamma(q_{2})$ (6.21)
+for some $Q_{1}^{\prime}|q_{1}$, $Q_{2}^{\prime}|q_{2}$, $Q_{1}^{\prime}Q_{2}^{\prime}<(q_{1}q_{2})^{O(\theta)}$.
+Applying Lemma 3.33 to $W_{1}^{2880}$, we obtain a set $W_{2}=[W_{1}^{2880},W_{1}^{2880}]^{O(\log\frac{1}{\theta})}$, such that
+$W_{2}\supset\Gamma(Q_{1}^{\prime\prime})/\Gamma(q_{1})\times\Gamma(Q_{2}^{\prime\prime})/\Gamma(q_{2}),$
+where $Q_{1}^{\prime\prime}=\text{gcd}(2(Q_{1}^{\prime})^{2},q_{1})$ and $Q_{2}^{\prime\prime}=\text{gcd}(2(Q_{2}^{\prime})^{2},q_{2})$, and
+$\pi_{p_{j}^{2[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]}}(\psi_{j}(W_{2}))=1,\hskip 14.22636pt\forall j\in U_{1}.$ (6.22)
+Note that we have doubled the exponent of $p_{j}$ on LHS of (6.22) compared to (6.20).
+Apply Lemma 3.33 again to $W_{2}$ and keep iterating for a total of $\log_{2}({\theta^{-\frac{1}{4}}})]+1$ times, then we obtain a set $W^{*}\subset W_{1}^{O((\log\frac{1}{\theta})^{2})}$ such that
+$\mathbb{P}_{1,2}(\psi(W^{*}))=W^{*}=\Gamma(Q_{1}^{\prime\prime\prime})/\Gamma(q_{1})\times\Gamma(Q_{2}^{\prime\prime\prime})/\Gamma(q_{2}),$ (6.23)
+$\pi_{q^{U_{1}}}\circ\mathbb{P}_{3}(\psi(W^{*}))=1,$ (6.24)
+where $Q_{1}^{\prime\prime\prime}Q_{2}^{\prime\prime\prime}\leq(Q_{1}^{\prime}Q_{2}^{\prime})^{[\theta^{-\frac{1}{4}}]}=(q_{1}q_{2})^{O{(\theta)\cdot\theta^{-\frac{1}{4}}}}=(q_{1}q_{2})^{O(\theta^{\frac{3}{4}})}$.
+Multiplying $W^{*}$ with $B_{2}$ then gives Proposition 6.1, with $q_{3}^{*}=q^{U_{1}}$, where
+$|\pi_{q_{1}q^{U_{1}},q_{2}}(B_{2}\cdot\psi(W^{*}))|=|\mathbb{P}_{1,2}(\psi(W^{*}))|\cdot|\pi_{q^{U_{1}}}(\mathbb{P}_{3}(B_{2}))|\geq(q_{1}q_{2})^{1-O(\theta^{\frac{3}{4}})}(q^{U_{1}})^{1-\theta^{\frac{1}{2}}}.$
+6.3. The case $q^{J_{2}}\geq(q_{3}^{\prime})^{\frac{1}{2}},q^{J_{4}}>(q^{J_{2}})^{\frac{1}{2}}$.
+Recall in §6.1 that for each $j\in J_{2}$, we have a set $S_{j}\subset G$, $|S_{j}|\geq 0.99|G|$ such that $\psi_{j}$ agrees with a homomorphism $h_{j}$ on $S_{j}$.
+Following the reasoning for obtaining $U_{0}$ in §6.1, there is a set $J_{5}\subset J_{4}$, $W:=\cap_{i\in J_{5}}S_{i}$, such that
+$q^{J_{5}}>(q^{J_{2}})^{\frac{99}{200}},$
+$|W|>(q_{1}q_{2})^{0-}|G|>(q_{1}q_{2})^{1-40\theta-},$
+and $\psi_{j}(W)\equiv h_{j}(W)\not\equiv 1(\operatorname{mod}p_{j}^{[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]})$ for any $j\in J_{5}$.
+The homomorphisms $\{h_{j}:j\in J_{5}\}$ uniquely determine a homomorphism
+$h:G\rightarrow\Gamma_{{(q^{J_{5}})}^{\{\theta^{\frac{1}{4}}\}}}$ (6.25)
+such that $\pi_{p_{j}^{[n_{j}\theta^{\frac{1}{4}}]}}\circ h=h_{j},\forall j\in J_{5}$.
+Since $|W|>(q_{1}q_{2})^{1-O(\theta)}$, by Proposition 3.12,
+$G^{\prime}:=\Gamma(\bar{Q}_{1})/\Gamma(q_{1})\times\Gamma(\bar{Q}_{2})/\Gamma(q_{2})\subset W^{2880}\subset G^{2880}$ (6.26)
+for some $\bar{Q}_{1}|q_{1},\bar{Q}_{2}|q_{2}$, such that
+$\bar{Q}_{1}\bar{Q}_{2}=(q_{1}q_{2})^{O(\theta)}.$ (6.27)
+Recall $B_{1}$ is a set of representatives of $G$ and $\psi:G\rightarrow B_{1}$ is a section map.
+Let $B_{4}\subset B_{1}^{2880}$ be a set of representatives of $G^{\prime}$.
+Thus we can define a section map $\bar{\psi}:G^{\prime}\rightarrow B_{4}$ such that $\mathbb{P}_{1,2}\circ\bar{\psi}$ is the identity map.
+We claim that
+$\pi_{(q^{J_{5}})^{\{\theta^{\frac{1}{4}}\}}}\circ\mathbb{P}_{3}\circ\bar{\psi}:G^{\prime}\rightarrow\Gamma_{{(q^{J_{5}})}^{\{\theta^{\frac{1}{4}}\}}}$ (6.28)
+agrees with the homomorphism $h$ at (6.25).
+Indeed, for any $g\in G^{\prime}$, write $\bar{\psi}(g)=b_{1}b_{2}\cdots b_{2880}$, where $b_{1},b_{2},\cdots,b_{2880}\in B_{1}$, and for each $1\leq i\leq 2880$.
+Write $b_{i}=\psi(g_{i})$, where $g_{i}\in G$.
+So we have $g=g_{1}\cdots g_{2880}$.
+Then
+$\pi_{(q^{J_{5}})^{\{\theta^{\frac{1}{4}}\}}}\circ\mathbb{P}_{3}\circ\bar{\psi}(g)=\pi_{(q^{J_{5}})^{\{\theta^{\frac{1}{4}}\}}}\circ\mathbb{P}_{3}(b_{1}b_{2}\cdots b_{2880})$
+$=$ $\prod_{j=1}^{2880}\pi_{(q^{J_{5}})^{\{\theta^{\frac{1}{4}}\}}}\circ\mathbb{P}_{3}\circ\psi(g_{j})=\prod_{j=1}^{2880}h(g_{j})=h(g).$
 Write
-$\Gamma_{q_{1}q_{3}}\times\Gamma_{q_{2}}\cong(\Gamma_{q_{1}}\times\Gamma_{q_{2}})\times(\Gamma_{q_{3}}\times 1).$
-Since $B$ satisfies (6.4) and (6.5), by Proposition 3.12, there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{3}^{\prime}|q_{3},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{40\theta},q_{3}^{\prime}<(q_{3})^{40\theta}$, such that
-$\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})\subset\pi_{q_{1},q_{2}}(B^{2880}).$ (6.8)
-$\Gamma(q_{3}^{\prime})/\Gamma(q_{3})\subset\pi_{q_{3},1}(B^{2880}).$ (6.9)
+$G^{\prime}=\Gamma(\bar{Q}_{1})/\Gamma(q_{1})\times\Gamma(\bar{Q}_{2})/\Gamma(q_{2})\cong\prod_{j\in I_{1}}\Gamma(p_{i}^{m_{i}^{(1)}})/\Gamma(p^{n_{i}})\times\prod_{j\in I_{2}}\Gamma(p_{i}^{m_{i}^{(2)}})/\Gamma(p^{n_{i}}).$
+The homomorphism (6.28) (which is equal to $h$) is completely factorizable, i.e. is a product of homomorphsims between local factors:
+$h_{i,j}^{(1)}:\Gamma(p_{i}^{m_{i}^{(1)}})/\Gamma(p_{i}^{n_{i}})\rightarrow\Gamma(p_{j})/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]}),i\in I_{1},j\in J_{5},$
+$h_{i,j}^{(2)}:\Gamma(p_{i}^{m_{i}^{(2)}})/\Gamma(p_{i}^{n_{i}})\rightarrow\Gamma(p_{j})/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]}),i\in I_{2},j\in J_{5}.$
+Since $\pi_{p_{j}^{[\frac{1}{2}n_{j}\theta^{\frac{1}{4}}]}}\circ h$ is nontrivial for each $j\in J_{5}$ and $(q_{1},q_{3})=1$, by Lemma 3.44, this can happen only if $j\in I_{2}$, so $q^{J_{5}}|q_{2}$.
+Moreover, the density condition (6.27) implies the set
+$J_{6}=:\{i\in J_{5}:m_{i}^{(2)}\leq\theta^{\frac{1}{2}}n_{i}\}$
+satisfies
+$q^{J_{6}}>(q^{J_{5}})^{1-O(\theta^{\frac{1}{4}})}.$ (6.29)
+Indeed, suppose not. Then
+[…]
+a contradiction since $\theta<(c_{0}\varepsilon)^{10}$.
+Now we take an element $g_{0}\in G^{\prime}$, such that
+$\mathbb{P}_{1}(g_{0})=1,$
+$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(g_{0})=1,\forall i\in I_{2}-J_{6},$
+$\mathbb{\pi}_{p_{i}^{[\frac{1}{2}n_{i}\theta^{\frac{1}{4}}]}}\circ h(g_{0})\neq 1,\forall i\in J_{6}.$
+Then from Lemma 3.44, we have
+$v_{p_{i}}(g_{0}-1)=O(\theta^{\frac{1}{4}}n_{i}),\forall i\in J_{6}.$ (6.30)
+The element $g_{0}$ is almost satisfactory for our purpose, but we want to turn the big $O$ condition in (6.30) into a big $\Theta$ condition.
+For this, we take a proper power of $g_{0}$ to control the valuations.
+Let
+$g=\left(g_{0}^{[(q^{J_{6}})^{\{\frac{1}{4}\theta^{\frac{1}{4}}\}}]}\right).$
+Then $g$ satisfies,
+$\mathbb{P}_{1}(\bar{\psi}(g))=1,$ (6.31)
+$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(\bar{\psi}(g))=1,\forall i\in I_{2}-J_{6}$ (6.32)
+$v_{p_{i}}(\mathbb{P}_{2}(\bar{\psi}(g)-1))=\Theta(\theta^{\frac{1}{4}}n_{i}),\forall i\in J_{6}.$ (6.33)
+$\frac{1}{4}n_{i}\theta^{\frac{1}{4}}\leq v_{p_{i}}(\mathbb{P}_{3}(\bar{\psi}(g)-1))\leq\frac{3}{4}n_{i}\theta^{\frac{1}{4}},\forall i\in J_{6}.$ (6.34)
 Write
-$G=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})$
-From (6.8), we can construct a section map
-$\psi:G\rightarrow B^{2880},$
-such that $\pi_{q_{1},q_{2}}(\psi(x))=x$.
-Since $q_{3}^{\prime}<q_{3}^{40\theta}$, there is $q_{4}||q_{3},q_{4}>q_{3}^{1-\theta^{\frac{1}{2}}}$ such that for every $p^{n}||q_{4}$, we have $p^{[40\theta^{\frac{1}{2}}n]}\nmid q_{3}^{\prime}$.
-Write $q_{4}=\prod_{j\in J}p_{j}^{n_{j}}$. For each $p_{j}^{n_{j}}||q_{4}$, we consider $\psi_{j}=\pi_{p_{j}^{[n_{j}\theta^{1/4}]}}\circ\mathbb{P}_{1}\circ\psi$.
-Define
-$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)\}$
-According to Proposition 3.1, there are two scenarios:
-  1. (1)
-    $|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (6.10)
-  2. (2)
-    There is a subset $S_{j}\in G$, $|S_{j}|\geq\frac{99}{100}|G|$ such that $\psi_{j}=h_{j}$ over $S_{j}$ where $h_{j}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]})$.
-Let $J=J_{1}\sqcup J_{2}$ where $J_{1},J_{2}$ is the collection of indices falling into Case (1) and Case (2), respectively. Write $q_{4}=q_{4}^{\prime}q_{4}^{\prime\prime}$, where
-$q_{4}^{\prime}=\prod_{j\in J_{1}}p_{j}^{n_{j}},\hskip 28.45274ptq_{4}^{\prime\prime}=\prod_{j\in J_{2}}p_{j}^{n_{j}}.$
-We further write $q_{4}^{\prime\prime}=q_{5}q_{5}^{\prime}$, where
-$q_{5}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}=1\end{subarray}}p_{j}^{n_{j}},\hskip 28.45274ptq_{5}^{\prime}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}\neq 1\end{subarray}}p_{j}^{n_{j}}.$
-We divide our analysis into three cases.
-6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
-In this case we have
-$\sum_{j\in J_{1}}(\log p_{j}^{n_{j}})|\mathcal{G}_{j}|>\log\left((q_{4}^{\prime})^{10^{-4}}\right)|G|^{2}$ (6.11)
-The left side of (6.11) is equal to
-$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(\prod_{j\in U}p_{j}^{n_{j}})|\cap_{j\in U}\mathcal{G}_{j}\bigcap\cap_{j\in J_{1}-U}\mathcal{G}_{j}^{c}|$ (6.12)
-Since the number of subsets of $J_{1}$ is $<q_{4}^{0+}$, there exists $J_{1}^{\prime}\subset J_{1}$, $\tilde{q}=\prod_{i\in J_{1}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime})^{\frac{1}{2}10^{-4}}$, such that
-$|\cap_{i\in J_{1}}\mathcal{G}_{i}|>q^{0-}|G|^{2}.$
-Take any $(g_{1},g_{2})\in\cap_{i\in J_{1}^{\prime}}\mathcal{G}_{i}$, and consider $\gamma_{0}=\psi(g_{1})\psi(g_{2})\psi(g_{1}g_{2})^{-1}$. Then $\gamma_{0}$ satisfies,
-$\pi_{q_{1},q_{2}}(\gamma_{0})=1,$
-and for any $p^{n}||\tilde{q}$,
-$\pi_{p^{[n\theta^{{1}/{4}}]}}\circ\mathbb{P}_{1}(\gamma_{0})\neq 1.$
-Take a set $B^{\prime}\subset B^{2880}$ which maps bijectively to $\Gamma(q_{3}^{\prime})/\Gamma(q_{3})$ under the projection $\pi_{q_{3},1}$. Take
-$B^{\prime\prime}=\{b_{1}\gamma_{0}b_{1}^{-1}\gamma_{0}^{-1}\cdot(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})\gamma_{0}(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})^{-1}\gamma_{0}^{-1},b_{1},b_{2}\in B^{\prime}\}$
-It follows from a quantitative open mapping result due to Salehi-Golsefidy (see Propostions 49, Proposition 49’ of [Gol19]) that
-$\pi_{q_{3}}\circ\mathbb{P}_{1}\left(B^{\prime\prime}\right)\supset\Gamma(q_{3}^{\{O(\theta^{1/4})\}}).$ (6.13)
-by first taking the logarithmic map to work on the Lie algebra level, then taking the exponential map back.
-The definition for $B^{\prime\prime}$ is motivated by the fact that for any $v\in\mathfrak{sl}_{2}(\mathbb{Q})$, $\mathfrak{sl}_{2}(\mathbb{Q})=[\mathfrak{sl}_{2}(\mathbb{Q}),v]+[[\mathfrak{sl}_{2}(\mathbb{Q}),v],v]$, where $+$ here means set addition.
-Obviously, $\pi_{q_{1},q_{2}}(B^{\prime\prime})=1$. Multiplying $B^{\prime\prime}$ with $B^{2880}$ and observing (6.8) we obtain Proposition 6.1 in this case by setting $q_{3}^{*}=\tilde{q}$.
-6.2. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
-The local homomorphisms $h_{j},j\in J_{2}$ can be lifted to a homomorphism
-$h:G\rightarrow\Gamma(\prod_{j\in J_{2}}p_{j})/\Gamma((q_{4}^{\prime\prime})^{\{\theta^{\frac{1}{4}}\}}).$
-Following the previous reasoning for obtaining $J_{1}^{\prime}$, there is a set $J_{2}^{\prime}\subset J_{2}$, such that
-$\bar{q}=\prod_{i\in J_{2}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime\prime})^{\frac{99}{200}},$
-$|\cap_{i\in J_{2}}S_{i}|>(q_{1}q_{2})^{-\theta}|G|,$
-and $\psi\equiv h$ on $S=\cap_{i\in J_{2}}S_{i}$.
-By Proposition 3.12, we have $S^{2880}\supset G^{\prime}=\Gamma(q_{1}^{\prime\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime\prime})/\Gamma(q_{2})$ for some $q_{1}^{\prime\prime}||q_{1}$, $q_{2}^{\prime\prime}||q_{2}$, $q_{1}^{\prime\prime}q_{2}^{\prime\prime}<(q_{1}q_{2})^{80\theta}$.
-This implies the existence of a subgroup $G^{\prime\prime}$ of $G^{\prime}$ of the form
-$G^{\prime\prime}=\Gamma(\tilde{q}_{1}(q_{1}^{*})^{\{\theta^{\frac{1}{2}}\}})/\Gamma(q_{1})\times\Gamma(\tilde{q}_{2}(q_{2}^{*})^{\{\theta^{{1}/{2}}\}})/\Gamma(q_{2}),$ (6.14)
-where
-$q_{1}=\tilde{q_{1}}q_{1}^{*},q_{2}=\tilde{q_{2}}q_{2}^{*},\tilde{q}_{1},q_{1}^{*}||q_{1},\tilde{q}_{2},q_{2}^{*}||q_{2},\tilde{q}_{1}\tilde{q}_{2}<(q_{1}q_{2})^{80\theta^{{1}/{2}}}.$ (6.15)
-Then we can define a map $\tilde{\psi}:G^{\prime\prime}\rightarrow B^{2880^{2}}$ as, for any $x\in G^{\prime\prime}$, choose a word $x=s_{1}s_{2}\cdots s_{2880}$ and let
-$\tilde{\psi}(x)=\psi(s_{1})\psi(s_{2})\cdots\psi(s_{2880}).$ (6.16)
-Then we have $\pi_{(q_{4}^{\prime\prime})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=h(x)$, and thus $\pi_{(q_{5})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=1$.
-Recall Lemma 5.43. By iteratively taking commutator of $\tilde{\psi}(G^{\prime\prime})$ $O(\log(1/\theta))$ times, we recover a group
-$\Gamma((q_{1}^{*})^{\{{2\theta^{{1}/{4}}}\}}q_{5})/\Gamma(q_{1}^{*}q_{5})\times\Gamma((q_{2}^{*})^{\{{2\theta^{{1}/{4}}}\}})/\Gamma(q_{2}^{*})\subset\pi_{q_{1}^{*}q_{5},q_{2}^{*}}(B^{O(\log(1/\theta))}),$ (6.17)
-which is identity when projected to $q_{5}$ and is large when projected to $(q_{1},q_{2}^{\prime})$.
-Recall $q_{5}||q_{3}$ and $v_{p}(q_{3}^{\prime})<v_{p}(q_{3})$ for every $p^{n}||q_{5}$. It follows from (6.15), (6.17), and (6.9) that
-$|\pi_{q_{1}q_{5},q_{2}}(B^{O(\log 1/\theta)})|>(q_{1}q_{2}q_{5})^{3-{O(\theta^{{1}/{4}})}}.$ (6.18)
-In this case, we take $q_{3}^{*}=q_{5}$ and the proposition is proved.
-6.3. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}^{\prime}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
-We begin with a lemma.
-Lemma 6.19.
-Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$.
-Then $p_{1}=p_{2}$ and $p^{x}||\xi-1$ for some $x\leq m_{1}+n_{2}-m_{2}$.
-Proof.
-We first observe that given $0\leq m_{1}<n_{1},1\leq m_{2}<n_{2}$ and two primes $p_{1},p_{2}$, if there is a nontrivial homomorphism $f$ from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{1}^{n_{2}})$, this will force $p_{1}=p_{2}$.
-To see this, if $m_{1}\neq 0$, then all elements in $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ have orders powers of $p_{1}$, and all elements in $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ have orders powers of $p_{2}$.
-Since $f$ is nontrivial, we have $p_{1}=p_{2}$.
-If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$ and we obtaine a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$.
-Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor thorough the center $Z$.
-$f^{\prime}$ can not be injective because there are elements in $\Gamma_{p_{1}}$ of order $p_{1}$, which is coprime to $p_{2}$.
-Likewise, $Z$ can not be the kernel of $f^{\prime}$ because one can also find an element $x\in\Gamma_{p_{1}}/Z$ such that the order of $x$ is coprime to $p_{2}$.
-Hence, $f^{\prime}$ as well as $f$ must be trivial and we arrive at a contradiction, so $p_{1}=p_{2}$.
-Suppose $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$.
-Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $x\leq m_{1}+n_{2}-m_{2}$.
+$\bar{\psi}(g)\equiv(1+Q_{5}X_{2},1+Q_{4}X_{1},)(\operatorname{mod}Q_{5}^{2},Q_{4}^{2}),$
+with $X_{1},X_{2}\in\text{Mat}_{2}{(\mathbb{Z})}$ primitive and traceless, $Q_{4}|q^{J_{6}}$, $Q_{5}|q^{J_{6}}$. From (6.33) and (6.34), we have
+$v_{p_{j}}(Q_{4})=\Theta(n_{i}\theta^{\frac{1}{4}})$
+$\frac{1}{4}n_{i}\theta^{\frac{1}{4}}\leq v_{p_{i}}(Q_{5})\leq\frac{3}{4}n_{i}\theta^{\frac{1}{4}},\hskip 14.22636pt\forall i\in J_{6}.$
+Let $Q_{6}=\text{gcd}(Q_{5}^{2},{({q^{J_{6}}})^{\{\theta^{\frac{1}{4}}\}}})$, recalling that $\mathbb{P}_{3}\circ\bar{\psi}$ is a homomorphism up to reduction by $({q^{J_{6}}})^{\{\theta^{\frac{1}{4}}\}}$}.
+We consider the family
+$\Xi=\{\bar{\psi}(g^{n}):n\in\mathbb{Z}\}\subset B_{4}.$ (6.35)
+Each $\bar{\psi}(g^{n})$ satisfies
+$\mathbb{P}_{1}(\bar{\psi}(g^{n}))=1$
+$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(\bar{\psi}(g^{n}))=1,\hskip 14.22636pt\forall i\in I_{2}-J_{6}$
+$\bar{\psi}(g^{n})\equiv(1+nQ_{5}X_{2},1+nQ_{4}X_{1})(\operatorname{mod}Q_{6},Q_{4}^{2})$
+with
+$v_{p_{j}}(Q_{4}),v_{p_{j}}(Q_{5})=\Theta(n_{j}\theta^{\frac{1}{4}})$
+and
+$\frac{1}{3}v_{p_{j}}(Q_{5})\leq v_{p_{j}}(Q_{6})-v_{p_{j}}(Q_{5})\leq v_{p_{j}}(Q_{5}),\hskip 14.22636pt\forall j\in J_{6}.$
+With $\Xi$ at hand, which plays the same role as the one-parameter group given by Lemma 5.50, we can run the same arguments in §5.5 and §5.6.
+Recall $\delta<\theta$.
+We start by applying Proposition 4.2 to find $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ to conjugate $X=(X_{2},X_{1})$ so that
+$\text{gcd}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}),q^{J_{6}})=(q^{J_{6}})^{O(\theta)},$
+which implies a set $J_{7}\subset J_{6}$, such that
+$q^{J_{7}}>(q^{J_{6}})^{1-O(\theta^{\frac{1}{2}})}$
+$v_{p_{j}}(\text{Det}(X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}))<n_{j}\theta^{\frac{1}{2}},\forall j\in J_{7}.$
+Then we apply the argument §5.6 to create a set $B_{5}\subset\{B_{4}\cup A\}^{O(\log\frac{1}{\theta})}$ such that
+$|\pi_{q^{J_{7}},q^{J_{7}}}(B_{5})|>(q^{J_{7}})^{6-O(\theta^{\frac{1}{4}})},$
+recalling $\pi_{q^{J_{7}},q^{J_{7}}}$ is the reduction map from $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ to $\Gamma_{q^{J_{7}}}\times\Gamma_{q^{J_{7}}}$. Therefore, $B_{5}$ has the following property:
+$\mathbb{P}_{1}(B_{5})=1$
+$\mathbb{\pi}_{p_{i}^{n_{i}}}\circ\mathbb{P}_{2}(B_{5})=1,\forall i\in I_{2}-J_{6}$
+$|\pi_{q^{J_{6}},q^{J_{6}}}(B_{5})|\geq|\pi_{q^{J_{7}},q^{J_{7}}}(B_{5})|>(q^{J_{7}})^{6-O(\theta^{\frac{1}{4}})}=(q^{J_{6}})^{6-O(\theta^{\frac{1}{4}})}.$
+So there must be some element $x_{0}\in\Gamma_{q^{J_{6}}}$ such that $B_{x_{0}}:=\{g\in B_{5}:\pi_{q^{J_{6}}}\circ\mathbb{P}_{2}(g)=x_{0}\}$ satisfies
+$|\mathbb{P}_{1,2}(B_{x_{0}})|=1$
+$|\pi_{q^{J_{6}}}\circ\mathbb{P}_{3}(B_{x_{0}})|=(q^{J_{6}})^{3-O(\theta^{\frac{1}{4}})}.$
+Then $|\pi_{q_{1}q^{J_{6}},q_{2}}(B_{1}\cdot B_{x_{0}})|=|\mathbb{P}_{1,2}(B_{1})||\pi_{q^{J_{6}}}\circ\mathbb{P}_{3}(B_{x_{0}})|>(q_{1}q_{2}q^{J_{6}})^{3-O(\theta^{\frac{1}{4}})}$.
+Proposition 6.1 is thus proved in this case as well with $q_{3}^{*}=q^{J_{6}}$.
 ∎
-Lemma 6.19 implies that $q_{5}^{\prime}||q_{2}^{*}$, since $(q_{5}^{\prime},q_{1}^{*})=1$.
-Let $G^{\prime\prime}$ and $\widehat{\psi}$ be the same as in Case 6.2.
-The following isomorphism is clear
-$G^{\prime\prime}\cong\Big{(}\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\Big{)}\times\Big{(}\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})\Big{)}\times\Big{(}\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\Big{)}$
-Since $\pi_{(q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}}\circ\mathbb{P}_{1}\circ\tilde{\psi}$ is a homomorphism and factors through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, it induces a homomorphism
-$\widehat{\psi}:\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\rightarrow\Gamma/\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}).$
-Since $\widehat{\psi}$ is nontrivial after reduction by each exact divisor of $(q_{5}^{\prime})^{\left\{\frac{1}{2}\theta^{1/4}\right\}}$, we can find $g_{1}\in\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}})/\Gamma(q_{5}^{\prime})$ such that for each $p|q_{5}^{\prime}$,
-$v_{p}(g_{1})=O(n\theta^{1/4}),\hskip 14.22636ptv_{p}(\widehat{\psi}(g_{1}))<(\frac{1}{2}n\theta^{1/4}).$
-Let $g_{2}=g_{1}^{p^{\{\frac{1}{4}n\theta^{1/4}\}}}$, then
-$v_{p}(g_{2})=\Theta(n\theta^{1/4}),\hskip 14.22636pt\frac{1}{4}n\theta^{1/4}<v_{p}(\widehat{\psi}(g_{2}))<\frac{3}{4}n\theta^{1/4}.$
-Write $g_{2}\equiv 1+Q_{1}X_{1}(\operatorname{mod}Q_{1}^{2}),\widehat{\psi}(g_{2})\equiv 1+Q_{2}X_{2}(\operatorname{mod}Q_{2}^{2})$ for some primitive $X_{1},X_{2}\in\text{Lie}(\text{SL}_{2}(\mathbb{Z}))$.
-Collecting all powers of $g_{2}$, pulling them back to $G^{\prime\prime}$ and using $\widehat{\psi}$ is induced from $\pi_{(q_{5}^{\prime})^{\theta^{1/4}}}\circ\mathbb{P}_{1}\circ\psi$ by factoring through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, we obtain
-$\pi_{q_{1}^{*}Q_{2}^{\{{4}/{3}\}},(q_{2}^{*}/q_{5}^{\prime})Q_{1}^{2}}(B^{2880^{2}})\cong\left(\pi_{q_{1}^{*},q_{2}^{*}/q_{5}^{\prime}}\times\pi_{Q_{2}^{\{4/3\}},Q_{1}^{2}}\right)(B^{2880^{2}})$
-$\supset$ $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma(({q_{2}^{*}}/{q_{5}^{\prime}})^{\{\theta^{1/2}\}})/\Gamma({q_{2}^{*}}/{q_{5}^{\prime}})\times\left(\pi_{Q_{2}^{\{{4}/{3}\}},Q_{1}^{2}}\left((1,1)+\mathbb{Z}(Q_{2}X_{2},Q_{1}X_{1})\right)\right)$ (6.20)
-with $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta^{1/4})$.
-With (6.3) at hand, which plays the same role as Theorem B creating a one-parameter group, we run the same arguments in Section 5.5, and we give a sketch here.
-We first apply Proposition 4.2 with $Q=q_{5}^{\prime}$ to find elements from $A$ to conjugate $X=(X_{1},X_{2})$ to other directions.
-By considering five primitive linear forms, we can produce $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ such that the following holds: There exists $\bar{q_{5}}||q_{5}^{\prime}$ and
-$\bar{q_{5}}>(q_{5}^{\prime})^{1-5\theta^{\frac{1}{2}}},$ (6.21)
-such that
-$(\bar{q_{5}})^{\{5\theta^{\frac{1}{2}}\}}\pi_{\bar{q_{5}}}(V\times V)\subset\pi_{\bar{q_{5}}}\text{Span}_{\mathbb{Z}}\{X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}\}$ (6.22)
-as long as $\delta<c_{2}\eta\theta^{1/2}/2$.
-Multiplying the RHS of (6.3) with its conjugates by $g_{1},g_{2},g_{3},g_{4},g_{5}$, followed by taking commutators of this product iteratively, and finally multiplying all these commutators together, we obtain
-$\Gamma((q_{1}^{*}\bar{q_{5}}))^{\{5\theta^{1/2}\}}/\Gamma(q_{1}^{*}\bar{q_{5}})\times\Gamma(((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})^{\{5\theta^{1/2}\}})/\Gamma((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})$
-$\subset$ $\pi_{q_{1}^{*}\bar{q_{5}},(q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}}}\{B\cup A\}^{O(\log 1/\theta)}$
-and thus
-$|\pi_{q_{1}{q_{5}^{\prime}},q_{2}}\{B\cup A\}^{O(\log 1/\theta)}|>(q_{1}^{*}q_{2}^{*}\bar{q_{5}})^{3(1-5\theta^{1/2})-}>(q_{1}q_{2}q_{5}^{\prime})^{3-O(\theta^{1/2})}$
-In this case, we let $q_{3}^{*}={q_{5}^{\prime}}$ and the proposition is proved.
 7. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$
-Let $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ to be determined at (7.12).
-Let $A$ be given in Proposition 2.2 but we assume
-$|\pi_{q}(A\cdot A\cdot A)|<|A|^{1+\delta}.$ (7.1)
+We free all the notations newly introduced in §5 and §6 except those in the statements of Proposition 5.1 and Proposition 6.1.
+We first prove Proposition 2.2 for all $q=\prod_{i\in I}p_{i}^{n_{i}}$ with $n_{i}>L,\forall i\in I$, where $L>0$ depends only on $S$ and $\varepsilon$ and is determined at (7.4).
+Suppose $A\subset\Lambda$ satisfies (2.3), but fails (2.4), that is, $A$ satisfies
+$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta}.$ (7.1)
 We will arrive at a contradiction when $\delta$ is sufficiently small.
-We divide our proof into two cases.
-7.1. The case $q_{s}>q^{\frac{\epsilon}{2}}$
-Take
-$\delta<\frac{\log\frac{1}{c_{0}}}{2}$ (7.2)
-for $c_{0}=c_{0}(L)$ the implied constant from Theorem A, so that
-$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|<\frac{1}{2}$
+We first observe if the first two conditions of (2.3) and (7.1) hold, then they also hold with $q$ replaced by any exact divisor $q^{\prime}$ of $q$ such that $q^{\prime}>q^{\frac{\varepsilon}{2}}$, and $\delta$ replaced by any
+$\delta_{0}\geq\frac{48\delta}{\varepsilon c_{1}}.$ (7.2)
+Indeed, (7.1) implies a similar product bound with modulus $q^{\prime}$.
+Divide $\pi_{q}(A)$ into congruence classes mod $q^{\prime}$, then at least one congruence class has cardinality $\geq\frac{|\pi_{q}(A)|}{|\pi_{q^{\prime}}(A)|}$.
+Recall by Proposition 4.1, $|\pi_{q^{\prime}}(A)|>(q^{\prime})^{\frac{c_{1}}{2}}$ if
+$\delta\leq\frac{c_{1}}{2}.$ (7.3)
+Therefore,
+$|\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)|\cdot\frac{|\pi_{q}(A)|}{|\pi_{q^{\prime}}(A)|}\leq|\pi_{q}(A)\cdot\pi_{q}(A)\cdot\pi_{q}(A)\cdot\pi_{q}(A)|\leq|\pi_{q}(A)|^{1+2\delta}$
+$\Rightarrow$ $\frac{|\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)|}{|\pi_{q^{\prime}}(A)|}\leq|\pi_{q}(A)|^{2\delta}<q^{12\delta}<(q^{\prime})^{\frac{24\delta}{\varepsilon}}<|\pi_{q^{\prime}}(A)|^{\frac{48\delta}{\varepsilon c_{1}}}\leq|\pi_{q^{\prime}}(A)|^{\delta_{0}}.$
+We take $\delta_{0}$ small enough so it satisfies the “sufficiently small” requirement for $\delta$ at Proposition 5.1. A proper choice for $\delta_{0}$ is made at (7.8). Set
+$L=\frac{3}{\delta_{0}},$ (7.4)
+so that Proposition 5.1 is applicable with $q$ replaced by all exact divisors $q^{\prime}>q^{\frac{\varepsilon}{2}}$ and $\delta$ replaced by $\delta_{0}$.
+We denote the constants $\rho_{0}=\rho(\delta_{0}),C_{0}=C(\delta_{0}),c_{0}=c$ implied by Proposition 5.1.
+We first apply Proposition 5.1 to obtain $q_{1}\|q,q_{1}\geq q^{c_{0}}$ such that
+$\Gamma(q_{1}^{\{\rho_{0}\}})/\Gamma(q_{1})\subset\mathbb{P}_{1}(A)^{C_{0}}(\operatorname{mod}q_{1}).$
+If $q_{1}>q^{1-\frac{\varepsilon}{2}}$, we are done.
+Otherwise, apply Proposition 5.1 again, with $q$ replaced by $\frac{q}{q_{1}}$, we obtain $q_{2}^{\prime}\|\frac{q}{q_{1}}$, $q_{2}^{\prime}\geq(\frac{q}{q_{1}})^{c_{0}}\geq q^{\frac{\varepsilon c_{0}}{2}}$, such that
+$\Gamma((q_{2}^{\prime})^{\{\rho_{0}\}})/\Gamma(q_{2}^{\prime})\subset\mathbb{P}_{1}(A)^{C_{0}}(\operatorname{mod}q_{2}^{\prime}).$
+We will take
+$\delta\leq c_{0}\epsilon\rho_{0},$ (7.5)
+so we can apply Proposition 6.1 to obtain $q_{2}^{\prime\prime}\|q_{2}$, $q_{2}^{\prime\prime}>(q_{2}^{\prime})^{\frac{1}{4}10^{-4}}$, such that
+$\left|\pi_{q_{1}q_{2}^{\prime\prime}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}[\log\frac{1}{\rho_{0}}]^{2}})\right|>(q_{1}q_{2}^{\prime\prime})^{3-K_{2}\rho_{0}^{\frac{1}{4}}}$
+for some two absolute constants $K_{1},K_{2}>0$ implied by the two big O notations in (6.4).
+Write $q_{2}=q_{1}q_{2}^{\prime\prime}$.
+If $q_{2}>q^{1-\frac{\varepsilon}{2}}$, we are done.
+Otherwise, we apply Proposition 5.1 and Proposition 6.1 again to find $q_{2}\|q_{3}\|q$, with $\frac{q_{3}}{q_{2}}>q^{\frac{10^{-4}\varepsilon c_{0}}{8}}$, such that
+$\left|\pi_{q_{3}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}^{2}[\log\frac{1}{\rho_{0}}]^{4}})\right|>q_{3}^{3-K_{2}^{\frac{5}{4}}\rho_{0}^{\frac{1}{16}}},$
+and we keep going, until we find an integer $q_{T}\|q$, $q_{T}>q^{1-\frac{\varepsilon}{2}}$, with
+$\left|\pi_{q_{T}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}^{T}[\log\frac{1}{\rho_{0}}]^{2T}})\right|=|\pi_{q_{T},1}(\cdots)|>q_{T}^{3-K_{2}^{2}\rho_{0}^{(\frac{1}{4})^{T}}}.$ (7.6)
+The number of iterations $T$ is bounded by $\frac{8\cdot 10^{4}}{\varepsilon c_{0}}$.
+Now we apply Proposition 5.1 and Proposition 6.1 to the $A$-power at 7.6 to expand on the second modulus.
+Iterate for another $T^{\prime}\leq\frac{8\cdot 10^{4}}{\varepsilon c_{0}}$ times, we obtain an integer $q_{T^{\prime}}^{\prime}\|q$, $q_{T^{\prime}}^{\prime}>q^{1-\frac{\varepsilon}{2}}$, such that
+$\left|\pi_{q_{T},q_{T^{\prime}}^{\prime}}(A^{C_{0}K_{1}^{T+T^{\prime}}[\log\frac{1}{\rho_{0}}]^{2(T+T^{\prime})}})\right|>(q_{T}q_{T^{\prime}}^{\prime})^{3-K_{2}^{2}\rho_{0}^{(\frac{1}{4})^{T+T^{\prime}}}}.$ (7.7)
+The exponent for $A$ at (7.7) is upper bounded by
+$T_{0}:=C_{0}K_{1}^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}[\log\frac{1}{\rho_{0}}]^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}},$
+and the exponent for $q_{T}q_{T^{\prime}}^{\prime}$ is lower bounded by
+$3-K_{2}^{2}\rho_{0}^{[(\frac{1}{4})^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}]}.$
+We take $\delta_{0}$ small enough so that $\delta_{0}$ satisfies the “sufficiently small” requirement for $\delta$ at Proposition 5.1 and the implied constant $\rho_{0}=\rho_{0}(\delta_{1})=O(\delta_{1})$ satisfies
+$K_{2}^{2}\rho_{0}^{[(\frac{1}{4})^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}]}<\frac{\varepsilon}{2}.$ (7.8)
+This determines $T_{0}$ and $L$ subsequently.
+(7.7) then implies
+$|\pi_{q}(A)|^{T_{0}}>q^{6-5\varepsilon}>q^{\varepsilon}|\pi_{q}(A)|>|\pi_{q}(A)|^{1+\frac{\varepsilon}{6}}$
+[…]
+Thus recalling (7.2), (7.3) and (7.5), if we set
+$\delta=\min\{\frac{\varepsilon c_{1}\delta_{0}}{48},\varepsilon c_{0}\rho_{0},\frac{c_{1}}{2},\frac{\varepsilon}{6T_{0}}\},$
+then (7) contradicts (7.1). So (2.4) has to hold for the above choice of $\delta$. We have thus proved Proposition 2.2 for all sufficiently large $q$ with exponents of all prime divisors $\geq L$.
+For a general $q$, let $q=q_{s}q_{l}$, where
+$q_{s}=\prod_{i\in I,n_{i}\leq L}p_{i}^{n_{i}},\hskip 28.45274ptq_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}.$
+If $q_{s}\leq q^{\frac{\varepsilon}{2}}$, then the modulus $q_{s}$ can be ignored. We can work with $q_{l}$ and run the previous argument.
+If $q_{s}>q^{\frac{\varepsilon}{2}}$, take
+$\delta<\frac{c_{s}}{4}$ (7.10)
+for $c_{s}=c_{s}(L)$ the implied spectral gap from Theorem A for $\Lambda$ and for all integer moduli with exponents of prime divisors $\leq L$. Then
+$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|\leq\frac{1}{2|\Gamma_{q_{s}}|}$
 for any $x\in\Gamma_{q_{s}}$. Since $\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(A)>q^{-\delta}$, we have
-$|\pi_{q_{s}}(A)|>|\Gamma_{q_{s}}|q^{-\delta}>q_{s}^{6-\frac{2\delta}{\epsilon}-}>q_{s}^{6-2\rho}$ (7.3)
+$|\pi_{q_{s}}(A)|>\frac{|\Gamma_{q_{s}}|}{2}q^{-\delta}>q_{s}^{6-\frac{2\delta}{\varepsilon}-}>q_{s}^{6-2\rho_{0}}$ (7.11)
 if we take
-$\delta<\min\{\frac{\log\frac{1}{c_{0}}}{2},{\epsilon\rho}\}.$ (7.4)
-for some $\rho\ll\epsilon$ to be determined at (7.9).
-Furthermore, if $q_{s}>q^{1-\frac{\epsilon}{2}}$, then (7.3) implies the assumption in (2.3) is void and so Proposition 2.2 automatically holds. So now we assume $q_{s}<q^{1-\epsilon/2}$.
-Now we apply Proposition 5.1 with $q$ replaced by $q_{l}$, then we obtain some $q^{\prime}||q_{l},q^{\prime}>q_{l}^{c}>q^{\frac{\epsilon c}{2}}$ for some $c$ only depending on the generating set $S$, and $C^{\prime}=O_{S}(\log\frac{1}{\rho})$ such that
-$|\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C^{\prime}})|>(q^{\prime})^{3-\rho}$
-if
-$\delta<c^{\prime}\epsilon\rho$ (7.5)
-for some $c^{\prime}$ depending only on $S$.
-Now we apply Proposition 6.1 with $B=A^{C^{\prime}}$, $q_{1}=q_{2}=q_{s}$, $q_{3}=q^{\prime}$ and $\theta=\rho$.
-We then obtain $q^{\prime\prime}|q^{\prime},q^{\prime\prime}>(q^{\prime})^{\frac{1}{4}\times 10^{-4}}$, and
-$|\pi_{q_{s}q^{\prime\prime},q_{s}}(A^{C^{\prime}C^{\prime\prime}})|>(q_{s}^{2}q^{\prime\prime})^{3-O(\theta^{\frac{1}{4}})},$ (7.6)
-for some
-$C^{\prime\prime}=O(\log{1}/{\rho}).$ (7.7)
-By (7.7), we have increased the modulus of the first component from $q_{s}$ to $q_{s}q^{\prime\prime}$, at the cost of a density loss from $3-\rho$ to $3-O(\rho^{{1}/{4}})$.
-If $q_{s}q^{\prime\prime}<q^{1-\frac{\epsilon}{2}}$, we apply Proposition 5.1 again with $q$ replaced by $\frac{q}{q_{s}q^{\prime\prime}}$ and $\theta=O(\rho^{1/4})$, after that we can apply Proposition 6.1 again to increase the modulus to a larger one.
-We apply Proposition 5.1 and Proposition 6.1 iteratively until we reach a modulus $q_{1}^{\star}||q,q_{1}^{\star}>q^{1-\frac{\epsilon}{2}}$.
-Next, we go through the same procedure to increase the modulus of the second component to $q_{2}^{\star}||q,q_{2}^{\star}>q^{1-\frac{\epsilon}{2}}$.
-The number of such iterations is bounded by by $\frac{K_{0}\log{(1/\epsilon)}}{c}$ for some absolute constant $K_{0}$.
-In the end, we obtain
-$\left|\pi_{q_{1}^{\star},q_{2}^{\star}}\left(A^{O_{S}(\log\frac{1}{\rho})}\right)\right|>(q_{1}^{\star}q_{2}^{\star})^{3-O(\rho^{(\frac{1}{4})^{\frac{K_{0}\log(1/\epsilon)}{c}}})}.$ (7.8)
-Let $K_{1},K_{2}$ be the implied constants on the LHS and RHS of (7.8), respectively. If we take
-$\rho=e^{-(\log(3K_{2}/\epsilon)4^{{\frac{K_{0}\log(1/\epsilon)}{c}}})},$ (7.9)
-then from (7.8),
-$\left|\pi_{q_{1},q_{2}}\left(A^{K_{1}(\log\frac{1}{\rho})}\right)\right|>q^{6-5\epsilon}.$ (7.10)
-Recall also $\pi_{q}(A)<q^{6-6\epsilon}$. Then (7.10) and (5.13) imply
-$|\pi_{q}(A\cdot A\cdot A)|>|A|^{1+\frac{\epsilon}{6K_{1}\log(1/\rho)}}.$ (7.11)
-We take any $\delta_{0}$ sufficiently small satisfying (7.2), (7.4) and (7.5) for $\delta$. Then we take
-$L=\left[\frac{13}{\delta_{0}}\right]$ (7.12)
-in view of the condition (6.3). Finally, we set
-$\delta=\min\{\delta_{0},\frac{\epsilon}{6K_{1}\log(1/\rho)}\}.$
-Then (7.11) contradicts (7.1). Proposition 2.2 is thus proved in this case.
-7.2. The case $q_{s}<q^{\frac{\epsilon}{2}}$
-In this case the modulus $q_{s}$ can be ignored. We can directly apply Proposition 5.1 and start the iterating process. The analysis is identical to the previous case.
+$\delta<\min\{\frac{c_{s}}{4},\frac{{\varepsilon\rho_{0}}}{2}\}.$ (7.12)
+where $\rho_{0}$ is given after (7.4).
+In case $q_{s}>q^{1-\frac{\varepsilon}{2}}$, (7.11) implies the third assumption in (2.3) is void and so Proposition 2.2 automatically holds.
+In case $q_{s}\leq q^{1-\frac{\varepsilon}{2}}$, we can run the argument in the large exponent case to grow the modulus $(q_{s},q_{s})$ to $(q_{1}^{*},q_{2}^{*})$, where $q_{1}^{*},q_{2}^{*}>q^{1-\frac{\varepsilon}{2}}$ and the projection of product set of $A$ to $(q_{1}^{*},q_{2}^{*})$ is very large.
 8. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$
@@
 We give a sketch.
-We write $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ depending only on the generating set $S$ and $\epsilon$.
-We take $\delta$ sufficiently small and assume
-$|\pi_{q}(A\cdot A\cdot A)|<|\pi_{q}(A)|^{1+\delta}.$ (8.1)
-If $q_{s}>q^{1-\epsilon/2}$, $q_{l}$ can be ignored and we can apply Theorem A to conclude Proposition 2.2.
-If $q^{\epsilon/2}<q_{s}<q^{1-\epsilon/2}$, we first apply Theorem A to find some $C_{1}$ so that $A^{C_{1}}(\operatorname{mod}q_{s})$ is large in $\Lambda/\Lambda(q_{s})$.
-Then we use the method in Section 7 to find some $C_{2}$ so that $\pi_{q,q_{s}}(A^{C_{2}})$ is large in $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q_{s}\mathbb{Z})^{2}$, and so by Proposition 3.20, there exists $q_{1}||q,q_{2}||(q_{1},q_{s}),q_{1}q_{2}>(qq_{s})^{1-\epsilon/2}$ such that
-$\pi_{q_{1},q_{2}}(A^{C_{2}})\supset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ (8.2)
-for some $\rho=O(\delta)$. Choose $B\subset A^{C_{2}}$ a representative set of RHS of (8.2) and consider the section map
-$\psi:(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})\rightarrow B.$
-If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is not a homomorphism, then one can apply the argument in Section 6.1 find a large divisor $q_{3}||\text{gcd}(q_{l},q_{1})$ and an element $(g_{0},w_{0})$ from a bounded product of $B$ such that $\pi_{q_{1},q_{2}}((g_{0},w_{0}))=(1,0)$, and $v_{p}(w_{0})=O(\rho n^{1/2})$ for any $p^{n}||q_{3}$.
-Conjugating $(g_{0},w_{0})$ by $B$ using the formula $(g,x)\cdot(1,y)\cdot(g,x)^{-1}=(1,g(y)),$ and multiplying these conjugates by $B$, we obtain a set which is large in $\Lambda/\Lambda(q_{1})\ltimes(\mathbb{Z}/q_{2}q_{3})^{2}$ under the reduction $\pi_{q_{1},q_{2}q_{3}}$.
-If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is a homomorphism, then we can find a large divisor $q_{4}||\text{gcd}(q_{l},q_{1})$ and a large subset $Z\subset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ such that $\psi|_{Z}$ is a homomorphism.
-But then Proposition 3.24 implies $\psi(Z)(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})\subset H(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})$ for some proper algebraic subgroup $H\subset\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ defined by two linear equations.
-Then using Proposition 4.4 and arguing as in Section 5.1 for creating the set $A_{3}$, we can find a large exact divisor $q_{5}||q_{4}$ and $(g_{1},w_{1})$ from a bounded product of $B\cup(A\cdot A)$, such that $\pi_{q_{1},q_{2}}((g_{1},w_{1}))=(1,0)$, and $v_{p}(w_{1})=O(\rho^{1/2}n)$ for any $p^{n}||q_{5}$.
-With the help of $(g_{1},w_{1})$, we can boundedly generate a set which is large in $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}q_{5}\mathbb{Z})^{2}$.
-In either of the above two cases, we increase the modulus $q_{2}$ by a factor of size at least $q^{\kappa}$ where $\kappa>0$ depending only on $S$ and $\epsilon$, Iterating the above procedure by controlled number of times, we can grow $q_{2}$ to a modulus $q_{6}||q_{1}$ and $q_{6}>q_{1}^{1-\epsilon/2}$, which will lead to a contradiction to (8.1) when $\delta$ is sufficiently small.
-If $q_{s}<q^{\epsilon/2}$, then $q_{s}$ can be ignored and the expanding procedure is virtually identical to the previous case.
+Let $\mathbb{P}_{0}:\Lambda\rightarrow\Gamma$ be the projection to the semisimple part and $\mathbb{P}_{u}:\Lambda\rightarrow\mathbb{Z}^{2}$ the projection to the unipotent part.
+We assume $q=\prod_{i\in I}p_{i}^{n_{i}}$ with each $n_{i}$ large.
+Assume (2.3) holds but (2.4) fails.
+Then by taking $\delta$ sufficiently small, we can apply Proposition 5.1 and Proposition 6.1 iteratively to obtain
+$G:=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\subset\mathbb{P}_{0}(A^{O((\log\frac{1}{\rho})^{2})}),$ (8.1)
+where $\mathbb{P}_{0}$ is the projection to the semisimple part, $\rho$ is a small quantity, $q_{1}=q^{I_{1}}$ for some $I_{1}\subset I$ with $q_{1}>q^{1-\frac{\varepsilon}{2}}$, and for any $i\in I_{1}$, $v_{p_{i}}(q_{1}^{\prime})<\rho v_{p_{i}}(q_{1})$.
+Let $B\subset A^{O((\log\frac{1}{\rho})^{2})}$ be a set of representatives of $G$ implied by (8.1). Let $\psi:G\rightarrow B$ such that $\mathbb{P}_{0}\circ\psi:G\rightarrow G$ is the identity map.
+For each $i\in I_{1}$, consider $\psi_{i}=\pi_{p_{i}^{\{\theta n_{i}\}}}\circ\mathbb{P}_{u}\circ\psi$ for another small quantity $\theta$ which is larger than $\rho$.
+According to Proposition 3.1, there are two scenarios:
+Event 1: Define
+$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{i}(xy)\neq\psi_{i}(x)\psi_{i}(y)\}$
+We have,
+$|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (8.2)
+Event 2: There is a subset $S_{i}\in G$, $|S_{i}|\geq\frac{99}{100}|G|$ such that $\psi_{i}\equiv h_{i}$ over $S_{i}$}, where $h_{i}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{i}^{[n_{i}\theta]})$.
+We split $I_{1}=I_{2}\sqcup I_{3}$, where $I_{2}$ consists of all $i$ such that Event 1 happens, and $I_{3}$ is the complement of $I_{2}$.
+We divide our analysis into two cases according to whether $q^{I_{2}}$ is large or not.
+Case 1: $q^{I_{2}}\geq(q^{I_{1}})^{\frac{1}{2}}$.
+We can find $(1,w)\in B^{2}B^{-1}$ such that $w\in\mathbb{Z}^{2}$ satisfies $v_{p_{i}}(w)<\theta n_{i}$ for all $i\in I_{2}$.
+Conjugate $(1,w)$ by $B$ using the formula
+$(g,x)\cdot(1,w)\cdot(g,x)^{-1}=(1,g(w)),$
+we obtain a set $B_{1}$ such that $\mathbb{P}_{0}(B_{1})=1$ and $\mathbb{P}_{u}(B_{1})(\operatorname{mod}q^{I_{2}})$ is large in $(\mathbb{Z}/q^{I_{2}}\mathbb{Z})^{2}$.
+Therefore, $\pi_{q,q^{I_{2}}}(B\cdot B_{1})$ is a large subset of $\Gamma_{q}\ltimes(\mathbb{Z}/q^{I_{2}}\mathbb{Z})^{2}$.
+Case 2: $q^{I_{3}}\geq(q^{I_{1}})^{\frac{1}{2}}$.
+In this case $\pi_{q,q^{I_{3}}}\circ\psi$ agrees with a homomorphism $h:G\rightarrow\Gamma_{q}\ltimes(\mathbb{Z}/(q^{I_{3}})^{\{\theta\}}\mathbb{Z})^{2}$ on a large subset $T$ of $G$, where $T$ boundedly generates a large subgroup $G^{\prime}$ of $G$.
+According to Proposition 3.24,
+$\pi_{(q^{I_{3}})^{\{\frac{\theta}{4}\}},(q^{I_{3}})^{\{\frac{\theta}{4}\}}}(h(G))\subset\pi_{(q^{I_{3}})^{\{\frac{\theta}{4}\}},(q^{I_{3}})^{\{\frac{\theta}{4}\}}}(H_{u,v})$
+for some $u,v\in\mathbb{Q}$.
+Next, we apply Proposition 4.3 to find $(g_{0},w_{0})$ from $A\cdot A$ such that $g_{0}\in G^{\prime}$ and
+$\pi_{p_{i}^{[\frac{1}{4}\theta n_{i}]}}(g_{0},w_{0})\not\in H_{u,v}(\operatorname{mod}p_{i}^{[\frac{1}{4}\theta n_{i}]}),\hskip 14.22636pt\forall i\in I_{3}^{\prime},$
+where $I_{3}^{\prime}\subset I_{3}$ and $q_{3}^{\prime}\geq(q^{I_{3}})^{\frac{1}{2}}$.
+On the other hand, since $T$ boundedly generates $G^{\prime}$, take $(g_{0},w_{1})\in(\psi(T))^{O(1)}\subset B^{O(1)}$. Let $(1,w^{\prime})=(g_{0},w_{0})(g_{0},w_{1})^{-1}$. Then
+$w^{\prime}\not\equiv 0(\operatorname{mod}p_{j}^{[\frac{1}{4}\theta n_{j}]}),\forall j\in I_{3}^{\prime}.$
+With $(1,w^{\prime})$ playing the same role as $(1,w)$ in the previous case, we can argue in the same way to construct a large subset of $\Gamma_{q}\ltimes(\mathbb{Z}/q_{3}^{\prime}\mathbb{Z})$.
+In either case, we manage to construct a subset from a bounded product of $\{A\cup B\}$ whose projection to $\Gamma_{q}\ltimes(\mathbb{Z}/q^{\prime}\mathbb{Z})^{2}$ is very large for some large $q^{\prime}\|q$, starting with a set $B$ only known to have large projection to $\Gamma_{q}\ltimes\{\vec{0}\}$.
+We can keep iterating the above argument to grow the modulus of the unipotent factor, and thus prove Proposition 2.2.